### v2

# The Julia-Wolff-Carathéodory theorem in convex finite type domains

Abstract.
Rudin’s version of the classical Julia-Wolff-Carathéodory theorem is a cornerstone of holomorphic function theory in the unit ball of $\mathbb{C}^{d}$. In this paper we obtain a complete generalization of Rudin’s theorem for a holomorphic map $f\colon D\to D^{\prime}$ between convex domains of finite type. In particular, given a point $\xi\in\partial D$ with finite dilation we show that the $K$-limit of $f$ at $\xi$ exists and is a point $\eta\in\partial D^{\prime}$, and we obtain asymptotic estimates for all entries of the Jacobian matrix of the differential $df_{z}$ in terms of the multitypes at the points $\xi$ and at $\eta$. We introduce a generalization of Bracci-Patrizio-Trapani’s pluricomplex Poisson kernel which, together with the dilation at $\xi$, gives a formula for the restricted $K$-limit of the normal component of the normal derivative $\langle df_{z}(n_{\xi}),n_{\eta}\rangle$. Our principal tools are methods from Gromov hyperbolicity theory, a scaling in the normal direction, and the strong asymptoticity of complex geodesics. To obtain our main result we prove a conjecture by Abate on the Kobayashi type of a vector $v$, proving that it is equal to the reciprocal of the line type of $v$, and we give new extrinsic characterizations of both $K$-convergence and restricted convergence to a point $\xi\in\partial D$ in terms of the multitype at $\xi$.
2020 Mathematics Subject Classification. Primary: 32A40. Secondary: 32H40, 32F18.
Key words and phrases: convex domain, finite type, boundary behavior of holomorphic maps
1 Partially supported by INdAM, by PRIN Real and Complex Manifolds: Geometry and Holomorphic Dynamics n. 2022AP8HZ9, and by the MUR Excellence Department Project MatMod@TOV CUP:E83C23000330006
2 Partially supported by the European Union (ERC Advanced grant HPDR, 101053085 to Franc Forstnerič), the research program P1-0291 from ARIS, Republic of Slovenia, and by the MUR Excellence Department Project MatMod@TOV CUP:E83C23000330006
1. Introduction
The Julia–Wolff–Carathéodory theorem is a classical result in the theory of one complex variable collecting results obtained by the three authors in the decade 1920-1930 [44, 62, 29]. The starting point is what is nowadays known as the Julia lemma, which is a boundary version of the Schwarz lemma. Let $\mathbb{D}\subset\mathbb{C}$ be the unit disc, let $\xi\in\partial\mathbb{D}$, and let $f\colon\mathbb{D}\to\mathbb{D}$ be a holomorphic map. Assume that, as $z\to\xi$, the image $f(z)$ goes to the boundary $\partial\mathbb{D}$ at least as fast as $z$, meaning that the dilation
$\lambda_{\xi}:=\liminf_{z\to\xi}\frac{1-|f(z)|}{1-|z|}$
is finite [^1] Notice that the dilation is never 0.. The Julia lemma says that every horosphere of radius $R>0$ centered at $\xi$ (which is the open disc of Euclidean radius $R/(R+1)$ internally tangent to $\xi$) is mapped by $f$ in a horosphere of radius $\lambda_{\xi}R$ centered at a point $\eta\in\partial\mathbb{D}$, and as an immediate consequence $f$ has non-tangential limit
$\angle\lim_{z\to\xi}f(z)=\eta,$
that is, $f$ has limit $\eta$ along any sequence converging to $\xi$ inside a cone with vertex $\xi$ and aperture $\pi$. One could say that the dilation is playing the role of the absolute value of the derivative of $f$ at $\xi$, even if $f$ is in general not even continuous at $\xi$. The Julia–Wolff–Carathéodory theorem shows that this intuition is correct.
Theorem 1.1 (Julia–Wolff–Carathéodory).
Let $f\colon\mathbb{D}\to\mathbb{D}$ be a holomorphic self-map and let $\xi\in\partial\mathbb{D}$ be a point with finite dilation $\lambda$. Then there exists $\eta\in\partial\mathbb{D}$ such that
$\angle\lim_{z\to\xi}f(z)=\eta,\quad\angle\lim_{z\to\xi}f^{\prime}(z)=\lambda_{\xi}\eta\bar{\xi}.$
In his book on function theory in the unit ball [59], Rudin proved in 1980 a remarkable generalization of this theorem for holomorphic maps $f\colon\mathbb{B}^{d}\to\mathbb{B}^{q}$. There are two main issues which make this setting radically different from the case of the unit disc. The first (obvious) one, is that instead of dealing with just a derivative, we have now a whole Jacobian matrix to consider, and as we will see the asymptotics of different entries will be different.
The second issue, more subtle, is that in $\mathbb{B}^{d}$ there are two relevant generalizations of the concept of non-tangential convergence to a boundary point, and both are different from the concept one would at first expect, that is, convergence inside a cone with vertex $\xi$ and aperture $\pi$. Indeed, as first remarked in 1969 by Korányi [47] (see also Stein [48, 60]) in his work on Fatou’s theorem in several complex variables, such definition of convergence is not natural and too restrictive in the unit ball, ultimately due to fact that the Kobayashi distance $k_{\mathbb{B}^{d}}$ of the ball behaves differently in the complex tangential directions w.r.t. the complex normal direction, in other words it is anisotropic. Instead, Korányi introduced a natural concept of convergence to $\xi$: a sequence $K$-converges to $\xi$ if it is contained in a $K$-region with vertex $\xi$. Such regions have an anisotropic shape: they look like cones near $\xi$ if intersected with the complex normal line, but they are tangent to the boundary $\partial\mathbb{B}^{d}$ in the complex tangential directions. $K$-convergence of a sequence $(z_{n})$ can be simply characterized by looking at the normal and complex tangent components of $z_{n}-\xi$, which we denote $N(z_{n})$ and $T(z_{n})$ respectively. Indeed $z_{n}\stackrel{{\scriptstyle K}}{{\to}}\xi$ iff
$N(z_{n})=O(\delta_{\mathbb{B}^{d}}(z_{n})),\quad T(z_{n})=O(\delta_{\mathbb{B}^{d}}(z_{n})^{\frac{1}{2}}),$ (1.1)
where $\delta_{\mathbb{B}^{d}}$ denotes the distance from the boundary (see Remark 7.12 for a discussion about this characterization). The Julia lemma easily generalizes to this setting: if a point $\xi$ has finite dilation then there exists a point $\eta\in\partial\mathbb{B}^{q}$ such that $f(z_{n})\to\eta$ for every sequence $z_{n}\stackrel{{\scriptstyle K}}{{\to}}\xi$, that is, $f$ has $K$-limit $\eta$ at $\xi$.
The second relevant notion of convergence is slightly more restrictive but still more general than convergence inside cones. It stems out of the Lindelöf principle, which is a main ingredient in the proof of the Julia–Wolff–Carathéodory theorem. The classical Lindelöf principle in one variable states that a bounded holomorphic function defined on $\mathbb{D}$ which admits limit along a continuous curve with endpoint $\xi\in\partial\mathbb{D}$ actually admits non-tangential limit at $\xi$. Surprisingly enough, the $K$-limit is not the right generalization of the non-tangential limit to consider for a Lindelöf principle in the ball. There are indeed bounded holomorphic functions defined on $\mathbb{B}^{d}$ which have limit on the radial segment with endpoint $\xi\in\partial\mathbb{B}^{d}$ but do not have $K$-limit at $\xi$, a simple example with $\xi=(1,0)$ is given by
$f(z,w)=\frac{w^{2}}{1-z^{2}},$ (1.2)
since
$f(t,\lambda\sqrt{1-t^{2}})=\lambda^{2},\quad t\in[0,1),\lambda\in\mathbb{D}.$
However Čirka [30] proved that if $f$ admits limit along some “special” curve (e.g. the radial segment) with endpoint $\xi$, then there exists $L\in\mathbb{C}$ such that $f(z_{n})\to L$ for every sequence $(z_{n})$ which $K^{\prime}$-converge [^2] Such sequences are also called restricted. to $\xi$, that is every sequence $(z_{n})$ such that
$N(z_{n})=O(\delta_{\mathbb{B}^{d}}(z_{n})),\quad T(z_{n})=o(\delta_{\mathbb{B}^{d}}(z_{n})^{\frac{1}{2}}).$ (1.3)
One says that $h$ has $K^{\prime}$-limit $L$ at $\xi$. It is easy to see that the existence of the $K$-limit implies the existence of the $K^{\prime}$-limit, which in turn implies the existence of the non-tangential limit defined with cones.
Given two functions $f,g\colon\mathbb{B}^{d}\to\mathbb{C}$, with $g$ never zero, we write $f=O_{K}(g)$ if the quotient $f/g$ is bounded on every $K$-region with vertex $\xi$, and we write $f=o_{K^{\prime}}(g)$ if $f/g$ has $K^{\prime}$-limit 0 at $\xi$. We are now in condition to state Rudin’s generalization of the Julia–Wolff–Carathéodory theorem, describing the asymptotics of all entries of the Jacobian matrix of $f$ and of its Jacobian determinant.
Theorem 1.2 (Rudin [59]).
Let $f\colon\mathbb{B}^{d}\to\mathbb{B}^{q}$ be a holomorphic map and let $\xi\in\partial\mathbb{B}^{d}$ such that
$\lambda_{\xi}:=\liminf_{z\to\xi}\frac{1-\|f(z)\|}{1-\|z\|}<+\infty.$
Then there exists $\eta\in\partial\mathbb{B}^{q}$ such that $f$ has $K$-limit $\eta$ at $\xi$. Moreover, if $\xi,v_{1},\dots,v_{d-1}$ and $\eta,u_{1},\dots,u_{q-1}$ are orthonormal bases of $\mathbb{C}^{d}$ and $\mathbb{C}^{q}$ respectively, then
  • (i)
    $\langle df_{z}(\xi),\eta\rangle=O_{K}(1)$ and has $K^{\prime}$-limit $\lambda_{\xi}$ at $\xi$;
  • (ii)
    $\langle df_{z}(v_{j}),\eta\rangle$ is both $O_{K}(\delta_{\mathbb{B}^{d}}(z)^{\frac{1}{2}})$ and $o_{K^{\prime}}(\delta_{\mathbb{B}^{d}}(z)^{\frac{1}{2}})$ for all $j$;
  • (iii)
    $\langle df_{z}(\xi),u_{i}\rangle$ is both $O_{K}(\delta_{\mathbb{B}^{d}}(z)^{-\frac{1}{2}})$ and $o_{K^{\prime}}(\delta_{\mathbb{B}^{d}}(z)^{-\frac{1}{2}})$ for all $i$;
  • (iv)
    $\langle df_{z}(v_{j}),u_{i}\rangle=O_{K}(1)$ for all $i,j$.
Finally if $d=q$, then $\det(df_{z})=O_{K}(1).$
Notice that the vectors $(v_{j})_{j=1}^{d-1}$ (resp. $(u_{i})_{i=1}^{q-1}$) span the complex tangent space $T_{\xi}^{\mathbb{C}}\partial\mathbb{B}^{d}$ (resp. $T_{\eta}^{\mathbb{C}}\partial\mathbb{B}^{q}$). Rudin also shows that the exponents $\frac{1}{2}$ and $\frac{1}{2}$ appearing in his result are sharp. One may wonder about their geometrical significance, and we will give an answer to this question below.
It is natural to try to generalize this result to other bounded domains in several complex variables. In 1990 Abate [2, 3] generalized Rudin’s theorem to the case of a holomorphic map $f\colon D\to D^{\prime}$ between bounded strongly convex domains with $C^{3}$ boundary. A crucial point of his approach is the following: instead of working with objects defined in Euclidean terms, he redefines the dilation, horospheres, $K$-regions and $K^{\prime}$-convergence in terms of the Kobayashi distance $k_{D}$, so as to fully exploit the fact that $f$ does not expand the Kobayashi distance. As an example, the dilation $\lambda_{\xi,p,p^{\prime}}$ at a point $\xi\in\partial D$ is defined by
$\log\lambda_{\xi,p,p^{\prime}}=\liminf_{z\to\xi}k_{D}(p,z)-k_{D^{\prime}}(p^{\prime},f(z)),$ (1.4)
where $p\in D,p^{\prime}\in D^{\prime}$ are given base-points. Notice that, if $f\colon\mathbb{B}^{d}\to\mathbb{B}^{q}$ is a holomorphic map and $\xi\in\partial\mathbb{B}^{d}$, then a simple computation shows
$\lambda_{\xi}:=\liminf_{z\to\xi}\frac{1-\|f(z)\|}{1-\|z\|}={\rm exp}\left(\liminf_{z\to\xi}{k_{\mathbb{B}^{d}}}(z,0)-k_{\mathbb{B}^{q}}(f(z),0)\right)=\lambda_{\xi,0,0}.$ (1.5)
Another important tool is Lempert’s theory of complex geodesics, which for example yields the existence of well-defined horospheres in strongly convex domains with $C^{3}$ boundary. Other domains of $\mathbb{C}^{d}$ have been considered, such as polydiscs [4] and strongly pseudoconvex domains [3, 25]. A Julia-Wolff-Carathéodory theorem was recently obtained [26, 7] for infinitesimal generators of holomorphic semigroups in the unit ball, see also [56].
We are interested in the case of convex domains of finite type. The D’Angelo finite type condition [32] plays a capital role in several complex variables, due to its equivalence to the existence of subelliptic estimates for the $\bar{\partial}$-Neumann problem [46, 34, 28]. For a convex domain with $C^{\infty}$ boundary, McNeal [52] (see also Boas-Straube [17]) showed that the D’Angelo finite type condition is equivalent to the finite line type condition, which we now recall. If $\xi\in\partial D$ and $v$ is a nonzero vector in $\mathbb{C}^{d}$, then the line type $m_{\xi}(v)$ of $v$ is the order of vanishing of the defining function of $D$ along the complex affine line $\zeta\mapsto\xi+\zeta v$. The line type of $\xi$ is defined as $\sup_{v\in\mathbb{C}^{d}\setminus\{0\}}m_{\xi}(v)$. The domain $D$ has finite line type if there exists $L\geq 2$ such that the line type of all $\xi\in\partial D$ is $\leq L$. Every vector in $\mathbb{C}^{d}\setminus T^{\mathbb{C}}_{\xi}\partial D$ has line type 1, while every nonzero complex tangent vector has line type $\geq 2$. In strongly convex domains every nonzero complex tangent vector has line type 2. In this paper we consider only the line type.
A partial Julia–Wolff–Carathéodory theorem in this setting was obtained by Abate-Tauraso [8] in 2002. They studied the case of a holomorphic function $f\colon D\to\mathbb{D}$ defined on a bounded convex domain of finite type and with values in the unit disc. Their result needs technical assumptions: $D$ is assumed to be a convex domain of finite type with $C^{\infty}$ boundary which is also strictly linearly convex, meaning that the complex tangent $T_{\xi}^{\mathbb{C}}\partial D$ intersects the boundary $\partial D$ only at the point $\xi$. They need to assume the existence of a complex geodesic $\varphi\colon\mathbb{D}\to D$ with $\varphi(1)=\xi$ such that the radial limit $\lim_{t\to 1}\varphi^{\prime}(t)$ exists finitely. Notice that even in the “egg” domain $\mathbb{E}_{m}:=\{z\in\mathbb{C}^{2}:|z_{0}|^{2}+|z_{1}|^{m}<1\}\subset\mathbb{C}^{2}$ there are complex geodesics with $\varphi(1)=(1,0)$ such that $\|\varphi^{\prime}(t)\|$ explodes as $\mathbb{R}\ni t\to 1^{-}$, see Example 4.8. Finally their result is expressed in terms of regions which are smaller than the $K$-regions (which they call $T$-regions). Their result highlights however a very interesting phenomenon: as $z\to\xi$ the asymptotic behavior of the partial derivative $\frac{\partial f}{\partial v}(z)$ with $v\in\mathbb{C}^{d}\setminus\{0\}$ is controlled by the asymptotic behavior of the Kobayashi–Royden metric $\kappa_{D}(z,v)$, that is, if $s>0$ is such that $\kappa_{D}(z,v)=O_{K}(\delta_{D}(z)^{-s}),$ then
$\frac{\partial f}{\partial v}(z)=O_{K}(\delta_{D}(z)^{1-s}).$
Define the Kobayashi type at $s_{\xi}(v)$ of a nonzero vector $v\in\mathbb{C}^{d}$ as the infimum of the set
$\{s>0:\kappa_{D}(z,v)=O_{K}(\delta_{D}(z)^{-s})\}.$
Abate–Tauraso [8] show that every vector in $\mathbb{C}^{d}\setminus T^{\mathbb{C}}_{\xi}\partial D$ has Kobayashi type 1, and that for all nonzero complex tangent vector $v$ one has $\frac{1}{L}\leq s_{\xi}(v)\leq 1-\frac{1}{L},$ where $L$ is the line type of the point $\xi$. In [5] Abate conjectures that the Kobayashi type equals the reciprocal of the line type [^3] Abate considers the D’Angelo type in the direction $v$ but by a result of Boas–Straube [17] this coincides with the line type in the direction $v$ $m_{\xi}(v)$, and leaves as an open question whether the infimum is attained.
In this paper we prove Abate’s conjecture, we show that the infimum is attained, and we give a complete generalization of Rudin’s theorem for a holomorphic map $f\colon D\to D^{\prime}$ between convex domains of finite line type, with minimal regularity assumptions: we only require that there exists $L\geq 2$ such that $\partial D$ is $C^{L}$ and has line type $\leq L$ at all points of the boundary, and analogously for $D^{\prime}$ (we actually need these assumptions only locally near $\xi$ and $\eta$ but for the sake of clarity we will ignore this here). As in Rudin’s theorem we are able to describe the asymptotics of all entries of the Jacobian matrix of $f$ and of its Jacobian determinant. Before stating our main result, we need to choose the right bases in which to write the Jacobian matrix. Given $\xi\in\partial D$, denote $n_{\xi}$ the outer normal versor of $\partial D$ at $\xi$. Set $v_{0}=n_{\xi}$ and complete it to an orthonormal basis $(v_{j})_{j=0}^{d-1}$ of $\mathbb{C}^{d}$ which respects the multitype flag introduced by Yu in [61] (see Section 6 for a detailed description of this construction). For a point $\eta\in\partial D^{\prime}$ we choose analogously an orthonormal basis $(u_{i})_{i=0}^{q-1}$ of $\mathbb{C}^{q}$ with $u_{0}=n_{\eta}$. We denote the multitypes at $\xi$ and $\eta$ as $(m_{j})_{j=0}^{d-1}:=(m_{\xi}(v_{j}))$ and $(n_{i})_{i=0}^{q-1}:=(m_{\eta}(u_{i}))$. They coincide with the multitypes in the sense of Catlin if $\partial D$ is $C^{\infty}$ [61].
Theorem 1.3.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be bounded convex domains of finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map. Assume that the dilation at $\xi$ is finite. Then there exists $\eta\in\partial D$ such that $f$ has $K$-limit $\eta$ at $\xi$, and for all $i,j$,
$\langle df_{z}(v_{j}),u_{i}\rangle=O_{K}\biggl(\delta_{D}(z)^{\frac{1}{n_{i}}-\frac{1}{m_{j}}}\biggr).$ (1.6)
Moreover
  • (i)
    $\langle df_{z}(n_{\xi}),n_{\eta}\rangle$ has $K^{\prime}$-limit $\alpha>0$ at $\xi$;
  • (ii)
    $\langle df_{z}(v_{j}),n_{\eta}\rangle=o_{K^{\prime}}\biggl(\delta_{D}(z)^{1-\frac{1}{m_{j}}}\biggr)$ for all  $1\leq j\leq d-1$;
  • (iii)
    $\langle df_{z}(n_{\xi}),u_{i}\rangle=o_{K^{\prime}}\biggl(\delta_{D}(z)^{\frac{1}{n_{i}}-1}\biggr)$ for all  $1\leq i\leq q-1$.
Finally if $d=q$, then $\det(df_{z})=O_{K}\biggl(\delta_{D}(z)^{\sum_{j=0}^{d-1}\frac{1}{n_{j}}-\frac{1}{m_{j}}}\biggr).$
Some remarks are in order. We prove the following formula for the $K^{\prime}$-limit $\alpha$ of $\langle df_{z}(n_{\xi}),n_{\eta}\rangle$:
$\alpha=\lambda_{\xi,p,p^{\prime}}\frac{\Omega^{D}_{\xi}(p)}{\Omega^{D^{\prime}}_{\eta}(p^{\prime})},$
where $\Omega^{D}_{\xi}(z)\colon D\to\mathbb{R}$ is a generalization of Bracci–Patrizio–Trapani’s pluricomplex Poisson kernel in strongly convex domains [22, 23]. We define
$\Omega^{D}_{\xi}(z):=-\frac{1}{\angle\lim_{\zeta\to 1}\langle\varphi^{\prime}(\zeta),n_{\xi}\rangle},$
where $\varphi$ is a complex geodesic such that $\varphi(0)=z$ and endpoint $\xi$. Notice that in general the non-tangential limit of $\varphi^{\prime}$ does not necessarily exist finite, but we show that the non-tangential limit of its normal component does. Moreover, it is not known whether such complex geodesic is unique, so we have to prove that this is well-defined.
Our result gives an interpretation for the geometrical significance of the exponents $\frac{1}{2}$ and $\frac{1}{2}$ in Rudin’s theorem: they are obtained as $1-\frac{1}{2}$ and $\frac{1}{2}-1$ respectively, due to the fact that in the ball the type of all nonzero complex tangential vectors is 2. It is also interesting to notice that for convex domains of finite type the Jacobian $\det df_{z}$ and the entries $\langle df_{z}(v_{j}),u_{i}\rangle$ with $i,j\geq 1$ are in general no longer $O_{K}(1)$, instead their behavior now depends on the multitypes at $\xi$ and $\eta$. As an example, if the harmonic mean of the multitype at $\eta$ is strictly smaller than the harmonic mean of the multitype at $\xi$, then
$K\textrm{-}\lim_{z\to\xi}\det df_{z}=0.$
Another point to stress is that the choice of the bases $(v_{j})$ and $(u_{i})$ is essential in this result. One could indeed expect that for any nonzero vectors $v\in\mathbb{C}^{d},u\in\mathbb{C}^{q}$ one has
$\langle df_{z}(v),u\rangle=O_{K}\biggl(\delta_{D}(z)^{\frac{1}{m_{\eta}(u)}-\frac{1}{m_{\xi}(v)}}\biggr),$
but this is false in general (see Example 12.9). Instead it follows from our result that
$\langle df_{z}(v),u\rangle=O_{K}\biggl(\delta_{D}(z)^{\frac{1}{M_{\eta}(u)}-\frac{1}{m_{\xi}(v)}}\biggr),$
where $M_{\eta}(u)$ is a dual notion of type called cotype that we introduce.
Finally, as a consequence of the minimality of our regularity assumptions, Theorem 1.3 extends to the case of $C^{2}$ boundary Abate’s Julia–Wolff–Carathéodory theorem for strongly convex domains [2]. Notice that in this case Lempert’s theory of complex geodesics is not available.
Our tools consist principally in methods from Gromov hyperbolic theory and a scaling technique along the normal direction, and both are new in the context of the Julia–Wolff–Carathéodory theorem. Gromov hyperbolicity theory was introduced in several complex variables by Balogh-Bonk [14]. Interactions between the two fields have been recently flourishing [37, 16, 64, 20, 21, 35, 11], we cite in particular Zimmer’s proof of the Gromov hyperbolicity of convex domains of finite type endowed with the Kobayashi distance [63], and the proof of the Julia lemma for proper geodesic Gromov hyperbolic metric spaces ([12], see also [13]). The scaling technique is taken from Gaussier [36], and is weighted with the multitype at the boundary point. This is used in [12] to prove strong asymptoticity of complex geodesics and the existence of horospheres in convex domains of finite type. Due to the fact that we assume finite type only locally around $\xi\in\partial D$$\eta\in\partial D^{\prime}$ we need to prove local versions of these statements. The study of strong asymptoticity also allows us to perform several key computations along the normal direction instead of along a complex geodesic, bringing considerable simplification. We will use the scaling method in particular to prove Abate’s conjecture in [5] (Theorem 8.2) and the following extrinsic characterization of both $K$-convergence and $K^{\prime}$-convergence, which will be important in the proof of our main theorem. Let $D\subset\mathbb{C}^{d}$ be a bounded convex domain of finite type, $\xi\in\partial D$, and let be $(v_{j})_{j=0}^{d-1}$ be a basis of $\mathbb{C}^{d}$ chosen as in Theorem 1.3.
Theorem 1.4 (see Theorem 7.9).
Let $(z_{n})$ be a sequence in $D$ converging to $\xi$. Then
  • (i)
    $z_{n}\stackrel{{\scriptstyle K}}{{\to}}\xi$ if and only if
    $\langle z_{n}-\xi,v_{j}\rangle=O\left(\delta_{D}(z_{n})^{1/m_{j}}\right),\quad\forall\,0\leq j\leq d-1;$
  • (ii)
    $z_{n}\stackrel{{\scriptstyle K^{\prime}}}{{\to}}\xi$ if and only if $\langle z_{n}-\xi,n_{\xi}\rangle=O(\delta_{D}(z_{n}))$, and
    $\langle z_{n}-\xi,v_{j}\rangle=o\left(\delta_{D}(z_{n})^{1/m_{j}}\right),\quad\forall\,1\leq j\leq d-1.$
Compare this with (1.1) and (1.3). Once again, the choice of the basis $(v_{j})$ is essential in this result.
Acknowledgements. We want to thank the anonimous Referee for the careful reading of the paper and for many useful comments.
2. Preliminaries
We use the following notations:
  • We denote by $(e_{j})_{j=0}^{d-1}$ the canonical basis of $\mathbb{C}^{d}$.
  • For $u,v\in\mathbb{C}^{d}$ we denote with $\langle u,v\rangle$ the standard Hermitian product of $\mathbb{C}^{d}$ and $\|u\|:=\sqrt{\langle u,u\rangle}$ the standard Euclidean norm of $\mathbb{C}^{d}$.
  • Let $\mathbb{D}:=\{\zeta\in\mathbb{C}:|\zeta|<1\}$ and $\mathbb{B}^{d}:=\{z\in\mathbb{C}^{d}:\|z\|<1\}.$ We denote by $\mathbb{H}:=\{\zeta\in\mathbb{C}:{\operatorname{Re}\,}\zeta<0\}$ the left half-plane. We will use the left half-plane instead of the more conventional right (or upper) half-plane because in this way the normal outer vector at the origin is 1, which makes the computations cleaner.
  • Denote
    $\mathscr{C}(\zeta)=\frac{\zeta-1}{\zeta+1},\quad\mathscr{C}^{-1}(\zeta)=\frac{1+\zeta}{1-\zeta}.$
    The biholomorphism $\mathscr{C}\colon\mathbb{D}\to\mathbb{H}$ is the Cayley transform which sends $1\in\partial\mathbb{D}$ to $0\in\partial\mathbb{H}$.
  • If $D\subsetneq\mathbb{C}^{d}$ is a domain and $p\in\mathbb{C}^{d}$ let
    $\delta_{D}(p):=\min\{\|p-\xi\|:\xi\in\partial D\}.$
  • If $D\subset\mathbb{C}^{d}$ is a Kobayashi hyperbolic domain, we denote by $k_{D}$ its Kobayashi distance and by $\kappa_{D}$ its Kobayashi–Royden metric. For all $z\in D$ and $r>0$, we denote
    $B_{D}(z,r):=\{w\in D:k_{D}(z,w)<r\}.$
  • By curve in $\mathbb{C}^{d}$ we always mean a continuous function $\gamma\colon I\to\mathbb{C}^{d}$ defined on an interval $I$.
We use the following normalizations (see e.g. Kobayashi [45]) for the Poincaré metric and distance in the disc $\mathbb{D}\subset\mathbb{C}$:
$\kappa_{\mathbb{D}}(z,v)=\frac{2|v|}{1-|z|^{2}},\quad k_{\mathbb{D}}(0,z)=\log\frac{1+|z|}{1-|z|}=2\,{\rm arctanh}(|z|).$
As a consequence, if $D\subset\mathbb{C}^{d}$ is a domain, the Kobayashi–Royden pseudometric is defined as follows: if $z\in D$, $v\in\mathbb{C}^{d}$,
$\kappa_{D}(z,v)=\inf\left\{\frac{2}{R}\,:\,R>0,\exists\varphi\in\textrm{Hol}(\mathbb{D},D),\varphi(0)=z,\,\varphi^{\prime}(0)=Rv\right\}.$ (2.1)
An important advantage of this choice of normalization is that, as a result, the definition of a central concept such as the dilation of a map at a boundary point is given by the same exact formula (1.4) in both the complex case and in the metric non-expanding case. In other words, the dilation of a holomorphic map $f$ coincides with the dilation of $f$ considered as a non-expanding map between metric spaces, see also [12, 13]. Some authors normalize differently the Poincaré metric in the disc, dividing it by 2. As a consequence they define the Kobayashi–Royden pseudometric $\kappa_{D}(z,v)$ as the above formula (2.1) divided by 2. Similar considerations hold for the Kobayashi distance.
We now review some basic definitions for convex domains of $\mathbb{C}^{d}$.
Definition 2.1 ($\mathbb{C}$-proper).
A convex domain $D\subset\mathbb{C}^{d}$ is $\mathbb{C}$-proper if $D$ does not contain a complex affine line.
Remark 2.2.
Harris [40] and Barth [15] (see also [24]) showed that for a convex domain $D\subset\mathbb{C}^{d}$ the following are equivalent:
  1. (1)
    $D$ is $\mathbb{C}$-proper,
  2. (2)
    $D$ is Kobayashi hyperbolic,
  3. (3)
    $D$ is complete Kobayashi hyperbolic.
Definition 2.3 (Line type).
Let $D\subseteq\mathbb{C}^{d}$ be a convex domain. Let $L\geq 2$, and assume that $\partial D$ is of class $C^{L}$ in a neighborhood of a boundary point $\xi\in\partial D$. Let $r$ be a defining function of class $C^{L}$ for $\partial D$ in a neighborhood of $\xi$. We say that the point $\xi$ has finite (line) type $L$ if:
  1. (1)
    for every $v\in\mathbb{C}^{d}\setminus\{0\}$ the function $\zeta\mapsto r(\xi+\zeta v)$, defined on $\mathbb{C}$, has order of vanishing smaller than or equal to $L$ at 0 (we denote such number as $m_{\xi}(v)$ and we call it the type of $v$ at $\xi$), and
  2. (2)
    there exists $v\in\mathbb{C}^{d}\setminus\{0\}$ with type $m_{\xi}(v)=L$.
Notice that the type $m_{\xi}(v)$ of a vector is necessarily an even number.
We say that a point $\xi\in\partial D$ has locally finite (line) type if there exist $L\geq 2$ and a neighborhood $U$ of $\xi$ such that $\partial D\cap U$ is of class $C^{L}$ and every point in $\partial D\cap U$ has line type at most $L$. Finally, we say that the convex domain $D$ has finite (line) type if there exists $L\geq 2$ such that the boundary $\partial D$ is of class $C^{L}$ and every point has line type at most $L$.
Remark 2.4.
At a strongly convex boundary point $\xi\in\partial D$ the line type is $2$. The same is true at a strongly pseudoconvex boundary point of $D$. If $\partial D$ is $C^{\infty}$ then McNeal [52] proved that $\xi$ has finite line type $L$ if and only if it has finite D’Angelo type $L$. It then follows from [33, Theorem 2 p. 131] that if $\partial D$ is $C^{\infty}$ then $\xi$ has finite line type if and only if it has locally finite line type.
In this paper we consider only the line type, hence we will call it type tout-court.
3. Gromov hyperbolicity methods
Zimmer [63] proved that if $D\subset\mathbb{C}^{d}$ is a bounded convex domain of finite type, then the metric space $(D,k_{D})$ is Gromov hyperbolic and its Gromov compactification is canonically homeomorphic to $\overline{D}$. Gromov hyperbolicity is a coarse notion of negative curvature that has numerous applications in complex geometry (see for example [37, 16, 64, 20, 21, 35, 11, 12, 13]). In this section we review the metric properties of bounded convex domains of finite type inherited from Gromov hyperbolicity, and then we show how to adapt them to our setting (that is, near a point of locally finite type of a $\mathbb{C}$-proper convex domain) via a localization result.
Definition 3.1.
(Geodesics, almost-geodesics, quasi-geodesics) Let $D\subset\mathbb{C}^{d}$ be a Kobayashi hyperbolic domain. A geodesic is a map $\gamma$ from an interval $I\subset\mathbb{R}$ to $D$ which is an isometry with respect to the Euclidean distance on $I$ and the Kobayashi distance on $D$, that is for all $s,t\in I$,
$k_{D}(\gamma(s),\gamma(t))=|t-s|.$
If the interval is $\mathbb{R}_{\geq 0}$ (resp. closed and bounded) we call $\gamma$ a geodesic ray (resp. geodesic segment).
We say that a map $\gamma\colon\mathbb{R}_{\geq 0}\to D$ is an almost-geodesic ray if for all $\varepsilon>0$ there exists $T_{\varepsilon}>0$ such that for all $s,t\geq T_{\varepsilon}$,
$|t-s|-\varepsilon\leq k_{D}(\gamma(s),\gamma(t))\leq|t-s|.$
Now fix $A\geq 1$, $B\geq 0$. A (non-necessarily continuous) map $\gamma\colon\mathbb{R}_{\geq 0}\to D$ is a $(A,B)$ quasi-geodesic ray if for every $t,s\geq 0$
$A^{-1}|t-s|-B\leq k_{D}(\gamma(s),\gamma(t))\leq A|t-s|+B.$
Clearly every almost-geodesic is a $(1,B)$ quasi-geodesic for some $B\geq 0$.
Finally, a map $\gamma\colon[0,T)\to D$, where $T\in(0,+\infty]$, is said to have endpoint $\gamma(T)=\xi\in\partial D$ if $\lim_{t\to T}\gamma(t)=\xi$.
Remark 3.2.
If $D\subset\mathbb{C}^{d}$ is a $\mathbb{C}$-proper convex domain and $\xi\in\partial D$ is a point of locally finite type, then it follows by the generalized Hopf-Rinow theorem that any two points of $D$ can be connected by a geodesic segment, and it follows from Proposition 4.3 below that, given any point $p\in D$, there exists a geodesic ray $\gamma$ with endpoint $\xi$ and with $\gamma(0)=p$.
Definition 3.3.
(Gromov product) Let $D\subset\mathbb{C}^{d}$ be a Kobayashi hyperbolic domain. For all $z,w,p\in D$ define the Gromov product $(z|w)_{p}$ as follows:
$(z|w)_{p}:=\frac{1}{2}[k_{D}(z,p)+k_{D}(w,p)-k_{D}(z,w)].$
Theorem 3.4.
Let $D\subset\mathbb{C}^{d}$ be a bounded convex domain of finite type.
  1. (i)
    (Boundary and Gromov product) If $\xi,\eta$ are two distinct points in $\partial D$, then
    $\limsup_{(z,w)\to(\xi,\eta)}(z|w)_{p}<+\infty.$
  2. (ii)
    (4-point condition) there exists $\delta\geq 0$ such that for all $z_{1},z_{2},z_{3},p\in D$ we have
    $(z_{1}|z_{2})_{p}\geq\min\{(z_{1}|z_{3})_{p},(z_{2}|z_{3})_{p}\}-\delta.$
  3. (iii)
    (Asymptoticity of (1,B) quasi-geodesics) Let $\gamma_{1},\gamma_{2}\colon\mathbb{R}_{\geq 0}\to D$ be two $(1,B)$ quasi-geodesic rays with the same endpoint $\xi\in\partial D$. Then
    $\sup_{t\geq 0}k_{D}(\gamma_{1}(t),\gamma_{2}(t))<\infty.$
Proof.
It is proved in [63] that $(D,k_{D})$ is Gromov hyperbolic and that the identity map $\mathrm{id}_{D}\colon D\to D$ extends to an homeomorphism between the Gromov compactification and $\overline{D}$. Points (i) and (ii) are classical results for proper geodesic Gromov hyperbolic metric spaces, see e.g. [18, 31]. For a proof of (iii), see [12, Lemma 5.8]. ∎
In order to adapt these properties to our setting, we will use the following localization argument.
Remark 3.5.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $L\geq 2$ be given by Definition 2.3. Let $W$ be a neighborhood of $\xi$ such that
  1. (1)
    $\tilde{D}:=D\cap W$ is a bounded convex domain;
  2. (2)
    the boundary $\partial\tilde{D}$ is of class $C^{L}$;
  3. (3)
    every point of $\partial\tilde{D}$ has line type at most $L$.
The domain $\tilde{D}$ enjoys the visibility property (see for instance [21] for the definition of visibility), hence it follows from [21, Theorem 1.4] that for every neighborhood $V\subset\subset W$ of $\xi$ there exists $C\geq 0$ such that
$k_{D}(z,w)\leq k_{\tilde{D}}(z,w)\leq k_{D}(z,w)+C,\quad\forall\,z,w\in V.$ (3.1)
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain. We denote by $\partial^{*}D$ the boundary of $D$ in the one-point compactification of $\mathbb{C}^{d}$.
Proposition 3.6.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain, and let $\xi\in\partial D$ be a point of locally finite type.
  • (i)
    Let $\eta\in\partial^{*}D\setminus\{\xi\}$ and $p\in D$. Then
    $\limsup_{(z,w)\to(\xi,\eta)}(z|w)_{p}<+\infty.$
  • (ii)
    There exist $\delta^{\prime}\geq 0$ and a neighborhood $V$ of $\xi$ such that for all $z_{1},z_{2},z_{3},p\in D\cap V$ we have
    $(z_{1}|z_{2})_{p}\geq\min\{(z_{1}|z_{3})_{p},(z_{2}|z_{3})_{p}\}-\delta^{\prime}.$
  • (iii)
    If $\gamma_{1},\gamma_{2}\colon\mathbb{R}_{\geq 0}\to D$ are two $(1,B)$ quasi-geodesic rays with endpoint $\xi$, then
    $\sup_{t\geq 0}k_{D}(\gamma_{1}(t),\gamma_{2}(t))<+\infty.$
Proof.
Let $W$, $\tilde{D}$ and $C$ be given by Remark 3.5. Let $V\subset\subset W$ be a neighborhood of $\xi$.
(i) Denote by $(\cdot|\cdot)^{D}_{p}$ the Gromov product with respect to $k_{D}$. Notice that we can suppose that $p\in V$, indeed for all $z,w,p,q\in D$, by the triangle inequality we have $|(z|w)^{D}_{p}-(z|w)^{D}_{q}|\leq k_{D}(p,q)$.
[…]
$(z_{n}|w_{n})^{D}_{p}\to+\infty.$
For all $n\geq 0$ let $\gamma_{n}\colon[0,T_{n}]\to D$ be a geodesic segment with $\gamma_{n}(0)=z_{n}$, $\gamma_{n}(T_{n})=w_{n}$. Let $B\subset\subset V$ be a small Euclidean ball centered in $\xi$ such that $\eta\notin B$. Let $N\geq 0$ such that $z_{n}\in B$ for all $n\geq N$. Then for all $n\geq N$ there exists $t_{n}\in(0,T_{n}]$ such that $\gamma_{n}(t_{n})\in D\cap\partial B$. Set $w^{\prime}_{n}:=\gamma_{n}(t_{n})$. We have
$2(z_{n}|w_{n})^{D}_{p}$ $=k_{D}(z_{n},p)+k_{D}(w_{n},p)-k_{D}(z_{n},w_{n})$
$=k_{D}(z_{n},p)+k_{D}(w_{n},p)-k_{D}(z_{n},w^{\prime}_{n})-k_{D}(w^{\prime}_{n},w_{n})$
$\leq k_{D}(z_{n},p)+k_{D}(w^{\prime}_{n},p)-k_{D}(z_{n},w^{\prime}_{n})=2(z_{n}|w^{\prime}_{n})^{D}_{p}.$
Since $p\in V$ and $z_{n},w^{\prime}_{n}\in V$ for all $n\geq N$, by (3.1) we have
$(z_{n}|w_{n}^{\prime})^{\tilde{D}}_{p}\geq(z_{n}|w_{n}^{\prime})^{D}_{p}-\frac{C}{2},$
and thus $(z_{n}|w_{n}^{\prime})^{\tilde{D}}_{p}\to$$\infty$. It easily follows that the sequence $(w_{n}^{\prime})$ is not relatively compact in $\tilde{D}$, and thus, up to a subsequence, we can assume that it converges to a point $\eta^{\prime}\in\partial\tilde{D}\cap\partial B$. This contradicts (i) of Theorem 3.4 because $\eta^{\prime}\neq\xi$.
(ii) By (3.1) and (ii) in Theorem 3.4 there exists $\delta\geq 0$ such that for all $z_{1},z_{2},z_{3},p\in V$
$(z_{1}|z_{2})^{D}_{p}\geq\min\{(z_{1}|z_{3})^{D}_{p},(z_{2}|z_{3})^{D}_{p}\}-\frac{3C}{2}-\delta.$
(iii) Let $T\geq 0$ be such that $\gamma_{1}(t),\gamma_{2}(t)\in V$ for all $t\geq T$. By (3.1) the curves $\gamma_{1}|_{[T,+\infty)}$ and $\gamma_{2}|_{[T,+\infty)}$ are $(1,B+C)$ quasi-geodesics with respect to $k_{\tilde{D}}$, so we conclude with (iii) of Theorem 3.4. ∎
Point (i) of the previous proposition immediately implies the following.
Corollary 3.7.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain, and let $\xi\in\partial D$ be a point of locally finite type. If $(z_{n}),(w_{n})$ are sequences in $D$ such that $z_{n}\to\xi$ and $(k_{D}(z_{n},w_{n}))_{n}$ is bounded, then $w_{n}\to\xi$.
We end this section with an estimate of the Kobayashi distance near the boundary which will be useful later on.
Lemma 3.8.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain, and let $\xi\in\partial D$ be a point of locally finite type. There exists a neighborhood $V$ of $\xi$ such that for all $p\in D$ there exists $c\geq 0$ such that
$\log\delta_{D}(z)-c\leq k_{D}(z,p)\leq-\log\delta_{D}(z)+c,\quad\forall\,z\in D\cap V.$
Proof.
Let $\tilde{D}$ and $V$ be given by Remark 3.5. By [8, Lemma 1.1, Lemma 1.3] given $z_{0}\in V\cap D$ there exists a constant $c^{\prime}\geq 0$ such that
$\log\delta_{\tilde{D}}(z)-c^{\prime}\leq k_{\tilde{D}}(z,z_{0})\leq-\log\delta_{\tilde{D}}(z)+c^{\prime},\quad\forall z\in\tilde{D}$.
Up to taking a smaller neighborhood $V$ of $\xi$ we clearly have $\delta_{D}=\delta_{\tilde{D}}$ on $V\cap D$. We conclude using (3.1). ∎
4. Complex geodesics
Definition 4.1 (Complex geodesics).
Let $D\subset\mathbb{C}^{d}$ be a Kobayashi hyperbolic domain. An extremal map is a holomorphic map $\varphi\colon\mathbb{D}\to D$ such that there exist distinct points $\zeta_{1},\zeta_{2}\in\mathbb{D}$ satisfying
$k_{D}(\varphi(\zeta_{1}),\varphi(\zeta_{2}))=k_{\mathbb{D}}(\zeta_{1},\zeta_{2}).$
A complex geodesic is a holomorphic map $\varphi\colon\mathbb{D}\to D$ such that $k_{D}(\varphi(\zeta_{1}),\varphi(\zeta_{2}))=k_{\mathbb{D}}(\zeta_{1},\zeta_{2})$ for all $\zeta_{1},\zeta_{2}\in\mathbb{D}$. If $\varphi\colon\mathbb{D}\to D$ is a complex geodesic, a holomorphic map $\tilde{\rho}\colon D\to\mathbb{D}$ is called a left inverse of $\varphi$ if $\tilde{\rho}\circ\varphi=\mathrm{id}_{\mathbb{D}}.$ Analogous definitions hold for holomorphic maps defined in the left half-plane $\mathbb{H}$. In particular we say that a holomorphic map $\varphi\colon\mathbb{H}\to D$ is a complex geodesic if $k_{D}(\varphi(\zeta_{1}),\varphi(\zeta_{2}))=k_{\mathbb{H}}(\zeta_{1},\zeta_{2})$ for all $\zeta_{1},\zeta_{2}\in\mathbb{H}$.
Proposition 4.2.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain. Then
  • (i)
    every extremal map $\varphi\colon\mathbb{D}\to D$ is a complex geodesic;
  • (ii)
    for any two distinct points $z,w\in D$ there exists a complex geodesic $\varphi\colon\mathbb{D}\to D$ such that $\varphi(0)=z$ and $\varphi(\tanh(k_{D}(z,w)/2))=w$;
  • (iii)
    every complex geodesic admits a left inverse.
Proof.
Follows from [24] and Proposition 11.1.4, Proposition 11.1.7 and Theorem 11.2.1 in [42]. ∎
The following result was proved by Abate [6] in the case of bounded strongly convex domains with $C^{3}$ boundary. Our proof is based on a Gromov hyperbolicity method.
Proposition 4.3.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $p\in D$. Then there exists a complex geodesic $\varphi\colon\mathbb{D}\to D$ with $\varphi(0)=p$ and $\lim_{\mathbb{R}\ni t\to 1^{-}}\varphi(t)=\xi.$
Proof.
Let $p\in D$. Let $(z_{n})$ be a sequence in $D$ converging to $\xi$. For all $n\geq 0$ let $\varphi_{n}$ be a complex geodesic such that $\varphi_{n}(0)=p$ and $\varphi_{n}(t_{n})=z_{n}$, where $t_{n}:=\tanh(k_{D}(p,z_{n})/2)\to 1.$
Since $D$ is taut, up to a subsequence we can assume that $\varphi_{n}\to\varphi$ uniformly on compact sets. Clearly $\varphi$ is a complex geodesic of $D$. We show that $\lim_{t\to 1^{-}}\varphi(t)=\xi$. Assume by contradiction that this is false, that is, there exists a sequence $(s_{k})$ in $[0,1)$ converging to $1$ such that $\varphi(s_{k})$ converges to a point $\eta\in\partial^{*}D$ different from $\xi$. Since $\varphi_{n}\to\varphi$, it follows that for all $k\geq 0$ we can find $n_{k}$ such that
$\|\varphi(s_{k})-\varphi_{n_{k}}(s_{k})\|<\frac{1}{k},$
in particular $\varphi_{n_{k}}(s_{k})\to\eta$.
Now
$(\varphi_{n_{k}}(t_{n_{k}})|\varphi_{n_{k}}(s_{k}))_{p}=\min\{k_{\mathbb{D}}(0,t_{n_{k}}),k_{\mathbb{D}}(0,s_{k})\}\to+\infty,$
which contradicts (i) of Proposition 3.6.
Definition 4.4.
Denote by $\gamma_{\mathbb{D}}\colon\mathbb{R}_{\geq 0}\to\mathbb{D}$ the geodesic ray in $\mathbb{D}$ with starting point 0 and endpoint $1$, that is $\gamma_{\mathbb{D}}(t)={\rm\tanh}(t/2).$ Denote by $\gamma_{\mathbb{H}}\colon\mathbb{R}_{\geq 0}\to D$ the geodesic ray in $\mathbb{H}$ with starting point 1-1 and endpoint 0, that is $\gamma_{\mathbb{H}}(t)=\mathscr{C}\circ\gamma_{\mathbb{D}}=-e^{-t}.$
Let now $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain. If $\varphi\colon\mathbb{D}\to D$ (resp. $\varphi\colon\mathbb{H}\to D$) is a holomorphic map, we denote by $\tilde{\varphi}\colon\mathbb{R}_{\geq 0}\to D$ the curve $\tilde{\varphi}=\varphi\circ\gamma_{\mathbb{D}}$ (resp. $\varphi\circ\gamma_{\mathbb{H}}$). If the curve $\tilde{\varphi}$ has an endpoint $\xi\in\partial D$, then we say that $\xi$ is the endpoint of $\varphi$. Notice that if $\varphi$ is a complex geodesic, then $\tilde{\varphi}$ is a geodesic ray.
Definition 4.5.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain. Let $\varphi\colon\mathbb{D}\to D$ (resp. $\varphi\colon\mathbb{H}\to D$) be a holomorphic map. We say that the point $1\in\partial\mathbb{D}$ (resp. $0\in\partial\mathbb{H}$) is a regular contact point of $\varphi$ if the curve $\tilde{\varphi}$ is an almost-geodesic ray of $D$ with endpoint of locally finite type $\xi\in\partial D$. By Corollary 3.7 this implies that $\varphi$ has non-tangential limit $\xi$ at $1$ (resp. at 0).
Remark 4.6.
In Section 10 we will show that regular contact points can be defined equivalently as points of finite dilation (see Definition 10.8 and Corollary 10.11). We will also extend the definition of regular contact points to the case of holomorphic maps between two $\mathbb{C}$-proper convex domains.
Remark 4.7.
Clearly if $\varphi\colon\mathbb{D}\to D$ (resp. $\varphi\colon\mathbb{H}\to D$) is a complex geodesic in a $\mathbb{C}$-proper convex domain $D$ with endpoint of locally finite type $\xi\in\partial D$, then 1 (resp. 0) is a regular contact point.
Let $D\subset\mathbb{C}^{d}$ be a bounded strongly convex domain with $C^{3}$ boundary, and let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi\in\partial D$. Then $\varphi$ extends to $\overline{\mathbb{D}}$ as a $C^{1}$ map , and $\varphi^{\prime}(1)$ is transversal to $T_{\xi}D$ (see e.g. [1]). This result plays an important role in Abate’s proof of the Julia–Wolff–Carathéodory theorem in strongly convex domains. If $D$ is a bounded convex domain of finite type the situation is radically different. Indeed, in this case $\|\varphi^{\prime}(t)\|$ may explode and $\varphi(t)$ may converge to $\xi$ tangentially when $\mathbb{R}\ni t\to 1^{-}$, as the following example shows.
Example 4.8 (Complex geodesics of the egg domain).
Given an even integer $m\geq 2$ define the egg domain
$\mathbb{E}_{m}:=\{(z_{0},z_{1})\in\mathbb{C}^{2}:|z_{0}|^{2}+|z_{1}|^{m}<1\}\subset\mathbb{C}^{2}.$ (4.1)
The points of $\partial\mathbb{E}_{m}$ with $z_{1}=0$ have type $m$, and all other points have type 2. For all $a\in\mathbb{C}$ the map $\varphi_{a}\colon\mathbb{D}\to\mathbb{E}_{m}$ defined by
$\varphi_{a}(\zeta):=\left(\frac{\zeta+|a|^{m}}{1+|a|^{m}},a\left(\frac{1-\zeta}{1+|a|^{m}}\right)^{2/m}\right)$ (4.2)
is a complex geodesic with endpoint $(1,0)$, see [41] and [43]. We are using the principal value of the $m$-th root.
However, one can still say something about the derivative of $\varphi$: it is proved in [12, Lemma 5.12] that the normal component of $\varphi^{\prime}(z)$ admits a positive non-tangential limit as $z\to 1$. In the next result, which will be crucial in what follows, we show that the same is true with only a local finite type assumption around $\xi$ and for any holomorphic map $\varphi\colon\mathbb{D}\to D$ with a regular contact point at 1 and endpoint $\xi$.
Proposition 4.9.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ (resp. $\varphi\colon\mathbb{H}\to D$) be a holomorphic map with a regular contact point at 1 (resp. at 0) and endpoint $\xi$. Then the function
$z\mapsto\langle\varphi^{\prime}(z),n_{\xi}\rangle$
admits a positive non-tangential limit as $z\to 1$ (resp. as $z\to 0$), which we denote by $\varphi^{\prime}_{N}(1)$ (resp. $\varphi^{\prime}_{N}(0)$).
Proof.
Assume first that the domain of $\varphi$ is $\mathbb{H}$. Up to a translation we can assume that $\xi=0$. Moreover up to a unitary change of coordinates of $\mathbb{C}^{d}$, we can assume that the outer normal versor $n_{0}$ is $e_{0}$, and thus the real tangent plane of $\partial D$ at 0 is $\{{\rm Re}\,z_{0}=0$}. By convexity of $D$ it follows that $\pi(D)\subseteq\mathbb{H},$ where $\pi\colon\mathbb{C}^{d}\to\mathbb{C}$ denotes the projection to the coordinate $z_{0}$. Write $\varphi(\zeta)=(\varphi_{0}(\zeta),\varphi_{1}(\zeta))\in\mathbb{C}\times\mathbb{C}^{d-1}$ and notice that $\langle\varphi(\zeta),n_{\xi}\rangle=\varphi_{0}(\zeta)$.
By [63, Proposition 11.1] there exists $\varepsilon>0$ such that the inner normal segment at the origin $s\mapsto(-\varepsilon e^{-s},0)$ is a $(1,\log 2)$ quasi-geodesic ray [^4] Actually due to the different normalization of the Kobayashi distance, in [63, Proposition 11.1] it is shown that $s\mapsto(-\varepsilon e^{-2s},0)$ is a $(1,\log\sqrt{2})$ quasi-geodesic with respect to $k_{D}$, so by (iii) in Proposition 3.6 there exists $M\geq 0$ such that, for all $t\geq 0$,
$k_{\mathbb{H}}(\varphi_{0}(-e^{-t}),-e^{-t})$ $\leq k_{\mathbb{H}}(\varphi_{0}(-e^{-t}),-\varepsilon e^{-t})+k_{\mathbb{H}}(-\varepsilon e^{-t},-e^{-t})$
$\leq k_{D}(\tilde{\varphi}(t),(-\varepsilon e^{-t},0))+|\log\varepsilon|\leq M.$
Hence
$\liminf_{t\to 0^{-}}k_{\mathbb{H}}(t,-1)-k_{\mathbb{H}}(\varphi_{0}(t),-1)\leq M.$
Taking into account (1.5), this shows that the map $\mathscr{C}^{-1}\circ\varphi_{0}\circ\mathscr{C}\colon\mathbb{D}\to\mathbb{D}$ has finite dilation at $1$ and has non-tangential limit 1 at 1. By the one-dimensional Julia–Wolff–Carathéodory theorem (Theorem 1.1) it follows that $\varphi_{0}^{\prime}(z)$ has positive non-tangential limit at 0.
Now assume that $\varphi\colon\mathbb{D}\to D$ is a holomorphic map with a regular contact point at $1$ with non-tangential limit $\xi$. Define $\psi\colon\mathbb{H}\to D$ as $\psi=\varphi\circ\mathscr{C}^{-1}$. Since $(\mathscr{C}^{-1})^{\prime}(0)=2$ we have
$\varphi^{\prime}_{N}(1)=\angle\lim_{z\to 1}\langle\varphi^{\prime}(z),n_{\xi}\rangle=\frac{1}{2}\angle\lim_{z\to 0}\langle\psi^{\prime}(z),n_{\xi}\rangle=\frac{1}{2}\psi^{\prime}_{N}(0)>0.$ (4.3)
5. Multitype
We review the flag of complex subspaces of $\mathbb{C}^{d}$ introduced by Yu [61]. Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of finite type $L\geq 2$. Set $m_{\xi}(0):=+\infty$. For all integers $m\geq 1$ define
$S_{m}:=\{v\in\mathbb{C}^{d}:m_{\xi}(v)\geq m\}.$
Yu shows, using the convexity of $D$, that the sets $S_{m}$ are complex linear subspaces. It immediately follows from the definition that $S_{1}=\mathbb{C}^{d}$, $S_{m_{2}}\subseteq S_{m_{1}}$ for $m_{1}\leq m_{2}$ and that $S_{m}=\{0\}$ if $m>L$. Moreover $S_{2}$ coincides with the complex tangent subspace $T_{\xi}^{\mathbb{C}}\partial D$, indeed
$v\in S_{2}\iff\frac{\partial}{\partial\zeta}r(\xi+\zeta v)|_{\zeta=0}=0\iff\sum_{j=0}^{d-1}\frac{\partial r}{\partial z_{j}}(\xi)v_{j}=0.$
Now set $l_{1}=1$. For all $j\geq 1$ define recursively $l_{j}>l_{j-1}$ as the smallest integer such that $S_{l_{j}}\subsetneq S_{l_{j-1}}$. This procedure stops when we find an integer $j$ such that $l_{j}=L$. We denote such integer $k$. We thus obtain the multitype flag of complex subspaces of $\mathbb{C}^{d}$:
$\{0\}\subsetneq S_{L}\subsetneq S_{l_{k-1}}\subsetneq\cdots\subsetneq S_{1}=\mathbb{C}^{d}.$
We say that an orthonormal basis $v_{0},\dots,v_{d-1}$ of $\mathbb{C}^{d}$ is a multitype basis at $\xi$ if $v_{0}$ is the outer normal versor $n_{\xi}$ at $\xi$, and if the basis $v_{0},\dots,v_{d-1}$ is adapted to the multitype flag (and presented in reverse order), that is, for all $j=0,\dots,k-1$ the last ${\rm dim}\,S_{l_{k-j}}$ vectors are a basis of the subspace $S_{l_{k-j}}$. For all $j=1,\dots,d-1$ set $m_{j}:=m_{\xi}(v_{j})$.
Following Yu, we call the vector $(m_{0},m_{1},\cdots,m_{d-1})\in\mathbb{N}^{d}$ the (linear) multitype of $\partial D$ at $\xi$. If $\partial D$ is smoothly bounded, then it is proved in [61] that $(m_{0},m_{1},\cdots,m_{d-1})$ coincides with the multitype in the sense Catlin (see [27]). Notice that $m_{0}=1$ and $2\leq m_{j}\leq L$ if $j=1,\dots,d-1$.
By an affine unitary change of coordinates we can assume that $\xi=0$ and that $v_{j}=e_{j}$ for $j=0,\dots,d-1$. We call the new coordinates multitype coordinates. In multitype coordinates the outer normal versor at the origin is $n_{0}=e_{0}$ and the hypersurface $\partial D$ has the following defining function in a neighborhood of the origin (see e.g. [36])
$r(z)={\operatorname{Re}\,}z_{0}+H(z_{1},\cdots,z_{d-1})+R(z),$ (5.1)
where $H\colon\mathbb{C}^{d-1}\longrightarrow\mathbb{R}$ is a convex non-negative polynomial which is
  1. (1)
    non-degenerate, that is the set $\{H=0\}$ does not contain any complex line;
  2. (2)
    weighted homogeneous with respect to $(m_{1},\dots,m_{d-1})$, i.e. for all $t>0$ and $(z_{1},\cdots,z_{d-1})\in\mathbb{C}^{d-1}$ we have
    $H\left(t^{1/\penalty 50m_{1}}z_{1},\dots,t^{1/\penalty 50m_{d-1}}z_{d-1}\right)=tH(z_{1},\cdots,z_{d-1}).$
The remainder satisfies
$R(z)=o\bigg(|z_{0}|+\sum_{j=1}^{d-1}|z_{j}|^{m_{j}}\bigg).$
We end this section introducing a dual version of the type of a vector at $\xi$. Denote by $(\mathbb{C}^{d})^{*}$ the dual vector space of $\mathbb{C}^{d}$.
Definition 5.1.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of finite type $L\geq 2$. If $\theta\in(\mathbb{C}^{d})^{*}\setminus\{0\}$, then we define its cotype $M_{\xi}(\theta)$ at $\xi$ as the integer $1\leq l_{j_{0}}\leq L$, where
$j_{0}:=\max_{1\leq j\leq k}\{j:\theta\not\in{\rm Ann}(S_{j})\},$
and ${\rm Ann}(S_{j})$ denotes the annihilator of $S_{j}$ in $(\mathbb{C}^{d})^{*}$. Notice that, since $S_{j}$ is a decreasing sequence of subspaces, if $\theta\in{\rm Ann}(S_{j})$, then $\theta\in{\rm Ann}(S_{i})$ for all $i\geq j$. If $v\in\mathbb{C}^{d}\setminus\{0\}$ we define the cotype $M_{\xi}(v)$ of $v$ as the cotype of the linear functional $w\mapsto\langle w,v\rangle$.
Remark 5.2.
If $v_{0},\dots,v_{d-1}$ is a multitype basis at $\xi$ and $v=\sum a_{j}v_{j}\in\mathbb{C}^{d}\setminus\{0\}$, then it is easy to see that $M_{\xi}(v)=\max\{m_{j}:a_{j}\neq 0\}$. Notice that $m_{\xi}(v)=\min\{m_{j}:a_{j}\neq 0\}$, so in particular we have $M_{\xi}(v_{j})=m_{\xi}(v_{j})$ for all $j=0,\dots,d-1$.
6. Scaling
In this section we recall Gaussier’s scaling in the normal direction and we use it to prove several properties of complex geodesics with endpoint of locally finite type.
Definition 6.1.
If $H\colon\mathbb{C}^{d-1}\to\mathbb{R}$ is a convex non-negative non-degenerate weighted homogeneous polynomial we call the domain
$D_{H}:=\{(z_{0},w)\in\mathbb{C}\times\mathbb{C}^{d-1}:{\operatorname{Re}\,}z_{0}+H(w)<0\},\quad$
a scaling model. If $D$ is a $\mathbb{C}$-proper convex domain with a point of locally finite type at the origin in multitype coordinates, then we say that the domain $D_{H}$, where $H$ is the polynomial given by (5.1), is the scaling model of $D$.
We recall that the Hausdorff distance between two compact sets $X,Y\subset\mathbb{C}^{d}$ is given by
$d_{H}(X,Y):=\max\left\{\sup_{x\in X}\inf_{y\in Y}\|x-y\|,\sup_{y\in Y}\inf_{x\in X}\|x-y\|\right\}.$
Moreover, a sequence of convex domains $(D_{n})$ in $\mathbb{C}^{d}$ converges in the local Hausdorff topology to a convex domain $D_{\infty}\subset\mathbb{C}^{d}$ if, for all $r>0$,
$\lim_{n\to+\infty}d_{H}(\overline{D_{n}}\cap\overline{r\mathbb{B}^{d}},\overline{D_{\infty}}\cap\overline{r\mathbb{B}^{d}})=0.$
The following theorem is proved in [36].
Theorem 6.2.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain with a point of locally finite type at the origin in multitype coordinates. Let $(\lambda_{n})$ be a sequence in $\mathbb{R}_{>0}$ converging to $\infty$ and for all $n\geq 0$ define the linear map $A_{n}\colon\mathbb{C}^{d}\to\mathbb{C}^{d}$ by
$A_{n}(z)=(\lambda_{n}z_{0},\lambda_{n}^{1/m_{1}}z_{1},\cdots,\lambda_{n}^{1/m_{d-1}}z_{d-1}).$
Then the domain $A_{n}D$ converges in the local Hausdorff topology to the scaling model $D_{H}$ of $D$.
Points (i) and (ii) of the following corollary are proved in [63, Lemma 4.4] and [38, Theorem 9.1] respectively. For lack of a reference of point (iii) we give a short proof.
Corollary 6.3.
In the assumptions of the previous theorem,
  • (i)
    every compact subset $K\subset D_{H}$ is eventually contained in the domain $A_{n}D$;
  • (ii)
    we have $\lim_{n\to+\infty}\kappa_{A_{n}D}=\kappa_{D_{H}}$ and $\lim_{n\to+\infty}k_{A_{n}D}=k_{D_{H}}$ uniformly on compact sets of $D_{H}\times\mathbb{C}^{d}$ and $D_{H}\times D_{H}$, respectively;
  • (iii)
    if $0<r<R$ and $z\in D_{H}$, then eventually $B_{A_{n}D}(z,r)\subseteq B_{D_{H}}(z,R).$
Proof.
Let $0<r<R$ and let $z\in D_{H}$. Notice that by (i) the point $z$ is eventually contained in $A_{n}D$. Fix $0<\delta<1$ such that $k_{\mathbb{D}}(0,\delta)>r$. Then by [63, Lemma 4.6] there exists $N_{1}\geq 0$ such that for all $n\geq N_{1}$, if $\varphi\colon\mathbb{D}\to A_{n}D$ is holomorphic with $\varphi(0)=z$, then $\varphi(\delta\mathbb{D})\subset D_{H}$.
We claim that there exists $C>0$ such that for all $n\geq N_{1}$ we have $\overline{B_{A_{n}D}(z,r)}\subseteq\overline{B_{D_{H}}(z,C)}$. Indeed, let $n\geq N_{1}$, let $w\in\overline{B_{A_{n}D}(z,r)}$ and consider a complex geodesic $\varphi\colon\mathbb{D}\to A_{n}D$ such that $\varphi(0)=z$ and $\varphi(t)=w$ with $t:=\tanh(k_{D}(z,w)/2)$. The map $\varphi_{\delta}:\mathbb{D}\to\mathbb{C}^{d}$ given by $\varphi_{\delta}(\zeta):=\varphi(\delta\zeta)$ has image in $D_{H}$. Notice that $t\leq\tanh(r/2)<\delta$, so
$k_{D_{H}}(z,w)=k_{D_{H}}(\varphi_{\delta}(0),\varphi_{\delta}(t/\delta))\leq k_{\mathbb{D}}(0,t/\delta)\leq k_{\mathbb{D}}(0,\tanh(r/2)/\penalty 50\delta)=:C,$
which proves the claim. The result now follows using (ii) on the compact subset $K:=\overline{B_{D_{H}}(z,C)}$. ∎
Notice that in multitype coordinates the inner normal segment $\sigma\colon[t_{0},1)\to\mathbb{C}^{d}$ with endpoint $\xi=0$ is given by $\sigma(t)=(t-1,0)\in\mathbb{C}\times\mathbb{C}^{d-1}.$
Corollary 6.4.
In the assumptions of the previous theorem, if $(z_{n})$ is a sequence in $D$ converging to 0 and there exists $M\geq 0$ such that $k_{D}(z_{n},\sigma)\leq M$ for all $n\geq 0$, then there exist $\lambda_{n}\to+\infty$ and $N\geq 0$ such that the sequence $(A_{n}z_{n})_{n\geq N}$ is relatively compact in $D_{H}$.
Proof.
Let $(a_{n})$ be a sequence of points in $\mathbb{R}_{>0}$ such that $k_{D}(z_{n},(-a_{n},0))\leq M$ for all $n\geq 0$. Clearly $a_{n}\to 0$. Set $\lambda_{n}:=\frac{1}{a_{n}}.$ Then for all $n\geq 0$ we have $A_{n}(-a_{n},0)=(-1,0)$ and $k_{A_{n}D}(A_{n}z_{n},(-1,0))\leq M$. By (iii) of Corollary 6.3, given $M^{\prime}>M$ there exists $N\geq 0$ such that for all $n\geq N$ we have
$A_{n}z_{n}\in B_{A_{n}D}((-1,0),M)\subset B_{D_{H}}((-1,0),M^{\prime})\subset\subset D_{H}.$
Definition 6.5.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi$ be a point of locally finite type in $\partial D$. Two holomorphic maps $\varphi,\psi\colon\mathbb{D}\to D$ with a regular contact point at 1 and with endpoint $\xi$ are strongly asymptotic if there exists $T\in\mathbb{R}$ such that
$\lim_{t\to+\infty}k_{D}(\tilde{\varphi}(t),\tilde{\psi}(t+T))=0.$
In the next proposition we prove that if we rescale any holomorphic map $\varphi:\mathbb{H}\to D$ with a regular contact point at 0 and endpoint at the origin we obtain in the limit the “slice” complex geodesic $\zeta\mapsto(\varphi^{\prime}_{N}(0)\zeta,0)$ of the model domain $D_{H}$. This considerably generalizes [12, Lemma 5.13], where $D$ is assumed to be a bounded convex domain of finite type and where $\varphi$ is assumed to be a complex geodesic.
Proposition 6.6.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain with a point of locally finite type at the origin in multitype coordinates. Let $D_{H}$ be the scaling model of $D$. Let $\varphi\colon\mathbb{H}\to D$ be a holomorphic map with a regular contact point at 0 and endpoint the origin. Let $(\lambda_{n})$ be a sequence in $\mathbb{R}_{>0}$ converging to $\infty$. Then the sequence $(A_{n}\varphi(\lambda_{n}^{-1}\cdot))$ converges uniformly on compact subsets to the complex geodesic $\hat{\varphi}\colon\mathbb{H}\to D_{H}$ defined by
$\hat{\varphi}(\zeta)=(\varphi^{\prime}_{N}(0)\zeta,0).$
Proof.
By the proof of [12, Lemma 5.13] the sequence $(A_{n}\varphi(\lambda_{n}^{-1}\cdot))$ converges (up to a subsequence) uniformly on compact subsets to a holomorphic map $\hat{\varphi}\colon\mathbb{D}\to D_{H}$ such that if we write
$\hat{\varphi}(\zeta)=(\hat{\varphi}_{0}(\zeta),\hat{\varphi}_{1}(\zeta))\in\mathbb{C}\times\mathbb{C}^{d-1},$
we have $\hat{\varphi}_{0}(\zeta)=\varphi^{\prime}_{N}(0)\zeta.$ Fix $s,t<0$. We claim that $k_{D_{H}}(\hat{\varphi}(s),\hat{\varphi}(t)))=k_{\mathbb{H}}(s,t).$ Indeed
$k_{D_{H}}(\hat{\varphi}(s),\hat{\varphi}(t))=\lim_{n\to+\infty}k_{A_{n}D}(A_{n}\varphi(\lambda_{n}^{-1}s),A_{n}\varphi(\lambda_{n}^{-1}t))\\ =\lim_{n\to+\infty}k_{D}(\varphi(\lambda_{n}^{-1}s),\varphi(\lambda_{n}^{-1}t)).$
Clearly for all $n\geq 0$ we have $k_{D}(\varphi(\lambda_{n}^{-1}s),\varphi(\lambda_{n}^{-1}t))\leq k_{\mathbb{H}}(\lambda_{n}^{-1}s,\lambda_{n}^{-1}t)=k_{\mathbb{H}}(s,t)$. On the other hand, fix $\varepsilon>0$. If $n$ is large enough, then $\lambda_{n}^{-1}s$ and $\lambda_{n}^{-1}t$ are >1,>-1, and thus
$k_{D}(\varphi(\lambda_{n}^{-1}s),\varphi(\lambda_{n}^{-1}t))=k_{D}(\tilde{\varphi}(-\log(-\lambda_{n}^{-1}s)),\tilde{\varphi}(-\log(-\lambda_{n}^{-1}t))).$
Since $\tilde{\varphi}$ is an almost geodesic, if $n$ is large enough we have
$k_{D}(\tilde{\varphi}(-\log(-\lambda_{n}^{-1}s)),\tilde{\varphi}(-\log(-\lambda_{n}^{-1}t)))\geq\left|\log\frac{s}{t}\right|-\varepsilon=k_{\mathbb{H}}(s,t)-\varepsilon,$
and thus $k_{D_{H}}(\hat{\varphi}(s),\hat{\varphi}(t)))\geq k_{\mathbb{H}}(s,t)-\varepsilon$, and the claim is proved. It follows that $\hat{\varphi}\colon\mathbb{H}\to D_{H}$ is an extremal map and thus by Proposition 4.2 the map $\hat{\varphi}$ is a complex geodesic. It now follows from the proof of [12, Lemma 5.13] that $\hat{\varphi}_{1}\equiv 0$, and thus $\hat{\varphi}(\zeta)=(\varphi^{\prime}_{N}(0)\zeta,0).$
Corollary 6.7.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi,\psi\colon\mathbb{D}\to D$ be two holomorphic maps with a regular contact point at 1 and endpoint $\xi$, and satisfying $\varphi^{\prime}_{N}(1)=\psi^{\prime}_{N}(1)$. Then
$\angle\lim_{z\to 1}k_{D}(\varphi(z),\psi(z))=0.$
Proof.
Consider multitype coordinates. We work with $\mathbb{H}$ instead of $\mathbb{D}$, the result will follow by composing with the Cayley transform. Let $\varphi,\psi\colon\mathbb{H}\to D$ be two holomorphic maps with a regular contact point at 0 and endpoint $\xi$, and satisfying $\varphi^{\prime}_{N}(0)=\psi^{\prime}_{N}(0)=:c$. By contradiction, suppose that there exists a sequence $(\zeta_{n})$ in $\mathbb{H}$ converging non-tangentially to 0 such that
$\lim_{n\to+\infty}k_{D}(\varphi(\zeta_{n}),\psi(\zeta_{n}))=M>0.$
For all $n\geq 0$ set $\lambda_{n}:=|{\operatorname{Re}\,}\zeta_{n}|^{-1}$ and define the maps $\varphi_{n}(z):=A_{n}\varphi(\lambda_{n}^{-1}z)$ and $\psi_{n}(z):=A_{n}\psi(\lambda_{n}^{-1}z)$. By Proposition 6.6 the two sequences converge uniformly on compact subsets to
$\hat{\varphi}(\zeta)=(c\zeta,0)=\hat{\psi}(\zeta).$
[…]
$0<M=\lim_{n\to+\infty}k_{D}(\varphi(\zeta_{n}),\psi(\zeta_{n}))=\lim_{n\to+\infty}k_{A_{n}D}(A_{n}\varphi(\zeta_{n}),A_{n}\psi(\zeta_{n}))=k_{D_{H}}(\hat{\varphi}(\hat{\zeta}),\hat{\psi}(\hat{\zeta}))=0,$
which gives a contradiction. ∎
Corollary 6.8.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi,\psi\colon\mathbb{D}\to D$ be two holomorphic maps with a regular contact point at 1 and with endpoint $\xi$. Then $\varphi$ and $\psi$ are strongly asymptotic.
Proof.
Let $\lambda:=\varphi^{\prime}_{N}(0)/\psi^{\prime}_{N}(0)$, and define a holomorphic map $\theta\colon\mathbb{H}\to D$ by $\theta(z)=\psi(\lambda z)$. Then
$\tilde{\theta}(t)=\psi(-e^{\log\lambda-t})=\tilde{\psi}(t-\log\lambda),$
and $\theta^{\prime}_{N}(0)=\varphi^{\prime}_{N}(0)$, hence the result follow from Corollary 6.7. ∎
As another application of the scaling method we prove the following lemma, which will be an important tool in the proof of our main Theorem 1.3. The result is interesting already in the particular case of a complex geodesic $\varphi\colon\mathbb{D}\to D$.
Lemma 6.9.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ be a holomorphic map with a regular contact point at 1 and with endpoint $\xi$, let $n_{\xi}$ denote the outer normal versor in $\xi$. Then if we parametrize the inner normal segment at $\xi$ as
$\sigma(t)=\xi+(t-1)\varphi_{N}^{\prime}(1)n_{\xi},\quad t\in[t_{0},1),$
we have
$\lim_{t\to 1^{-}}k_{D}(\varphi(t),\sigma(t))=0.$
Proof.
Consider multitype coordinates. Then $\sigma(t)=(\varphi^{\prime}_{N}(1)(t-1),0)$. Set $\psi:=\varphi\circ\mathscr{C}^{-1}$, $\theta(t):=\sigma\circ\mathscr{C}^{-1}(t)=\left(\frac{2t}{1-t}\varphi^{\prime}_{N}(1),0\right)$. We will show that
$\lim_{t\to 0^{-}}k_{D}(\psi(t),\theta(t))=0.$
Assume by contradiction that this is not the case. Then there exist a sequence $(t_{n})$ in $\mathbb{R}_{<0}$ converging to 0 and $M>0$ such that
$\lim_{n\to+\infty}k_{D}(\psi(t_{n}),\theta(t_{n}))=M>0.$
Set $\lambda_{n}:=|t_{n}|^{-1}$. By Proposition 6.6 the rescaled sequence $(A_{n}\psi(\lambda^{-1}_{n}z))$ converges uniformly on compact subsets to the map $\zeta\mapsto(2\varphi_{N}^{\prime}(1)\zeta,0)$ (since by (4.3) $\psi^{\prime}_{N}(0)=2\varphi_{N}^{\prime}(1)$). On the other hand, the rescaled sequence $(A_{n}\theta(\lambda^{-1}_{n}t))$ converges uniformly on compact subsets to the map $t\mapsto(2\varphi_{N}^{\prime}(1)t,0)$, so with an argument similar to the one in Corollary 6.7 we obtain a contradiction. ∎
The strong asymptoticity of complex geodesics yields the existence of horospheres centered at points of locally finite type of $\partial D$. The proof is similar to the proof of [12, Theorem 3.5], but due to its importance in this paper we include a shortened version for the convenience of the reader.
Theorem 6.10 (Existence of horospheres).
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Fix a base-point $p\in D$. Then as $w\to\xi$ the function $k_{D}(\cdot,w)-k_{D}(w,p)$ converges uniformly on compact subsets of $D$ to a function $h_{\xi,p}$.
Proof.
Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. Then the family of functions
$(k_{D}(\cdot,\varphi(t))-k_{D}(\varphi(t),\varphi(0)))_{t\in[0,1)}=(k_{D}(\cdot,\varphi(t))-t)_{t\in[0,1)}$
is pointwise non-increasing, locally uniformly bounded family of 1-Lipschitz functions. Hence as $t\to 1^{-}$ it converges to a function which can be easily seen to be $1$-Lipschitz, and thus by Dini’s theorem the convergence is uniform on compact subsets. It follows that for every $p\in D$ the family
$(k_{D}(\cdot,\varphi(t))-k_{D}(\varphi(t),p))_{t\in[0,1)}$
converges uniformly on compact subsets to a function $B_{\varphi}(z,p)$. By Corollary 6.8 two complex geodesics of $D$ with endpoint $\xi$ are strongly asymptotic. It easily follows (see [12, Proposition 3.3]), that the function $B_{\varphi}(\,\cdot\,,p)$ does not depend on the choice of $\varphi$.
We show that for any sequence $w_{n}\to\xi$ in $D$ we have
$k_{D}(\,\cdot\,,w_{n})-k_{D}(w_{n},p)\stackrel{{\scriptstyle n\to+\infty}}{{\longrightarrow}}B_{\varphi}(\,\cdot\,,p),$
uniformly on compact subsets of $D$. By the Ascoli–Arzelà Theorem it is enough to prove that for all $z\in D$, every convergent subsequence $\left(k_{D}(z,w_{n_{k}})-k_{D}(w_{n_{k}},p)\right)$ of the sequence $(k_{D}(z,w_{n})-k_{D}(w_{n},p))$ converges to $B_{\varphi}(z,p)$. Let $(\varphi_{k}\colon\mathbb{D}\to D)$ be a sequence of complex geodesics with $\varphi(0)=z$ and $\varphi(r_{k})=w_{n_{k}}$ with $r_{k}\in[0,1)$. Up to extracting another subsequence we can assume that $(\varphi_{k})$ converges uniformly on compact subsets to a complex geodesic $\varphi$. Arguing as in Proposition 4.3 we see that $\varphi$ has endpoint $\xi$. Similarly we obtain a sequence $(\psi_{k}\colon\mathbb{D}\to D)$ of complex geodesics with $\psi_{k}(0)=p$ and $\psi_{k}(s_{k})=w_{n_{k}}$ with $s_{k}\in[0,1)$, converging uniformly on compact subsets to a complex geodesic $\psi$ with endpoint $\xi$. For $a\in(-1,1)$ define the automorphism of the disc $\tau_{a}(z)=\frac{z-a}{1-az}$. Since $\tau^{\prime}(1)=\frac{1+a}{1-a}$, there exists $a\in(-1,1)$ such that $\tau_{a}^{\prime}(1)=\varphi_{N}^{\prime}(1)/\psi^{\prime}_{N}(1).$ Define $\Psi=\psi\circ\tau_{a}$ and $\Psi_{k}=\psi_{k}\circ\tau_{a}$ for all $k$. Then $\Psi^{\prime}_{N}(1)=\varphi^{\prime}_{N}(1)$ and thus by Corollary 6.7 we have
$\lim_{t\to 1^{-}}k_{D}(\varphi(t),\Psi(t))=0.$
Fix $|a|\leq t<1$. For $k$ large enough we have
$k_{D}(z,w_{n_{k}})$ $=k_{D}(z,\varphi_{k}(t))+k_{D}(\varphi_{k}(t),w_{n_{k}}),$
$k_{D}(w_{n_{k}},p)$ $\leq k_{D}(w_{n_{k}},\varphi_{k}(t))+k_{D}(\varphi_{k}(t),p),$
$k_{D}(w_{n_{k}},p)$ $=k_{D}(w_{n_{k}},\Psi_{k}(t))+k_{D}(\Psi_{k}(t),p)$
$\geq k_{D}(w_{n_{k}},\varphi_{k}(t))+k_{D}(\varphi_{k}(t),p)-2k_{D}(\varphi_{k}(t),\Psi_{k}(t)),$
and thus
$k_{D}(z,w_{n_{k}})-k_{D}(w_{n_{k}},p)$ $\geq k_{D}(z,\varphi_{k}(t))-k_{D}(\varphi_{k}(t),p),$
$k_{D}(z,w_{n_{k}})-k_{D}(w_{n_{k}},p)$ $\leq k_{D}(z,\varphi_{k}(t))-k_{D}(\varphi_{k}(t),p)+2k_{D}(\varphi_{k}(t),\Psi_{k}(t)).$
Taking the limit as $k\to\infty$ we obtain
$k_{D}(z,\varphi(t))-k_{D}(\varphi(t),p)\leq\lim_{k\to\infty}k_{D}(z,w_{n_{k}})-k_{D}(w_{n_{k}},p)\leq k_{D}(z,\varphi(t))-k_{D}(\varphi(t),p)+2k_{D}(\varphi(t),\Psi(t)),$
and by letting $t\to 1^{-}$ it follows $\lim_{k\to\infty}k_{D}(z,w_{n_{k}})-k_{D}(w_{n_{k}},p)=B_{\varphi}(z,p)$, as desired.
Definition 6.11.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. The horosphere centered at $\xi$ of radius $R>0$ and with base-point $p\in D$ is the level set
$E_{p}(\xi,R)=\{z\in D\colon h_{\xi,p}(z)<\log R\}.$
Remark 6.12.
Notice that if $p,q\in D$, then
$h_{\xi,q}=h_{\xi,p}+h_{\xi,q}(p),$
hence changing the base-point leaves the family of horospheres centered in $\xi$ invariant, and amounts to multiplying the radius by a fixed constant.
Proposition 6.13.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Then, for all $R>0$,
$\overline{E_{p}(\xi,R)}^{*}\cap\partial^{*}D=\{\xi\},$ (6.1)
where $\overline{E_{p}(\xi,R)}^{*}$ denotes the closure of $E_{p}(\xi,R)$ in the one-point compactification of $\mathbb{C}^{d}$. Moreover,
$\bigcap_{R>0}\overline{E_{p}(\xi,R)}^{*}=\{\xi\}.$ (6.2)
Proof.
The proof is similar to [12, Proposition 6.5], but we include it here for completeness. Fix $R>0$ and $p\in D$. Let $\gamma\colon\mathbb{R}_{\geq 0}\to D$ be a geodesic ray with $\gamma(0)=p$ and endpoint $\xi$. Then for all $t\geq 0$
$h_{\xi,p}(\gamma(t))=-t,$
[…]
$(z_{n}|w_{m})_{p}:=\frac{1}{2}[k_{D}(z_{n},p)+k_{D}(w_{m},p)-k_{D}(z_{n},w_{m})]\leq M.$
Now for all $n\geq N$
$h_{\xi,p}(z_{n}):=\lim_{m\to\infty}k_{D}(z_{n},w_{m})-k_{D}(w_{m},p)\geq-2M+k_{D}(z_{n},p)$
so $\lim_{n\to+\infty}h_{\xi,p}(z_{n})=+\infty$. This proves (6.1), and (6.2) easily follows. ∎
7. $K$-convergence and $K^{\prime}$-convergence
In this section we discuss $K$-convergence and $K^{\prime}$-convergence (or restricted convergence) to a boundary point of locally finite type of a $\mathbb{C}$-proper convex domain. We first deal with $K$-convergence. Given a boundary point $\xi$ of a bounded strongly convex domain with $C^{3}$ boundary $D$, Abate (see e.g. [1]) introduced a generalization of the classical Korányi regions in the ball, namely the $K$-region with vertex $\xi$, base-point $p\in D$ and amplitude $M>1$ as the set
$K_{p}(\xi,M):=\{z\in D\colon h_{\xi,p}(z)+k_{D}(z,p)<2\log M\}.$
Notice that $h_{\xi,p}(z)+k_{D}(p,z)\geq 0$. $K$-regions are comparable to Korányi–Stein admissible regions [60, 48] and to Krantz admissible regions [50]. Thanks to the existence of horospheres (Theorem 6.10) this definition carries on verbatim to the case where $D$ is a $\mathbb{C}$-proper convex domain and $\xi\in\partial D$ is a point of locally finite type, and allows us to define $K$-convergence and $K$-limits in such domains.
Definition 7.1.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $(z_{n})$ be a sequence in $D$ converging to $\xi$. We say that $(z_{n})$ $K$-converges to $\xi$ if it is contained in a $K$-region with vertex $\xi$. Let $f\colon D\to\mathbb{C}$ be a function. We say that $f$ has $K$-limit $L\in\mathbb{C}$ at $\xi$ if for every sequence $(z_{n})$ $K$-converging to $\xi$ we have $f(z_{n})\to L$. In this case we write
$K\textrm{-}\lim_{z\to\xi}f(z)=L.$
Finally we say that $f$ is $K$-bounded at $\xi$ if it is bounded in every $K$-region with vertex $\xi$.
Notice that the previous definitions do not depend on the chosen base-point. Indeed by Remark 6.12 for all $p,q\in D$
$K_{q}(\xi,M)\subseteq K_{p}\left(\xi,Me^{\frac{h_{\xi,p}(q)+k_{D}(q,p)}{2}}\right).$ (7.1)
The notion of geodesic region was introduced in [9] (for bounded strongly convex domains with $C^{3}$ boundary) as regions which are comparable to $K$-regions and thus define the same notions of convergence, limit and boundedness, but are often much easier to work with. See also [12, 13], where such notion is used to define a generalization of the $K$-limit to the context of proper geodesic Gromov hyperbolic metric spaces.
Definition 7.2 (Geodesic region).
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\gamma\colon\mathbb{R}_{\geq 0}\to D$ be a geodesic ray with endpoint $\xi$, and let $R>0$. We define the geodesic region $A(\gamma,R)$ as the open connected subset defined by
$A(\gamma,R)=\left\{z\in D:\inf_{t\geq 0}k_{D}(z,\gamma(t))<R\right\}.$
We say that the point $\xi$ is the vertex of $A(\gamma,R)$. If $\varphi$ is a complex geodesic with endpoint $\xi$, by a slight abuse of notation we denote by $A(\varphi,R)$ the geodesic region $A(\tilde{\varphi},R),$ where $\tilde{\varphi}\colon\mathbb{R}_{\geq 0}\to D$ is the geodesic ray $\tilde{\varphi}(t)=\varphi({\rm tanh}(t/2)).$
Lemma 7.3 (Properties of geodesic regions).
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\gamma\colon\mathbb{R}_{\geq 0}\to D$ be a geodesic ray with endpoint $\xi$.
  • (i)
    If $\theta\colon\mathbb{R}_{\geq 0}\to D$ is another geodesic ray with endpoint $\xi$, then there exists $S>0$ such that, for all $R>0$,
    $A(\gamma,R)\subseteq A(\theta,R+S),\quad A(\theta,R)\subseteq A(\psi,R+S).$
  • (ii)
    For all $R>0$ we have that $\overline{A(\gamma,R)}^{*}\cap\partial^{*}D=\{\xi\}$, where $\overline{A(\gamma,R)}^{*}$ denotes the closure of $A(\gamma,R)$ in the one-point compactification of $\mathbb{C}^{d}$.
  • (iii)
    Let $R>0$, then for all $z\in A(\gamma,R)$
    $h_{\xi,\gamma(0)}(z)<2R-k_{D}(z,\gamma(0));$
    in particular a sequence converging to $\xi$ inside a geodesic region is eventually contained in every horosphere centered at $\xi$.
Proof.
By (iii) of Proposition 3.6 there exists $S>0$ such that $k_{D}(\gamma(t),\theta(t))\leq S$ for each $t\geq 0$, which implies (i). Point (ii) immediately follows from Corollary 3.7. We now prove (iii). Let $z\in A(\gamma,R)$ and consider $t^{*}\geq 0$ such that $k_{D}(z,\gamma(t^{*}))<R$, then
$h_{\xi,\gamma(0)}(z)$ $=\lim_{t\to 1^{-}}k_{D}(z,\gamma(t))-k_{D}(\gamma(t),\gamma(0))$
$\leq\limsup_{t\to 1^{-}}k_{D}(z,\gamma(t^{*}))+k_{D}(\gamma(t^{*}),\gamma(t))-k_{D}(\gamma(t),\gamma(0))$
$=k_{D}(z,\gamma(t^{*}))-k_{D}(\gamma(t^{*}),\gamma(0))$
$\leq 2k_{D}(z,\gamma(t^{*}))-k_{D}(z,\gamma(0))$
$<2R-k_{D}(z,\gamma(0)).$
We now show that geodesic regions and $K$-regions are comparable. The proof of [9, Lemma 7.8] does not work in this context since it is based on the Gromov hyperbolicity of the domain $D$.
Proposition 7.4.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Then every geodesic region with vertex $\xi$ is contained in a $K$-region with vertex $\xi$, and vice versa.
Proof.
Let $\gamma\colon\mathbb{R}_{\geq 0}\to D$ be a geodesic ray with endpoint $\xi$ and set $p=\gamma(0)$. Let $z\in A(\gamma,R)$, then by (iii) in Lemma 7.3 we have
$h_{\xi,p}(z)+k_{D}(z,p)<2R,$
which implies $A(\gamma,R)\subseteq K_{p}(\xi,e^{R})$.
We now prove the converse.
Let $\delta^{\prime}>0$ and $V$ be given by (ii) of Proposition 3.6. Let $\gamma\colon\mathbb{R}_{\geq 0}\to D$ be a geodesic ray with endpoint $\xi$ that is contained in $V$ and set $q:=\gamma(0)$. We will show that every $K$-region $K_{q}(\xi,M)$ is contained in a geodesic region with vertex $\xi$. By (7.1) this will prove the result. Notice that $K_{q}(\xi,M)\cap V^{\complement}$ is relatively compact in $D$, so we can find $R>0$ such that $K_{q}(\xi,M)\cap V^{\complement}\subseteq A(\gamma,R)$. Now assume that $z\in K_{q}(\xi,M)\cap V$, and consider $z_{\gamma}:=\gamma(k_{D}(z,q))$. Let $t>k_{D}(z,q)$, then by (ii) in Proposition 3.6
$(z|z_{\gamma})_{q}\geq\min\{(z|\gamma(t))_{q},(z_{\gamma}|\gamma(t))_{q}\}-\delta^{\prime}.$ (7.2)
Now
$(z|\gamma(t))_{q}=\frac{1}{2}[k_{D}(z,q)+k_{D}(\gamma(t),q)-k_{D}(z,\gamma(t))]\stackrel{{\scriptstyle t\to+\infty}}{{\longrightarrow}}\frac{1}{2}[k_{D}(z,q)-h_{\xi,q}(z)]$
and
$(z_{\gamma}|\gamma(t))_{q}=\frac{1}{2}[k_{D}(z_{\gamma},q)+k_{D}(\gamma(t),q)-k_{D}(z_{\gamma},\gamma(t))]=k_{D}(z_{\gamma},q)=k_{D}(z,q).$
Expanding the Gromov product in left-hand side of (7.2) we have
$2k_{D}(z,q)-k_{D}(z,z_{\gamma})\geq\min\{k_{D}(z,q)-h_{\xi,q}(z),2k_{D}(z,q)\}-2\delta^{\prime},$
which implies
$k_{D}(z,\gamma)\leq k_{D}(z,z_{\gamma})\leq\max\{h_{\xi,q}(z)+k_{D}(z,q),0\}+2\delta^{\prime}\leq 2\log M+2\delta^{\prime},$
so $K_{q}(\xi,M)\subseteq A(\gamma,R^{\prime})$, where $R^{\prime}:=\max\{R,2\log M+2\delta^{\prime}\}.$
There are many instances in which $K$-convergence is the natural generalization to several complex variables of non-tangential convergence in the disc (see e.g. the Julia Lemma in Section 10). There is however a notable exception, namely the Lindelöf principle. Indeed, one can find a bounded holomorphic function defined on the ball $\mathbb{B}^{d}$ which admits radial limit at $e_{0}$ but does not admit $K$-limit at $e_{0}$, see e.g. (1.2). This is the main reason to consider a second, more restrictive definition of convergence to a boundary point $\xi$, for which the Lindelöf principle holds: $K^{\prime}$-convergence (also called restricted convergence). For the classical definition in the unit ball we refer to [59]. $K^{\prime}$-convergence has been generalized by Abate (see e.g. [1]) to the case of strongly convex domains, and by Abate–Tauraso [8, Section 3] to the case of convex domains of finite type. In our context we introduce the following definition.
Definition 7.5.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$, and let $(z_{n})$ be a sequence in $D$ converging to $\xi$. We say that $(z_{n})$ $K^{\prime}$-converges to $\xi$ (or is restricted) if it $K$-converges to $\xi$ and if
$\lim_{n\to+\infty}k_{D}(z_{n},\varphi(\mathbb{D}))=0.$ (7.3)
If $f\colon D\to\mathbb{C}$ is a function, we say that $f$ has $K^{\prime}$-limit (or restricted $K$-limit) $L\in\mathbb{C}$ at $\xi$ if $f(z_{n})\to L$ for any restricted sequence $(z_{n})$ converging to $\xi$. In this case we write
$K^{\prime}\textrm{-}\lim_{z\to\xi}f(z)=L.$
Remark 7.6.
Let $(z_{n})$ be a restricted sequence converging to $\xi$ and let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. Let $(\zeta_{n})$ be a sequence in $\mathbb{D}$ such that $k_{D}(z_{n},\varphi(\zeta_{n}))\to 0$. Then $\zeta_{n}\to 1$ (non-tangentially), since for all $R>0$,
$\varphi(A^{\mathbb{D}}({\sf id}_{\mathbb{D}},R))=A^{D}(\varphi,R)\cap\varphi(\mathbb{D}).$ (7.4)
Remark 7.7.
This notion of restricted convergence is a priori more general than the one given by Abate–Tauraso [8, Section 3] for a convex domain of finite type $D\subset\mathbb{C}^{d}$. Given a complex geodesic $\varphi$ with endpoint $\xi\in\partial D$ they choose a left inverse $\tilde{\rho}\colon D\to\mathbb{D}$ of $\varphi$. Their definition is (equivalent to) the following: a sequence $(z_{n})$ converging to $\xi$ is restricted if it $K$-converges to $\xi$ and if
$k_{D}(\gamma(t),\varphi(\tilde{\rho}(z_{n})))\to 0.$
Clearly a sequence with this property satisfies (7.3), but the converse is unclear. It is also unclear whether Abate–Tauraso’s definition depends on the choice of the complex geodesic $\varphi$. The next result show that Definition 7.5 does not depend on the choice of the complex geodesic $\varphi$.
Proposition 7.8.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi,\psi$ be two complex geodesics with the same endpoint $\xi$. Then a sequence $(z_{n})$ in $D$ converging to $\xi$ is restricted w.r.t. $\varphi$ if and only if it is restricted w.r.t. $\psi$.
Proof.
Let $\zeta_{n}\in\mathbb{D}$ such that $k_{D}(z_{n},\varphi(\zeta_{n}))\to 0$. By Remark 7.6 $\zeta_{n}\to 1$ non-tangentially. Up to a change of parametrization of $\psi$ we can suppose $\varphi^{\prime}_{N}(1)=\psi^{\prime}_{N}(1)$, so by Proposition 6.7
$k_{D}(z_{n},\psi(\zeta_{n}))\leq k_{D}(z_{n},\varphi(\zeta_{n}))+k_{D}(\varphi(\zeta_{n}),\psi(\zeta_{n}))\to 0.$
We now prove the following extrinsic characterization for $K$-convergence and for $K^{\prime}$-convergence. Notice the analogy with the classic definitions in the ball (see (2.2.33) and (2.2.34) in [1]).
Theorem 7.9.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $(v_{j})_{j=0}^{d-1}$ be a multitype basis at $\xi$, and let $(z_{n})$ be a sequence in $D$. Then
  • (i)
    $(z_{n})$ $K$-converges to $\xi$ if and only if
    $\langle z_{n}-\xi,v_{j}\rangle=O\left(\delta_{D}(z_{n})^{1/m_{j}}\right),\quad\forall\,0\leq j\leq d-1;$ (7.5)
  • (ii)
    $(z_{n})$ $K^{\prime}$-converges to $\xi$ if and only if (7.5) holds for $j=0$ and
    $\langle z_{n}-\xi,v_{j}\rangle=o\left(\delta_{D}(z_{n})^{1/m_{j}}\right),\quad\forall\,1\leq j\leq d-1.$ (7.6)
Proof.
For the sake of clarity, in this proof we denote the sequence in $D$ converging to the origin by $(z^{(n)})$ instead of $(z_{n})$. Consider multitype coordinates at $\xi$. Let $\varphi\colon\mathbb{H}\to D$ be a complex geodesic with endpoint the origin. Up to a change of parametrization, we may assume $\varphi^{\prime}_{N}(0)=1$. Set $\lambda_{n}:=\delta_{D}(z^{(n)})^{-1}$ and consider the scaling
$A_{n}(z)=(\lambda_{n}z_{0},\lambda_{n}^{1/m_{1}}z_{1},\cdots,\lambda_{n}^{1/m_{d-1}}z_{d-1}).$
Notice that in multitype coordinates the sequence $(A_{n}z^{(n)})$ is bounded in $\mathbb{C}^{d}$ if and only if (7.5) holds for all $j=0,\dots,d-1$.
We prove (i). Assume that the sequence $(A_{n}z^{(n)})$ is bounded in $\mathbb{C}^{d}$. Let $r$ be the local defining function of $D$ near the origin as in (5.1). We have that
$\lambda_{n}r(A_{n}^{-1}z)\to r_{D_{H}}(z):={\operatorname{Re}\,}z_{0}+H(z_{1},\cdots,z_{d-1})$
uniformly on compact sets of $\mathbb{C}^{d}$. It is well known (see, for example, [57, Chapter 2, Lemma 2.5]) that for $z\in D$ close to the origin the defining function $r$ is bi-Lipschitz to the distance function $\delta_{D}$, that is there exist $0<c<C$ such that
$C\delta_{D}(z)\leq r(z)\leq-c\delta_{D}(z).$
It follows that
$\limsup_{n\to+\infty}r_{D_{H}}(A_{n}z^{(n)})=\limsup_{n\to+\infty}\lambda_{n}r(A_{n}^{-1}A_{n}z^{(n)})=\limsup_{n\to+\infty}\frac{r(z^{(n)})}{\delta_{D}(z^{(n)})}\leq-c<0,$
which implies that the limit set of the sequence $(A_{n}z^{(n)})$ is contained in a compact subset $K$ of $D_{H}$. By Proposition 6.6 the sequence $(A_{n}\varphi(\lambda_{n}^{-1}\cdot))$ converges uniformly on compact subsets to $\hat{\varphi}(\zeta)=(\zeta,0)$, so by Corollary 6.3 we have
$\limsup_{n\to+\infty}k_{D}(z^{(n)},\varphi(-\lambda_{n}^{-1}))=\limsup_{n\to+\infty}k_{A_{n}D}(A_{n}z^{(n)},A_{n}\varphi(-\lambda_{n}^{-1}))\leq\max_{z\in K}k_{D_{H}}(z,(-1,0))<+\infty,$
which means that the sequence $(z^{(n)})$ is contained in a geodesic region.
Conversely, assume that the sequence $(z^{(n)})$ $K$-converges to $\xi$. Let $(t_{n})$ be a sequence in $(0,1]$ converging to 0 such that $k_{D}(z^{(n)},\varphi(-t_{n}))\leq M$ for all $n\geq 0$. We claim that there are two constants $0<c_{1}<C_{1}$ such that
$c_{1}\leq\lambda_{n}t_{n}\leq C_{1}.$ (7.7)
Indeed, by Lemma 3.8 we can find $c_{2}\geq 0$ such that
$c_{2}\geq|k_{D}(\varphi(-t_{n}),\varphi(-1))+\log\delta_{D}(\varphi(-t_{n}))|=|k_{\mathbb{H}}(-t_{n},-1)+\log\delta_{D}(\varphi(-t_{n}))|=\left|\log\frac{\delta_{D}(\varphi(-t_{n}))}{t_{n}}\right|.$
By Proposition 2.4 in [53] we have $k_{D}(p,q)\geq\left|\log\frac{\delta_{D}(p)}{\delta_{D}(q)}\right|$ for all $p,q\in D$, so
$\left|\log\frac{t_{n}}{\delta_{D}(z^{(n)})}\right|\leq\left|\log\frac{t_{n}}{\delta_{D}(\varphi(-t_{n}))}\right|+\left|\log\frac{\delta_{D}(\varphi(-t_{n}))}{\delta_{D}(z^{(n)})}\right|\leq c_{2}+k_{D}(\varphi(-t_{n}),z^{(n)})\leq c_{2}+M,$
which proves the claim. Thus the sequence $(-\lambda_{n}t_{n})$ is contained in $C_{1},-c_{1}]\subset\subset\mathbb{H}$. Hence the limit set of the sequence
$A_{n}\varphi(-t_{n})=A_{n}\varphi(\lambda_{n}^{-1}(-\lambda_{n}t_{n}))$
is contained in $C_{1},-c_{1}]\times\{0\}\subset\subset D_{H}$. Recalling that $k_{D}(z^{(n)},\varphi(-t_{n}))\leq M$, by (iii) in Corollary 6.3 it follows that the limit set of the sequence $(A_{n}z^{(n)})$ is relatively compact in $D_{H}$. Hence $(A_{n}z^{(n)})$ is bounded in $\mathbb{C}^{d}$.
We now prove (ii). Assume that the sequence $(z^{(n)})$ $K^{\prime}$-converges to $\xi$. By definition there exists $(\zeta_{n})$ in $\mathbb{H}$ such that $k_{D}(z^{(n)},\varphi(\zeta_{n}))\to 0$. By Remark 7.6 $\zeta_{n}\to 0$ non-tangentially, so there exists $A>0$ such that $|{\operatorname{Im}\,}\zeta_{n}|\leq A|{\operatorname{Re}\,}\zeta_{n}|$. It follows that
$k_{D}(z^{(n)},\varphi({\operatorname{Re}\,}\zeta_{n}))\leq k_{D}(z^{(n)},\varphi(\zeta_{n}))+k_{\mathbb{H}}(\zeta_{n},{\operatorname{Re}\,}\zeta_{n})\leq k_{D}(z^{(n)},\varphi(\zeta_{n}))+2\operatorname{arcsinh}\frac{A}{2},$
which is bounded. Arguing as above (setting $t_{n}:=-{\operatorname{Re}\,}\zeta_{n}$) we get that there exists $C_{1}>0$ such that
$c_{1}\leq-\lambda_{n}{\operatorname{Re}\,}\zeta_{n}\leq C_{1},$
which implies that the sequence $(\lambda_{n}\zeta_{n})$ is contained in a compact subset of $\mathbb{H}$. Let $(n_{k})$ be a subsequence such that $(A_{n_{k}}z^{(n_{k})})$ converges to a point $\hat{z}\in D_{H}$ and $(\lambda_{n_{k}}\zeta_{n_{k}})$ converges to a point $\hat{\zeta}\in\mathbb{H}$. We have
$0=\lim_{k\to+\infty}k_{D}(z^{(n_{k})},\varphi(\zeta_{n_{k}}))=\lim_{k\to+\infty}k_{A_{n_{k}}D}(A_{n_{k}}z^{(n_{k})},A_{n_{k}}\varphi(\zeta_{n_{k}}))=k_{D_{H}}(\hat{z},(\hat{\zeta},0)),$
thus $A_{n_{k}}z^{(n_{k})}\to(\hat{\zeta},0)$, which means that $\lambda_{n_{k}}^{1/m_{j}}z^{(n_{k})}_{j}\to 0$ for all $j=1,\cdots,d-1$, which implies (7.6).
Conversely, assume (7.6), and we show that $k_{D}(z^{(n)},\varphi(z^{(n)}_{0}))\to 0$. By (7.5) the sequence $(\lambda_{n}z_{0}^{(n)})$ is bounded in $\mathbb{C}$. Moreover, from the convexity of $D$, we have $\lambda_{n}:=\delta_{D}(z^{(n)})\leq-{\operatorname{Re}\,}z_{0}^{(n)}$, so $(\lambda_{n}z^{(n)}_{0})$ is contained in a compact subset of $\mathbb{H}$. Let $(n_{k})$ be a subsequence such that $(\lambda_{n_{k}}z^{(n_{k})}_{0})$ converges to a point $\hat{\zeta}\in\mathbb{H}$. By (7.6) we have $A_{n_{k}}z^{(n_{k})}\to(\hat{\zeta},0)$, and by Proposition 6.6 we have
$A_{n_{k}}\varphi(\lambda_{n_{k}}^{-1}\lambda_{n_{k}}z_{0}^{(n_{k})})\to(\hat{\zeta},0).$
Thus
$k_{D}(z^{(n_{k})},\varphi(z^{(n_{k})}_{0}))=k_{A_{n_{k}}D}(A_{n_{k}}z^{(n_{k})},A_{n_{k}}\varphi(z^{(n_{k})}_{0}))\to 0.$
Remark 7.10.
Notice that the (i) and (ii) in previous theorem immediately imply the following:
  • (i)
    $\langle z-\xi,v_{j}\rangle=O_{K}\left(\delta_{D}(z)^{1/m_{j}}\right)$ for all $0\leq j\leq d-1$;
  • (ii)
    $\langle z-\xi,v_{j}\rangle=o_{K^{\prime}}\left(\delta_{D}(z)^{1/m_{j}}\right)$ for all $1\leq j\leq d-1$.
Remark 7.11.
It is natural to wonder whether, if a sequence $(z^{(n)})$ $K$-converges to $\xi$, one has for all $v\in\mathbb{C}^{d}\setminus\{0\}$,
$\langle z^{(n)}-\xi,v\rangle=O\left(\delta_{D}(z^{(n)})^{1/m_{\xi}(v)}\right).$
This turns out to be false in general, as the following example shows: let $m\geq 2$ be an even integer and let $\mathbb{E}_{m}\subset\mathbb{C}^{2}$ be the egg domain defined in Example 4.8. Let $\xi:=(1,0)\in\partial\mathbb{E}_{m}$. Fix $\lambda\in\mathbb{D}^{*}$ and consider the curve $\gamma\colon[0,1)\to\mathbb{E}_{m}$ with endpoint $\xi$ defined as $\gamma(t)=(t,\lambda\sqrt[m]{1-t^{2}}),$ which satisfies
$\delta_{\mathbb{E}_{m}}(\gamma(t))\approx 1-t.$
Let $t_{n}\nearrow 1$ and set $z^{(n)}:=\gamma(t_{n}).$ By the previous theorem $z^{(n)}$ converges to $\xi$ inside a $K$-region. Notice that the outer normal versor $n_{\xi}$ at $\xi$ is $(1,0)$ and that the complex tangent at $\xi$ is $T_{\xi}^{\mathbb{C}}\partial\mathbb{E}_{m}=\{z_{0}=0\}$. Now let $v\in\mathbb{C}^{2}\setminus\{z_{0}z_{1}=0\}$. The type of $v$ is $m_{\xi}(v)=1$ while its cotype is $M_{\xi}(v)=m$. A simple computation shows that $\langle z^{(n)}-\xi,v\rangle$ is $O((1-t_{n})^{1/m})$ but is not $O(1-t_{n})$. In general it follows by simple linear algebra that for all $v\in\mathbb{C}^{d}\setminus\{0\}$ one has
$\langle z^{(n)}-\xi,v\rangle=O\left(\delta_{D}(z^{(n)})^{1/M_{\xi}(v)}\right),$
where $M_{\xi}(v)$ is the cotype of $v$. Similar considerations hold for the restricted convergence.
Remark 7.12.
In the unit ball $\mathbb{B}^{d}$ the conditions (7.5) for $j=1,\dots,d-1$ are redundant, that is, a sequence $(z_{n})$ in $\mathbb{B}^{d}$ $K$-converges to $\xi\in\partial\mathbb{B}^{d}$ if and only if $\langle z_{n}-\xi,n_{\xi}\rangle=O\left(\delta_{\mathbb{B}^{d}}(z_{n})\right).$ Indeed one has, for all $M>1$,
$K_{0}(\xi,M)=\left\{z\in\mathbb{B}^{d}:\frac{|1-\langle z,\xi\rangle|}{1-\|z\|}<M\right\}.$ (7.8)
The next result shows that the same is true in $\mathbb{C}$-proper convex domains at strongly linearly convex boundary points. We recall that a $C^{2}$ boundary point of a domain is called strongly linearly convex if the Hessian of some (equivalently, any) defining function is definite positive on the complex tangent space. Notice that every strongly linearly convex point of a convex domain has line type 2.
Corollary 7.13.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a strongly linearly convex boundary point. Let $(z_{n})$ be a sequence in $D$, then $(z_{n})$ $K$-converges to $\xi$ if and only if
$\langle z_{n}-\xi,n_{\xi}\rangle=O\left(\delta_{D}(z_{n})\right).$
Proof.
By (i) in Theorem 7.9 is enough to prove that, if
$\langle z_{n}-\xi,n_{\xi}\rangle=O\left(\delta_{D}(z_{n})\right),$
then
$\langle z_{n}-\xi,\tau\rangle=O\left(\delta_{D}(z_{n})^{1/2}\right)$
for all $\tau\in T_{\xi}^{\mathbb{C}}\partial D$. Since $\xi$ is a strongly linearly convex boundary point, by [49, Proposition 22] there exists $M\geq 0$ such that
$\|z_{n}-\xi\|\leq M|\langle z_{n}-\xi,n_{\xi}\rangle|^{1/2},$
hence for all $\tau\in T_{\xi}^{\mathbb{C}}\partial D$ we have
$|\langle z_{n}-\xi,\tau\rangle|\leq\|z_{n}-\xi\|\leq M|\langle z_{n}-\xi,n_{\xi}\rangle|^{1/2}=O\left(\delta_{D}(z_{n})^{1/2}\right).$
Remark 7.14.
The conditions (7.5) for $j=1,\dots,d-1$ are however not redundant in general. Consider as an example the tube domain
$D:=\{(z_{0},z_{1})\in\mathbb{C}^{2}:{\operatorname{Re}\,}z_{0}+2({\operatorname{Re}\,}z_{1})^{2}<0\},$
a $\mathbb{C}$-proper convex domain which is biholomorphic to the Siegel Half-space
$\mathbb{H}_{2}:=\{(z_{0},z_{1})\in\mathbb{C}^{2}:{\rm Re}\,z_{0}+|z_{1}|^{2}<0\}$
via the biholomorphism $\Psi\colon\mathbb{H}_{2}\to D$ given by
$\Psi(z_{0},z_{1})=\left(z_{0}-z_{1}^{2},z_{1}\right),\quad\Psi^{-1}(z_{0},z_{1})=\left(z_{0}+z_{1}^{2},z_{1}\right).$
In particular $D$ is biholomorphic to the ball $\mathbb{B}^{2}$. Every point of $\partial D$ is strongly pseudoconvex (and thus has line type 2) and no point of $\partial D$ is strongly linearly convex. Let $\alpha>0$ and $t_{n}\searrow 0$, and consider the sequence $z^{(n)}:=(-t_{n},it_{n}^{\alpha})$ converging to the origin. The outer normal versor at the origin is $n_{0}=(1,0)$. Notice that $\tau=(0,i)\in T_{0}^{\mathbb{C}}\partial D$ and that $\delta_{D}(z_{n})=t_{n}$ and
$\langle z^{(n)},n_{0}\rangle=-t_{n},\quad\langle z^{(n)},\tau\rangle=t_{n}^{\alpha}.$
Hence we have $\langle z^{(n)},n_{0}\rangle=O(\delta_{D}(z^{(n)}))$ for all $\alpha>0$, but $\langle z^{(n)},\tau\rangle=O(\delta_{D}(z^{(n)})^{1/2})$ if and only if $\alpha\geq\frac{1}{2}$.
We end this section with another corollary of Theorem 7.9.
Corollary 7.15.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain, let $\xi\in\partial D$ be a point of locally finite type, and let $f\colon D\to\mathbb{C}$ be a function. Then
  • (i)
    if $f$ has $K$-limit $L$ at $\xi$, then $f$ has $K^{\prime}$-limit $L$ at $\xi$;
  • (ii)
    if $f$ has $K^{\prime}$-limit $L$ at $\xi$, then $f$ has non-tangential limit $L$ at $\xi$ (in the sense of cones).
Proof.
(i) follows from the definition of restricted sequence. We prove (ii). Let $(z_{n})$ be a sequence in $D$ converging to $\xi$ inside a cone of aperture $\pi$. Let $B\subseteq D$ be an Euclidean ball internally tangent at $\xi$, then a simple geometric argument shows that there exists $A>1$ such that $\|z_{n}-\xi\|\leq A\delta_{B}(z_{n})$ and obviously $\delta_{B}(z_{n})\leq\delta_{D}(z_{n})$, thus by Theorem 7.9 the sequence $(z_{n})$ is restricted in $D$. ∎
8. Kobayashi type
The following definition was introduced by Abate in [5], see also Abate–Tauraso [8].
Definition 8.1 (Kobayashi type).
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and $\xi\in\partial D$. Let $v\in\mathbb{C}^{d}$, $v\neq 0$, then the Kobayashi type $s_{\xi}(v)$ of $v$ at $\xi$ is the number
$s_{\xi}(v):=\inf\{s>0:\kappa_{D}(z,v)=O_{K}(\delta_{D}(z)^{-s})\}.$
Abate conjectures the following in [5]: “the Kobayashi type might be the inverse of the D’Angelo type of $\partial D$ at $\xi$ along the direction $v$ (that is, the highest order of contact of $\partial D$ with a complex curve tangent to $v$ at $\xi$)”. He also leaves as an open question whether the infimum in the definition of $s_{\xi}(v)$ is always attained. In this section we prove the conjecture (the equality between D’Angelo type at $\xi$ along the direction $v$ and $m_{\xi}(v)$ can be proved as in [17]), and we answer the question in the positive.
Theorem 8.2.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be point of locally finite type. Let $v\in\mathbb{C}^{d}\setminus\{0\}.$ Then the function
$\delta_{D}(z)^{1/m_{\xi}(v)}\kappa_{D}(z,v)$
and its reciprocal are $K$-bounded at $\xi$. In particular, the infimum in the definition of $s_{\xi}(v)$ is attained and
$s_{\xi}(v)=\frac{1}{m_{\xi}(v)}.$
Proof.
Let $(m_{j})_{j=0}^{d-1}$ be the multitype at $\xi$ and consider multitype coordinates at $\xi$. Write $v=\sum_{j=0}^{d-1}a_{j}e_{j}$. Denote
$j_{0}:=\min\{j:a_{j}\neq 0\},$
and notice that $m_{\xi}(v)=m_{j_{0}}$. We first show that $\delta_{D}(z)^{1/m_{\xi}(v)}\kappa_{D}(z,v)$ is $K$-bounded. Assume by contradiction that there exists a sequence $z^{(n)}\to 0$ inside a geodesic region such that
$\delta_{D}^{1/m_{\xi}(v)}(z^{(n)})\kappa_{D}(z^{(n)},v)\to+\infty.$
Set $\lambda_{n}:=\delta_{D}(z^{(n)})^{-1}$ and consider the scaling maps
$A_{n}(z)=(\lambda_{n}z_{0},\lambda_{n}^{1/m_{1}}z_{1},\cdots,\lambda_{n}^{1/m_{d-1}}z_{d-1}).$
Arguing as in (i) in Theorem 7.9, we obtain that up to a subsequence
$A_{n}z^{(n)}\to z^{(\infty)}\in D_{H}.$
Now
$\infty=\lim_{n\to+\infty}\lambda_{n}^{-1/m_{\xi}(v)}\kappa_{D}(z^{(n)},v)=\lim_{n\to+\infty}\kappa_{A_{n}D}(A_{n}z^{(n)},\lambda_{n}^{-1/m_{\xi}(v)}A_{n}v).$
Notice that
$\lambda_{n}^{-1/m_{\xi}(v)}A_{n}v=\sum_{j=j_{0}}^{d-1}\lambda_{n}^{1/m_{j}-1/m_{\xi}(v)}a_{j}e_{j}.$
Since $m_{j}\geq m_{\xi}(v)$ for $j=j_{0},\cdots,d-1$, it follows that
$\lambda_{n}^{-1/m_{\xi}(v)}A_{n}v\to v_{\infty}:=\sum_{j\colon m_{j}=m_{\xi}(v)}a_{j}e_{j}.$
Finally
$\infty=\lim_{n\to+\infty}\kappa_{A_{n}D}(A_{n}z^{(n)},\lambda_{n}^{-1/m_{\xi}(v)}A_{n}v)=\kappa_{D_{H}}(z^{(\infty)},v_{\infty})<\infty,$
obtaining a contradiction. Now we show that the reciprocal is $K$-bounded. Assume by contradiction that there exists a sequence $z^{(n)}\to 0$ inside a geodesic region such that
$\delta_{D}^{1/m_{\xi}(v)}(z^{(n)})\kappa_{D}(z^{(n)},v)\to 0.$
Reasoning as before and noting that $v_{\infty}\neq 0$, we obtain
$0=\lim_{n\to+\infty}\lambda_{n}^{-1/m_{\xi}(v)}\kappa_{D}(z^{(n)},v)=\lim_{n\to+\infty}\kappa_{A_{n}D}(A_{n}z^{(n)},\lambda_{n}^{-1/m_{\xi}(v)}A_{n}v)=\kappa_{D_{H}}(z^{(\infty)},v_{\infty})>0,$
obtaining a contradiction. ∎
Remark 8.3.
The previous result can be rephrased as the following estimate for the Kobayashi metric: if $D\subset\mathbb{C}^{d}$ is a $\mathbb{C}$-proper convex domain and $\xi\in\partial D$ is a point of locally finite type, then for all $v\in\mathbb{C}^{d}\setminus\{0\}$, $p\in D$, $M>1$ there exist $c,C>0$ such that for all $z\in K_{p}(\xi,M)$ we have
$\frac{c}{\delta_{D}(z)^{1/m_{\xi}(v)}}\leq\kappa_{D}(z,v)\leq\frac{C}{\delta_{D}(z)^{1/m_{\xi}(v)}}.$
This estimate was obtained along the normal segment by Lee in [51] (see also [55] for the $\mathbb{C}$-convex case).
9. Lindelöf principle
The purpose of this section is to prove the following version of the Lindelöf principle.
Definition 9.1.
A restricted curve $\gamma:[0,1)\to D$ with endpoint $\xi$ is a curve such that for all sequences $(t_{n})$ in $[0,1)$ converging to 1 we have that the sequence $(\gamma(t_{n}))$ is restricted in the sense of Definition 7.5. Clearly this is equivalent to saying that $\gamma([0,1))$ is contained in a $K$-region with vertex $\xi$ and that $k_{D}(\gamma(t),\varphi(\mathbb{D}))\stackrel{{\scriptstyle t\to 1^{-}}}{{\longrightarrow}}0,$ where $\varphi$ is a complex geodesic with endpoint $\xi$. We say that curve $\gamma:[0,1)\to D$ with endpoint $\xi$ is strongly restricted if $\gamma([0,1))$ is contained in a $K$-region with vertex $\xi$ and if there exists a (continuous) curve $\tilde{\gamma}\colon[0,1)\to\mathbb{D}$ such that
$\lim_{t\to 1^{-}}k_{D}(\gamma(t),\varphi(\tilde{\gamma}(t)))=0,$ (9.1)
where $\varphi$ is a complex geodesic with endpoint $\xi$.
Remark 9.2.
If $\gamma$ is strongly restricted, the curve $\varphi\circ\tilde{\gamma}$ is contained in a geodesic region with vertex $\xi$, and thus by (7.4) it follows that the curve $\tilde{\gamma}(t)$ has endpoint 1 and is non-tangential. Notice also that the definition of strongly restricted curve does not depend on the choice of the complex geodesic $\varphi$.
The definition of strongly restricted curve is only slightly more stringent than the definition of restricted curve, as the following lemma shows.
Lemma 9.3.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Then any rectifiable restricted curve with endpoint $\xi$ is strongly restricted.
Proof.
Define a sequence $(t_{n})$ as follows. Set $t_{0}=0$. Then define inductively $t_{n+1}$ as the minimum $t\geq t_{n}$ such that $k_{D}(\gamma(t),\gamma(t_{n}))\geq 1/n.$ It easily follows from the fact that $\gamma$ is rectifiable that $t_{n}\to 1$. Set $z_{n}:=\gamma(t_{n})$. Since the sequence $(z_{n})$ is restricted, there exists a sequence $(\zeta_{n})$ in $\mathbb{D}$ such that $k_{D}(z_{n},\varphi(\zeta_{n}))\to 0$. Define the curve $\tilde{\gamma}\colon[0,1)\to\mathbb{D}$ as follows: for all $n\geq 0$ set $\tilde{\gamma}(t_{n}):=\zeta_{n},$ and then interpolate for all $t\in[0,1)$ using (suitably reparametrized) geodesic segments. It is easy to see that $\lim_{t\to 1^{-}}k_{D}(\gamma(t),\varphi(\tilde{\gamma}(t)))=0.$
Theorem 9.4 (Lindelöf Principle).
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $f\colon D\to\mathbb{C}$ be a holomorphic function which is $K$-bounded at $\xi$, and assume that there exists a strongly restricted curve $\gamma\colon[0,1)\to D$ such that $f(\gamma(t))\to L$ as $t\to 1$. Then $f$ has restricted $K$-limit $L$ at $\xi$.
Remark 9.5.
This statement is only apparently similar to [8, Theorem 3.2], given our more general notion of restricted convergence, see Remark 7.7 (moreover, we assume $K$-boundedness instead of $T$-boundedness). The proof is also different, as it is only based on the one variable Lindelöf Principle, the Royden localization lemma and metric arguments.
Lemma 9.6 (Royden Localization Lemma).
[58] Let $D\subset\mathbb{C}^{d}$ be a Kobayashi hyperbolic domain and let $U\subset D$ be a domain. Then
$\kappa_{U}(z,v)\leq\coth(k_{D}(z,D\backslash U)/2)\kappa_{D}(z,v)$
for any $z\in U$ and $v\in\mathbb{C}^{d}$, where $k_{D}(z,D\backslash U):=\inf_{w\in D\backslash U}k_{D}(z,w)$.
Lemma 9.7.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $(z_{n})$ be a restricted sequence converging to $\xi$, and let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. Let $R>0$ be such that $z_{n}\in A(\varphi,R)$ for all $n\geq 0$. Let $(\zeta_{n})$ be a sequence in $\mathbb{D}$ such that $k_{D}(z_{n},\varphi(\zeta_{n}))\to 0.$ Then for all $R^{\prime}>R$ the sequence $(\varphi(\zeta_{n}))$ is eventually contained in $A(\varphi,R^{\prime}$) and we have
$\lim_{n\to+\infty}k_{A(\varphi,R^{\prime})}(z_{n},\varphi(\zeta_{n}))=0.$
Proof.
For all $n\geq 0$, let $\gamma_{n}\colon[0,T_{n}]\to D$ be a geodesic segment between $z_{n}$ and $\varphi(\zeta_{n})$ with respect to $k_{D}$. Fix $R<S<R^{\prime}$. Clearly $\gamma_{n}([0,T_{n}])$ is eventually contained in $A(\gamma,S)$. By the Royden localization lemma, we have
$\kappa_{A(\varphi,R^{\prime})}(z,v)\leq B\kappa_{D}(z,v),\ \ \ z\in A(\varphi,S),v\in\mathbb{C}^{d},$
where $B:=\coth(R^{\prime}-S)>1$, which implies
$k_{A(\varphi,R^{\prime})}(z_{n},\varphi(\zeta_{n}))$ $\leq\int_{0}^{T_{n}}\kappa_{A(\varphi,R^{\prime})}(\gamma_{n}(t),\gamma_{n}^{\prime}(t))dt$
$\leq B\int_{0}^{T_{n}}\kappa_{D}(\gamma_{n}(t),\gamma_{n}^{\prime}(t))dt$
$=Bk_{D}(z_{n},\varphi(\zeta_{n}))\to 0.$
Remark 9.8.
It follows from the previous lemma that, if $\gamma\colon[0,1)\to D$ is a strongly restricted curve such that $\gamma(t)\in A(\varphi,R)$ for all $t\in[0,1)$ and $\tilde{\gamma}\colon[0,1)\to\mathbb{D}$ is a curve satisfying (9.1), then for all $R^{\prime}>R$ we have that $\varphi(\tilde{\gamma}(t))$ is eventually contained in $A(\varphi,R^{\prime})$ and
$\lim_{t\to 1^{-}}k_{A(\varphi,R^{\prime})}(\gamma(t),\varphi(\tilde{\gamma}(t)))=0.$
Proof of Theorem 9.4.
Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$ and let $\tilde{\gamma}\colon[0,1)\to\mathbb{D}$ be a curve satisfying (9.1). Let $R>0$ such that $\gamma(t)\in A(\varphi,R)$ for all $t\in[0,1)$, and let $R^{\prime}>R$. Since $f$ is $K$-bounded at $\xi$ there exists $S>0$ such that $f(A(\varphi,R^{\prime}))\subset\subset S\mathbb{D}$, and thus by Lemma 9.7
$k_{S\mathbb{D}}(f(\gamma(t)),f(\varphi(\tilde{\gamma}(t))))\leq k_{A(\varphi,R^{\prime})}(\gamma(t),\varphi(\tilde{\gamma}(t)))\stackrel{{\scriptstyle t\to 1^{-}}}{{\longrightarrow}}0,$
thus
$\lim_{t\to 1^{-}}|f(\gamma(t))-f(\varphi(\tilde{\gamma}(t)))|=0.$
Since $f(\gamma(t))\to L$ as $t\to 1^{-}$ it follows that $f(\varphi(\tilde{\gamma}(t)))\to L$. Applying the classical Lindelöf Theorem to the holomorphic function $f\circ\varphi\colon\mathbb{D}\to\mathbb{C}$ we have that $f\circ\varphi$ has non-tangential limit $L$ at $1$.
Let now $(z_{n})$ be a restricted sequence in $D$ converging to $\xi$, and let $(\zeta_{n})$ be a sequence in $\mathbb{D}$ such that $k_{D}(z_{n},\varphi(\zeta_{n}))\to 0.$ Arguing as above we obtain that
$\lim_{n\to+\infty}|f(z_{n})-f(\varphi(\zeta_{n}))|=0.$ (9.2)
Since $(\zeta_{n})$ converges to $1$ non-tangentially, we have $f(\varphi(\zeta_{n}))\to L$. By (9.2) we obtain $f(z_{n})\to L$ as $n\to+\infty$, and we are done. ∎
10. Julia lemma
The main result in this section is the generalization to our setting of the intrinsic Julia lemma proved by Abate for bounded strongly convex domains with $C^{3}$ boundary [1, Theorem 2.4.16, Proposition 2.7.15]. For bounded convex domains of finite type this is already done in [12, Theorem 6.28]. In our local finite type assumptions the statement and the proof have to be adapted since the classical argument which gives the existence of the $K$-limit at $\xi$ does not work. Lemma 10.4 below is the main tool to deal with this issue.
Definition 10.1 (Dilation).
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D$ be a point of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map, and fix base-points $p\in D,p^{\prime}\in D^{\prime}$. The dilation $\lambda_{\xi,p,p^{\prime}}\in(0,+\infty]$ of $f$ at $\xi$ is the number defined by
$\log\lambda_{\xi,p,p^{\prime}}=\liminf_{z\to\xi}k_{D}(z,p)-k_{D^{\prime}}(f(z),p^{\prime}).$
If $\lambda_{\xi,p,p^{\prime}}<+\infty,$ we call a sequence $(z_{n})$ in $D$ converging to $\xi$ dilation minimizing if
$\lim_{n\to+\infty}k_{D}(z_{n},p)-k_{D^{\prime}}(f(z_{n}),p^{\prime})=\log\lambda_{\xi,p,p^{\prime}},$
while we say that $(z_{n})$ has bounded dilation if the sequence $(k_{D}(z_{n},p)-k_{D^{\prime}}(f(z_{n}),p^{\prime}))$ is bounded from above. Notice that the notion of bounded dilation is base-point independent, that is, if a sequence has bounded dilation w.r.t base-points $p,p^{\prime}$, then by the triangle inequality it has bounded dilation also w.r.t. any $q,q^{\prime}$. In case $D^{\prime}=\mathbb{D}$ we simply denote $\lambda_{\xi,p}:=\lambda_{\xi,p,0}$.
Remark 10.2.
Notice that $\log\lambda_{\xi,p,p^{\prime}}$ cannot be -$\infty$. Indeed,
$k_{D}(z,p)-k_{D^{\prime}}(f(z),p^{\prime})\geq k_{D^{\prime}}(f(z),f(p))-k_{D^{\prime}}(f(z),p^{\prime})\geq-k_{D^{\prime}}(f(p),p^{\prime}).$
Remark 10.3.
A sequence $(z_{n})$ in $D$ satisfies $h_{\xi,p}(z_{n})\to$-$\infty$ if and only if it is eventually contained in every horosphere centered at $\xi$. This is the case for example if $(z_{n})$ $K$-converges to $\xi$ (by (iii) in Lemma 7.3).
Lemma 10.4.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type, and let $p\in D,p^{\prime}\in D^{\prime}$. Let $f\colon D\to D^{\prime}$ be a holomorphic map such that $\lambda_{\xi,p,p^{\prime}}<+\infty$. Let $(z_{n})$ be a sequence in $D$ converging to $\xi$ inside a horosphere centered at $\xi$ such that $f(z_{n})\to\eta$. If $(x_{n})$ is a sequence in $D$ converging to $\xi$ with bounded dilation, then we have $f(x_{n})\to\eta$.
Proof.
Let $A\geq 0$ be such that for all $n\geq 0$ we have
$k_{D}(x_{n},p)-k_{D^{\prime}}(f(x_{n}),p^{\prime})\leq A,$
so $k_{D^{\prime}}(f(x_{n}),p^{\prime})\geq k_{D}(x_{n},p)-A\to+\infty$. Hence, up to subsequence, $(f(x_{n}))$ converges to a point $\nu\in\partial^{*}D^{\prime}.$ Now
$2(f(x_{n})|f(z_{m}))_{p^{\prime}}$ $=k_{D^{\prime}}(f(x_{n}),p^{\prime})+k_{D^{\prime}}(f(z_{m}),p^{\prime})-k_{D^{\prime}}(f(x_{n}),f(z_{m}))$
$\geq k_{D}(x_{n},p)+k_{D^{\prime}}(f(z_{m}),p^{\prime})-k_{D^{\prime}}(f(x_{n}),f(z_{m}))-A$
$\geq k_{D}(x_{n},p)+k_{D^{\prime}}(f(z_{m}),p^{\prime})-k_{D}(x_{n},z_{m})-A.$
Since $h_{\xi,p}(z_{m}):=\lim_{n\to+\infty}k_{D}(z_{m},x_{n})-k_{D}(x_{n},p)$ is bounded from above there exists a constant $M\geq 0$ and a subsequence $(n_{m})$ such that for all $m\geq 0$ we have
$k_{D}(z_{m},x_{n_{m}})-k_{D}(x_{n_{m}},p)\leq M.$
Hence
$\lim_{m\to+\infty}(f(x_{n_{m}})|f(z_{m}))=+\infty,$
which implies by (i) in Proposition 3.6 that $\nu=\eta$. ∎
Lemma 10.5.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type, let $p,q\in D$ and $p^{\prime},q^{\prime}\in D^{\prime}$. Let $f\colon D\to D^{\prime}$ be a holomorphic map such that $\lambda_{\xi,p,p^{\prime}}<+\infty$ and assume that there exists a sequence $(z_{n})$ in $D$ converging to $\xi$ inside a horosphere centered at $\xi$ such that $f(z_{n})\to\eta$. Then
  • (i)
    $\log\lambda_{\xi,q,q^{\prime}}=\log\lambda_{\xi,p,p^{\prime}}+h^{D}_{\xi,p}(q)+h^{D^{\prime}}_{\eta,q^{\prime}}(p^{\prime});$
  • (ii)
    the notion of dilation minimizing is base-point independent, that is any dilation minimizing sequence w.r.t. base-points $p,p^{\prime}$ is dilation minimizing also w.r.t. any $q,q^{\prime}$.
Proof.
Let $w_{n}\to\xi$ be a dilation minimizing sequence w.r.t. $p,p^{\prime}$. We have
$k_{D}(q,w_{n})-k_{D^{\prime}}(q^{\prime},f(w_{n}))$ $=k_{D}(p,w_{n})-k_{D^{\prime}}(p^{\prime},f(w_{n}))+k_{D}(q,w_{n})-k_{D}(p,w_{n})$
$+k_{D^{\prime}}(p^{\prime},f(w_{n}))-k_{D^{\prime}}(q^{\prime},f(w_{n})).$
By Lemma 10.4 $f(w_{n})\to\eta$, and by Theorem 6.10 it follows that
$k_{D}(q,w_{n})-k_{D^{\prime}}(q^{\prime},f(w_{n}))\stackrel{{\scriptstyle n\to+\infty}}{{\longrightarrow}}\log\lambda_{\xi,p,p^{\prime}}+h^{D}_{\xi,p}(q)+h^{D^{\prime}}_{\eta,q^{\prime}}(p^{\prime}).$ (10.1)
Hence
$\log\lambda_{\xi,q,q^{\prime}}\leq\log\lambda_{\xi,p,p^{\prime}}+h^{D}_{\xi,p}(q)+h^{D^{\prime}}_{\eta,q^{\prime}}(p^{\prime}),$
and considering a minimizing sequence w.r.t. $q,q^{\prime}$ we obtain (i). Notice that (10.1) together with (i) implies (ii). ∎
Proposition 10.6 (Julia Lemma).
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map, and let $p\in D$ and $p^{\prime}\in D^{\prime}$. The following are equivalent:
  1. (1)
    $\lambda_{\xi,p,p^{\prime}}<+\infty$ and there exists a sequence $(z_{n})$ in $D$ converging to $\xi$ inside a horosphere centered at $\xi$ such that $f(z_{n})\to\eta$;
  2. (2)
    for all $z\in D$
    $h^{D^{\prime}}_{\eta,p^{\prime}}(f(z))-h^{D}_{\xi,p}(z)\leq\log\lambda_{\xi,p,p^{\prime}},$
    or equivalently,
    $f(E^{D}_{p}(\xi,R))\subseteq E^{D^{\prime}}_{p^{\prime}}(\eta,\lambda_{\xi,p,p^{\prime}}R),\quad\forall R>0;$
  3. (3)
    the function $h^{D^{\prime}}_{\eta,p^{\prime}}(f(z))-h^{D}_{\xi,p}(z)$ is bounded from above;
  4. (4)
    $\lambda_{\xi,p,p^{\prime}}<+\infty$ and $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$.
Proof.
We prove $(1)\Rightarrow(2)$. Let $(w_{n})$ be a sequence in $D$ converging to $\xi$ such that
$k_{D}(w_{n},p)-k_{D^{\prime}}(f(w_{n}),p^{\prime})\to\log\lambda_{\xi,p,p^{\prime}}.$
By Lemma 10.4 we have $f(w_{n})\to\eta$. Then (2) follows from the metric Julia Lemma [12, Lemma 6.14]. $(2)\Rightarrow(3)$ is trivial.
We prove $(3)\Rightarrow(4)$. If $(w_{n})$ is a sequence $K$-converging to $\xi$, then by (iii) in Lemma 7.3 $h^{D}_{\xi,p}(w_{n})\to$-$\infty$, and thus $h^{D^{\prime}}_{\eta,p^{\prime}}(f(w_{n}))\to-\infty.$ By Remark 6.12 it follows that $f(w_{n})\to\eta$, and thus $\eta$ is the $K$-limit of $f$ at $\xi$. Moreover, if $\gamma$ is a geodesic ray with $\gamma(0)=p$ and endpoint $\xi$, then for all $t\geq 0$ we have
$k_{D}(p,\gamma(t))-k_{D^{\prime}}(p^{\prime},f(\gamma(t)))\leq h^{D^{\prime}}_{\eta,p^{\prime}}(f(\gamma(t)))-h^{D}_{\xi,p}(\gamma(t)),$ (10.2)
and thus $\lambda_{\xi,p,p^{\prime}}<+\infty.$ $(4)\Rightarrow(1)$ is trivial.
Remark 10.7.
If the codomain $D^{\prime}\subset\mathbb{C}^{q}$ is a bounded convex domain of finite type, then it is easy to see that if $\lambda_{\xi,p,p^{\prime}}<+\infty,$ then there exists $\eta\in\partial D^{\prime}$ such that $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$.
Definition 10.8.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D$ be a point of locally finite type. Let $f\colon D\to D^{\prime}$ holomorphic, and fix base-points $p\in D,p^{\prime}\in D^{\prime}$. We say that $\xi$ is a regular contact point if $\lambda_{\xi,p,p^{\prime}}<+\infty$ and if $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$, where $\eta$ is a point of locally finite type in $\partial D^{\prime}$.
Remark 10.9.
By Lemma 10.5 the previous definition does not depend on the chosen base-points.
The next result is an important consequence of the Julia Lemma.
Proposition 10.10.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $f\colon D\to D^{\prime}$ be a holomorphic map, and let $p\in D$ and $p^{\prime}\in D^{\prime}$. Let $\xi\in\partial D$, $\eta\in\partial D^{\prime}$ be points of locally finite type and assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then, if $\gamma\colon\mathbb{R}_{\geq 0}\rightarrow D$ is a geodesic ray with endpoint $\xi$, we have
$\lim_{t\to+\infty}k_{D}(\gamma(t),p)-k_{D^{\prime}}(f(\gamma(t)),p^{\prime})=\log\lambda_{\xi,p,p^{\prime}}.$ (10.3)
As a consequence
$\sup_{z\in D}\left(h^{D^{\prime}}_{\eta,p^{\prime}}(f(z))-h^{D}_{\xi,p}(z)\right)=\log\lambda_{\xi,p,p^{\prime}}.$ (10.4)
Proof.
Set $q:=\gamma(0)$. Arguing as in [13, Lemma 3.12] using (2) in Proposition 10.6 we have that
$\lim_{t\to+\infty}k_{D}(\gamma(t),q)-k_{D^{\prime}}(f(\gamma(t)),p^{\prime})=\log\lambda_{\xi,q,p^{\prime}}.$
Then (10.3) follows from (ii) in Lemma 10.5. Equation (10.4) now follows using (10.2). ∎
The definition of regular contact point given above is different from the one given in Definition 4.5 for holomorphic maps with domain of definition $\mathbb{D}$. The next result shows that the two definitions are in fact equivalent. The proof is similar to [13, Proposition 3.8, Lemma 3.13] so we omit it.
Proposition 10.11.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points locally finite type, and let $p\in D,p^{\prime}\in D^{\prime}$. Let $f\colon D\to D^{\prime}$ be a holomorphic map. The following are equivalent.
  1. (1)
    $\lambda_{\xi,p,p^{\prime}}<+\infty$ and $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$;
  2. (2)
    there exists a geodesic ray $\gamma$ with endpoint $\xi$ such that $f\circ\gamma$ is a $(1,B)$ quasi-geodesic with endpoint $\eta$;
  3. (3)
    for all geodesic rays $\gamma$ with endpoint $\xi$ the curve $f\circ\gamma$ is an almost-geodesic with endpoint $\eta$.
We now prove other consequences of the Julia lemma. The first is that $K$-regions are mapped into $K$-regions.
Corollary 10.12.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type, and let $p\in D$. Let $f\colon D\to D^{\prime}$ be holomorphic, and assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then for all $M>1$,
$f(K_{p}(\xi,M))\subseteq K_{f(p)}(\eta,\lambda_{\xi,p,f(p)}^{1/2}M).$
Proof.
Let $z\in K_{p}(\xi,M)$, then by the Julia Lemma
$h_{\eta,f(p)}(f(z))+k_{D^{\prime}}(f(z),f(p))\leq h_{\xi,p}(z)+\log\lambda_{\xi,p,f(p)}+k_{D}(z,p)$
$<2\log M+\log\lambda_{\xi,p,f(p)}=2\log(\lambda_{\xi,p,f(p)}^{1/2}M).$
Another corollary of the Julia Lemma is a chain rule for the dilation.
Corollary 10.13.
Let $D\subset\mathbb{C}^{d}$, $D^{\prime}\subset\mathbb{C}^{d^{\prime}}$, and $D^{\prime\prime}\subset\mathbb{C}^{d^{\prime\prime}}$ be $\mathbb{C}$-proper convex domains. Let $f\colon D\to D^{\prime}$, $g\colon D^{\prime}\to D^{\prime\prime}$ be holomorphic maps and assume that $\xi\in\partial D$ is a regular contact point of $f$ with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta\in\partial D^{\prime}$, and that $\eta$ is a regular contact point of $g$ with $K\textrm{-}\lim_{w\to\eta}g(w)=\zeta\in\partial D^{\prime\prime}$. Then $\xi$ is a regular contact point of $g\circ f$ with $K\textrm{-}\lim_{z\to\xi}g(f(z))=\zeta$ and
$\lambda_{\xi,p,p^{\prime\prime}}(g\circ f)=\lambda_{\xi,p,p^{\prime}}(f)\cdot\lambda_{\eta,p^{\prime},p^{\prime\prime}}(g),$
where $p\in D,p^{\prime}\in D^{\prime},p^{\prime\prime}\in D^{\prime\prime}$.
Proof.
By the Julia Lemma we obtain
$h_{\zeta,p^{\prime\prime}}(g(f(z)))\leq h_{\xi,p}(z)+\log\lambda_{\xi,p,p^{\prime}}(f)+\log\lambda_{\eta,p^{\prime},p^{\prime\prime}}(g),\quad\forall z\in D.$
Again by the Julia Lemma this implies that $\xi$ is a regular contact point of the map $g\circ f$, and that the boundary point $\zeta$ is the $K$-limit of $g\circ f$ at $\xi$. By (10.4) we obtain
$\log\lambda_{\xi,p,p^{\prime\prime}}(g\circ f)\leq\log\lambda_{\xi,p,p^{\prime}}(f)+\log\lambda_{\eta,p^{\prime},p^{\prime\prime}}(g).$
On the other hand, if $\gamma$ is a geodesic ray in $D$ with endpoint $\xi$, by (10.3) we have
$\log\lambda_{\xi,p,p^{\prime\prime}}(g\circ f)=\lim_{t\to+\infty}k_{D}(\gamma(t),p)-k_{D^{\prime\prime}}(g(f(\gamma(t))),p^{\prime\prime})$
$\geq\lim_{t\to+\infty}k_{D}(\gamma(t),p)-k_{D^{\prime}}(f(\gamma(t)),p^{\prime})+\liminf_{t\to+\infty}k_{D^{\prime}}(f(\gamma(t)),p^{\prime})-k_{D^{\prime\prime}}(g(f(\gamma(t))),p^{\prime\prime})$
$\geq\log\lambda_{\xi,p,p^{\prime}}(f)+\log\lambda_{\eta,p^{\prime},p^{\prime\prime}}(g).$
We end this section with an extrinsic characterization of regular contact points.
Proposition 10.14.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map and assume $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. The following are equivalent.
  1. (1)
    $\xi$ is a regular contact point;
  2. (2)
    $\delta_{D^{\prime}}(f(z))/\delta_{D}(z)$ is $K$-bounded at $\xi$;
  3. (3)
    there exists a sequence $(z_{n})$ in $D$, $K$-converging to $\xi$, such that the sequence $(\delta_{D^{\prime}}(f(z_{n}))/\delta_{D}(z_{n}))$ is bounded.
Proof.
We prove $(1)\Rightarrow(2)$. Let $\gamma\colon\mathbb{R}_{\geq 0}\to D$ be a geodesic ray with endpoint $\xi$. Since $\xi$ is a regular contact point, by (10.3) there exists $A\geq 0$ such that for all $t\geq 0$
$k_{D}(\gamma(t),p)-k_{D^{\prime}}(f(\gamma(t)),p^{\prime})\leq A.$
Fix $R>0$ and let $z\in A(\gamma,R)$ be close to $\xi$, then by definition there exists $t^{*}\geq 0$ such that $k_{D}(z,\gamma(t^{*}))<R$. Now using Lemma 3.8 for $D$ near $\xi$ and $D^{\prime}$ near $\eta$ we can find $c\geq 0$ such that
$\log\frac{\delta_{D^{\prime}}(f(z))}{\delta_{D}(z)}$ $\leq k_{D}(z,p)-k_{D^{\prime}}(f(z),p^{\prime})+c$
$\leq k_{D}(z,\gamma(t^{*}))+k_{D}(\gamma(t^{*}),p)+k_{D^{\prime}}(f(z),f(\gamma(t^{*})))-k_{D^{\prime}}(f(\gamma(t^{*})),p^{\prime})+c$
$\leq 2R+A+c.$
The implication $(2)\Rightarrow(3)$ is trivial, and $(3)\Rightarrow(1)$ follows from Lemma 3.8. ∎
The previous result can be strengthened if the codomain $D^{\prime}\subset\mathbb{C}^{q}$ is a bounded convex domain of finite type.
Proposition 10.15.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain, and let $D^{\prime}\subset\mathbb{C}^{q}$ be a bounded convex domain of finite type. Let $\xi\in\partial D$ be a point of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map. Then the following are equivalent:
  1. (1)
    $\xi$ is a regular contact point;
  2. (2)
    $\delta_{D^{\prime}}(f(z))/\delta_{D}(z)$ is $K$-bounded at $\xi$;
  3. (3)
    $\liminf_{z\to\xi}\delta_{D^{\prime}}(f(z))/\delta_{D}(z)<+\infty.$
Proof.
If $\xi$ is a regular contact point then (2) holds by the previous proposition. (2) $\Rightarrow$ (3) is trivial. If (3) holds, let $(w_{n})$ be a sequence in $D$ converging to $\xi$ such that $\delta_{D^{\prime}}(f(w_{n}))/\delta_{D}(w_{n})$ is bounded. Up to a subsequence we may assume that the sequence $(f(w_{n}))$ converges to a point in $\eta\in\overline{D^{\prime}}$. Since $\delta_{D}(w_{n})\to 0$ we have $\delta_{D^{\prime}}(f(w_{n}))\to 0$, and thus $\eta\in\partial D^{\prime}$. By Lemma 3.8 and Remark 10.7 $\xi$ is a regular contact point. ∎
11. Pluricomplex Poisson kernel and normalized dilation
In our generalization of the Julia-Wolff-Carathéodory theorem we want to establish a relationship between the dilation $\lambda_{\xi,p,p^{\prime}}(f)$ of a holomorphic map at a regular contact point $\xi\in\partial D$ (with $K$-limit $\eta\in\partial D^{\prime}$) and the restricted $K$-limit of the normal component of the normal directional derivative $\langle df_{z}(n_{\xi}),n_{\eta}\rangle$. The fact that the dilation $\lambda_{\xi,p,p^{\prime}}(f)$ depends on the base-points $p\in D,p^{\prime}\in D^{\prime}$ is, from this point of view, a disturbance. We now introduce, in the setting of $\mathbb{C}$-proper convex domains of finite type, a function which generalizes the pluricomplex Poisson kernel originally defined (with completely different methods) by Bracci–Patrizio–Trapani [22, 23] on strongly convex domains with $C^{\infty}$ boundary, and we use it to normalize the dilation and eliminate the dependence on the base-points.
Recall that by Corollary 6.8 two complex geodesics with the same endpoint $\xi\in\partial D$ are strongly asymptotic.
Lemma 11.1.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be of locally finite type. Let $\varphi,\psi\colon\mathbb{D}\to D$ be two complex geodesics with the same endpoint $\xi$. Let $T\in\mathbb{R}$ be such that $k_{D}(\tilde{\varphi}(t),\tilde{\psi}(t+T))\to 0.$ Then
$T=h_{\xi,\varphi(0)}(\psi(0)).$
In particular, if $\varphi(0)=\psi(0)$, then $T=0.$
Proof.
We have
$h_{\xi,\varphi(0)}(\psi(0))=\lim_{t\to+\infty}k_{D}(\tilde{\psi}(0),\tilde{\varphi}(t))-k_{D}(\tilde{\varphi}(t),\tilde{\varphi}(0))=\lim_{t\to+\infty}k_{D}(\tilde{\psi}(0),\tilde{\psi}(t+T))-k_{D}(\tilde{\varphi}(t),\tilde{\varphi}(0))=T.$
Proposition 11.2.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi,\psi\colon\mathbb{D}\to D$ be two complex geodesics with the same endpoint $\xi$. Let $T\in\mathbb{R}$ be such that $k_{D}(\tilde{\varphi}(t),\tilde{\psi}(t+T))\to 0.$ Then
$T=\log\frac{\psi^{\prime}_{N}(1)}{\varphi_{N}^{\prime}(1)},$ (11.1)
and thus
$h_{\xi,\varphi(0)}(\psi(0))=\log\frac{\psi^{\prime}_{N}(1)}{\varphi_{N}^{\prime}(1)}.$ (11.2)
In particular if $\varphi(0)=\psi(0)$, then $\psi^{\prime}_{N}(1)=\varphi_{N}^{\prime}(1).$
Proof.
Let $a\in(-1,1)$ be such that $\frac{1-a}{1+a}=\frac{\varphi^{\prime}_{N}(1)}{\psi_{N}^{\prime}(1)}$ and let $\tau_{a}\colon\mathbb{D}\to\mathbb{D}$ be the automorphism of the disc given by $\tau_{a}(\zeta)=\frac{\zeta+a}{1+a\zeta}$. Notice that $\tau^{\prime}_{a}(1)=\frac{1-a}{1+a}.$ The complex geodesic $\psi_{a}:=\psi\circ\tau_{a}$ satisfies $(\psi_{a})^{\prime}_{N}(1)=\varphi^{\prime}_{N}(1)$, so by Theorem 6.7 we have
$\lim_{t\to+\infty}{k_{D}}(\tilde{\varphi}(t),\tilde{\psi}_{a}(t))=0.$
A simple computation shows that $\tilde{\psi}_{a}(t)=\tilde{\psi}\left(t+\log\frac{\psi^{\prime}_{N}(1)}{\varphi_{N}^{\prime}(1)}\right)$. Equation 11.2 follows combining (11.1) with Lemma 11.1. ∎
Definition 11.3.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. For all $z\in D$ we define the pluricomplex Poisson kernel $\Omega_{\xi}\colon D\to(-\infty,0)$ as
$\Omega_{\xi}(z)=-\frac{1}{\varphi^{\prime}_{N}(1)},$
where $\varphi\colon\mathbb{D}\to D$ is a complex geodesic such that $\varphi(0)=z$ and with endpoint $\xi$. Notice that it is not known whether there is a unique complex geodesic connecting the point $z$ to $\xi$, however $\Omega_{\xi}$ is well-defined thanks to Proposition 11.2.
Remark 11.4.
It follows from Proposition 11.2 that if $z,w\in D$, then
$h_{\xi,z}(w)=\log|\Omega_{\xi}(z)|-\log|\Omega_{\xi}(w)|,$ (11.3)
hence $\Omega_{\xi}$ is continuous and its level sets are exactly the horospheres centered in $\xi$.
The function $\Omega^{\mathbb{D}}_{1}$ coincides with the classical (negative) Poisson kernel in the disc:
$\Omega^{\mathbb{D}}_{1}(\zeta)=-\frac{1-|\zeta|^{2}}{|1-\zeta|^{2}}.$
Moreover, the function $\Omega_{\xi}$ is a multiple of the classical Poisson kernel in the unit disc when restricted to a complex geodesic with endpoint $\xi$.
Corollary 11.5.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. Then, for all $\zeta\in\mathbb{D}$,
$\Omega_{\xi}^{D}(\varphi(\zeta))=|\Omega_{\xi}^{D}(\varphi(0))|\Omega_{1}^{\mathbb{D}}(\zeta)=\frac{\Omega_{1}^{\mathbb{D}}(\zeta)}{\varphi^{\prime}_{N}(1)}.$
Proof.
By (11.3),
$h^{\mathbb{D}}_{1,0}(\zeta)=-\log|\Omega^{\mathbb{D}}_{1}(\zeta)|,\quad\zeta\in\mathbb{D}.$
Again by (11.3) we obtain, for all $\zeta\in\mathbb{D}$,
$\log|\Omega^{\mathbb{D}}_{1}(\zeta)|=h^{\mathbb{D}}_{1,0}(\zeta)=h^{D}_{\xi,\varphi(0)}(\varphi(\zeta))=\log|\Omega^{D}_{\xi}(\varphi(0))|-\log|\Omega^{D}_{\xi}(\varphi(\zeta))|.$
Example 11.6.
Using the previous corollary and formula (4.2) one easily obtains an explicit formula for the pluricomplex Poisson kernel of the egg domain $\mathbb{E}_{m}$ at the point $\xi=(1,0)$, that is
$\Omega^{\mathbb{E}_{m}}_{\xi}(z_{0},z_{1})=-\frac{1-|z_{0}|^{2}-|z_{1}|^{m}}{|1-z_{0}|^{2}}.$
Corollary 11.7.
If $D\subset\subset\mathbb{C}^{d}$ is strongly convex with $C^{\infty}$ boundary, the function $\Omega_{\xi}$ coincides with the pluricomplex Poisson kernel introduced by Bracci–Patrizio–Trapani [22, 23].
Proof.
See (1.2) in [23]. ∎
If $D\subset\subset\mathbb{C}^{d}$ is strongly convex with $C^{\infty}$ boundary, the pluricomplex Poisson kernel $\Omega_{\xi}$ solves the following homogeneous Monge-Ampère equation with a simple singularity at the boundary:
[…]
In a forthcoming paper [10] with F. Bracci we prove that in convex domains of finite type the function $\Omega_{\xi}$ solves an equation analogous to (11.4).
Definition 11.8 (Normalized dilation).
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains and $f\colon D\to D^{\prime}$ be a holomorphic map. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. We define the normalized dilation of $f$ at $\xi$ as
$\alpha_{\xi}:=\lambda_{\xi,p,p^{\prime}}\frac{\Omega^{D}_{\xi}(p)}{\Omega^{D^{\prime}}_{\eta}(p^{\prime})}\in(0,+\infty),$
where $p\in D,p^{\prime}\in D^{\prime}$.
The next result shows that the normalized dilation is well-defined, that is, it does not depend on the chosen base-points $p\in D,p^{\prime}\in D^{\prime}.$
Lemma 11.9.
Let $D\subset\mathbb{C}^{d},D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map, and assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then the positive real number
$\lambda_{\xi,p,p^{\prime}}\frac{\Omega^{D}_{\xi}(p)}{\Omega^{D^{\prime}}_{\eta}(p^{\prime})}\in(0,+\infty)$
is independent on $p\in D,p^{\prime}\in D^{\prime}$.
Proof.
Let $p,q\in D,p^{\prime},q^{\prime}\in D^{\prime}$. By Lemma 10.5 and (11.3) we have
$\log\lambda_{\xi,p,p^{\prime}}-\log\lambda_{\xi,q,q^{\prime}}+\log|\Omega^{D}_{\xi}(p)|-\log|\Omega^{D}_{\xi}(q)|-\log|\Omega^{D^{\prime}}_{\eta}(p^{\prime})|+\log|\Omega^{D^{\prime}}_{\eta}(q^{\prime})|$
$=$ $h^{D}_{\xi,q}(p)+h^{D^{\prime}}_{\eta,p^{\prime}}(q^{\prime})+h^{D}_{\xi,p}(q)+h^{D^{\prime}}_{\eta,q^{\prime}}(p^{\prime})=0.$
Remark 11.10.
Notice that, if $D^{\prime}=\mathbb{D}$, then $\alpha_{\xi}=\lambda_{\xi,p,0}|\Omega^{D}_{\xi}(p)|$.
Remark 11.11.
It immediately follows from Corollary 10.13 that the normalized dilation satisfies a chain rule:
$\alpha_{\xi}(g\circ f)=\alpha_{\xi}(f)\cdot\alpha_{\eta}(g).$
Lemma 11.12.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. Then $1$ is a regular contact point of $\varphi$ and
$\alpha_{1}(\varphi)=\frac{1}{|\Omega^{D}_{\xi}(\varphi(0))|}=\varphi_{N}^{\prime}(1).$
Moreover, if $\tilde{\rho}\colon D\to\mathbb{D}$ is a left inverse of $\varphi$, then $\xi$ is a regular contact point of $\tilde{\rho}$, and
$\alpha_{\xi}(\tilde{\rho})=|\Omega^{D}_{\xi}(\varphi(0))|=\frac{1}{\varphi_{N}^{\prime}(1)}.$
Proof.
Notice that since $\varphi$ is a complex geodesic, we have $\lambda_{1,0,\varphi(0)}(\varphi)=1.$ Hence
$\alpha_{1}(\varphi)=\lambda_{1,0,\varphi(0)}(\varphi)\frac{\Omega_{1}^{\mathbb{D}}(0)}{\Omega_{\xi}^{D}(\varphi(0))}=\frac{1}{|\Omega^{D}_{\xi}(\varphi(0))|}.$
Similarly one obtains the second statement.
12. The Julia–Wolff–Carathéodory theorem
12.1. Statement of the result
Our main theorem is the following generalization of the Julia–Wolff-Carathéodory theorem.
Theorem 12.1.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains. Let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type with multitype respectively $(m_{j})_{j=0}^{d-1}$ and $(n_{i})_{i=0}^{q-1}$. Let $(v_{j})_{j=0}^{d-1}$ (resp. $(u_{i})_{i=0}^{q-1}$) be an orthonormal multitype basis at $\xi$ (resp. at $\eta$). Let $f\colon D\to D^{\prime}$ be a holomorphic map and assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then for all  $0\leq j\leq d-1$ and  $0\leq i\leq q-1$,
$\langle df_{z}(v_{j}),u_{i}\rangle=O_{K}\biggl(\delta_{D}(z)^{\frac{1}{n_{i}}-\frac{1}{m_{j}}}\biggr)$ (12.1)
and
  • (i)
    $K^{\prime}\textrm{-}\lim_{z\to\xi}\langle df_{z}(v_{0}),u_{0}\rangle=\alpha_{\xi}:=\lambda_{\xi,p,p^{\prime}}\frac{\Omega^{D}_{\xi}(p)}{\Omega^{D^{\prime}}_{\eta}(p^{\prime})}\in(0,+\infty)$;
  • (ii)
    $\langle df_{z}(v_{j}),u_{0}\rangle=o_{K^{\prime}}\biggl(\delta_{D}(z)^{1-\frac{1}{m_{j}}}\biggr)$ for all  $1\leq j\leq d-1$;
  • (iii)
    $\langle df_{z}(v_{0}),u_{i}\rangle=o_{K^{\prime}}\biggl(\delta_{D}(z)^{\frac{1}{n_{i}}-1}\biggr)$ for all  $1\leq i\leq q-1$.
Remark 12.2.
(cf. Remark 7.11). It is natural to wonder whether
$\langle df_{z}(v),u\rangle=O_{K}\biggl(\delta_{D}(z)^{\frac{1}{m_{\eta}(u)}-\frac{1}{m_{\xi}(v)}}\biggr)$ (12.2)
for all $v,u\in\mathbb{C}^{d}\setminus\{0\}$. This turns out to be false in general (see e.g. Example 12.9). However, it follows from (12.1) using simple linear algebra that
$\langle df_{z}(v),u\rangle=O_{K}\biggl(\delta_{D}(z)^{\frac{1}{M_{\eta}(u)}-\frac{1}{m_{\xi}(v)}}\biggr).$
It is worth stating Theorem 12.1 in the particular case of functions with values in the unit disc. This statement will actually play an important role in the proof of Theorem 12.1.
Theorem 12.3.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $f\colon D\to\mathbb{D}$ be a holomorphic function, and assume that $\xi$ is a regular contact point. Then for all $v\in\mathbb{C}^{d}\setminus\{0\}$,
$\frac{\partial f}{\partial v}(z)=O_{K}\left(\delta_{D}(z)^{1-\frac{1}{m_{\xi}(v)}}\right).$
Moreover
  • (i)
    $K^{\prime}\textrm{-}\lim_{z\to\xi}\frac{\partial f}{\partial n_{\xi}}(z)=\eta\alpha_{\xi}=\eta|\Omega^{D}_{\xi}(p)|\lambda_{\xi,p},$ where $\eta=K\textrm{-}\lim_{z\to\xi}f(z)$,
  • (ii)
    $\frac{\partial f}{\partial v}(z)=o_{K^{\prime}}\left(\delta_{D}(z)^{1-\frac{1}{m_{\xi}(v)}}\right)$ if $v\in T^{\mathbb{C}}_{\xi}\partial D\setminus\{0\}$.
Hence if $v\in\mathbb{C}^{d}$,
$K^{\prime}\textrm{-}\lim_{z\to\xi}\frac{\partial f}{\partial v}(z)=\langle v,n_{\xi}\rangle\eta\alpha_{\xi}=\langle v,n_{\xi}\rangle\eta\,|\Omega^{D}_{\xi}(p)|\lambda_{\xi,p}.$
We divide the proof of Theorem 12.1 in three steps.
12.2. First step: proof of (12.1)
Remark 12.4.
If $\zeta\in\mathbb{D}$, then
$\log(1-|\zeta|)\leq k_{\mathbb{D}}(0,\zeta)\leq\log 2-\log(1-|\zeta|).$ (12.3)
Lemma 12.5.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$, and let $\tilde{\rho}\colon D\to\mathbb{D}$ be a left inverse of $\varphi$. Then the function
$\frac{1-\tilde{\rho}(z)}{\delta_{D}(z)}$
and its reciprocal are $K$-bounded at $\xi$.
Proof.
Let $p:=\varphi(0)$. Denote $\rho:=\varphi\circ\tilde{\rho}\colon D\to D$ and notice that it is a holomorphic retraction with image $\varphi(\mathbb{D})$. Fix $R>0$ and let $z\in A^{D}(\varphi,R)$. Let $t^{*}\in[0,1)$ such that $k_{D}(z,\varphi(t^{*}))<R$, then $k_{D}(\rho(z),\varphi(t^{*}))<R$ and thus $\tilde{\rho}(z)\in A^{\mathbb{D}}(\mathrm{id}_{\mathbb{D}},R)$. By the proof of Proposition 7.4 we have $\tilde{\rho}(z)\in K_{0}^{\mathbb{D}}(1,e^{R})$. Using (7.8) and (12.3) we have
$|1-\tilde{\rho}(z)|\leq e^{R}(1-|\tilde{\rho}(z)|)\leq 2e^{R-k_{\mathbb{D}}(0,\tilde{\rho}(z))}=2e^{R-k_{D}(p,\rho(z))}.$
Let $V$ be the neighborhood of $\xi$ given by Lemma 3.8. Notice that $\xi$ is a regular contact point for $\rho$ with $K\textrm{-}\lim_{z\to\xi}\rho(z)=\xi,$ hence there exists a neighborhood $U$ of $\xi$ such that
$\rho(U\cap A^{D}(\varphi,R))\subset V.$
Then for all $z\in U\cap A^{D}(\varphi,R)$ we have
$|1-\tilde{\rho}(z)|\leq 2e^{c+R}\delta_{D}(\rho(z)),$
where $c>0$ is the constant given by Lemma 3.8. Since by Proposition 10.14 the function $\delta_{D}(\rho(z))/\delta_{D}(z)$ is $K$-bounded, it follows that $(1-\tilde{\rho}(z))/\penalty 50\delta_{D}(z)$ is $K$-bounded.
To prove that the reciprocal is also $K$-bounded, notice that $k_{D}(p,z)\geq k_{D}(p,\rho(z))=k_{\mathbb{D}}(0,\tilde{\rho}(z))$. If $z\in V\cap D$, by Lemma 3.8 and (12.3) we have
$\delta_{D}(z)\leq e^{c-k_{D}(p,z)}\leq e^{c-k_{\mathbb{D}}(0,\tilde{\rho}(z))}\leq e^{c}(1-|\tilde{\rho}(z)|)\leq e^{c}|1-\tilde{\rho}(z)|,$
hence the function $\delta_{D}(z)/(1-\tilde{\rho}(z))$ is bounded in $V\cap D$, and thus it is $K$-bounded. ∎
Proposition 12.6.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map. Assume that $\xi\in\partial D$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then $f$ sends strongly restricted curves in $D$ with endpoint $\xi$ to strongly restricted curves in $D^{\prime}$ with endpoint $\eta$.
Proof.
We first prove the result in the case $D=\mathbb{D}$ and $\xi=1\in\partial\mathbb{D}$. First of all we recall that by Proposition 4.9 there exists $f^{\prime}_{N}(1)>0$. Now let $\psi\colon\mathbb{D}\to D^{\prime}$ be a complex geodesic with endpoint $\eta$ such that $\psi^{\prime}_{N}(1)=f^{\prime}_{N}(1)$. Let $\gamma\colon[0,1)\to\mathbb{D}$ be a curve converging to $1$ non-tangentially, then by Corollary 6.7
$\lim_{t\to 1^{-}}k_{D}(f(\gamma(t)),\psi(\gamma(t)))=0,$
which implies that $f\circ\gamma$ is a strongly restricted curve in $D^{\prime}$ with endpoint $\eta$.
In the general case, let $\gamma\colon[0,1)\to D$ be a strongly restricted curve with endpoint $\xi$, and let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. By definition there exists a non-tangential curve $\tilde{\gamma}\colon[0,1)\to\mathbb{D}$ with endpoint 1 such that $\lim_{t\to 1^{-}}k_{D}(\gamma(t),\varphi(\tilde{\gamma}(t)))=0.$ By the first part of the proof applied to the map $f\circ\varphi\colon\mathbb{D}\to D$ there exists a complex geodesic $\psi\colon\mathbb{D}\to D^{\prime}$ with endpoint $\eta$ such that
$\lim_{t\to 1^{-}}k_{D^{\prime}}(f(\varphi(\tilde{\gamma}(t))),\psi(\tilde{\gamma}(t)))=0,$
and thus
$k_{D^{\prime}}(f(\gamma(t)),\psi(\tilde{\gamma}(t)))$ $\leq k_{D^{\prime}}(f(\gamma(t)),f(\varphi(\tilde{\gamma}(t))))+k_{D^{\prime}}(f(\varphi(\tilde{\gamma}(t))),\psi(\tilde{\gamma}(t)))$
$\leq k_{D}(\gamma(t),\varphi(\tilde{\gamma}(t)))+k_{D^{\prime}}(f(\varphi(\tilde{\gamma}(t))),\psi(\tilde{\gamma}(t)))\stackrel{{\scriptstyle t\to 1^{-}}}{{\longrightarrow}}0.$
Moreover, by Corollary 10.12 the curve $t\mapsto f(\gamma(t))$ is contained in a geodesic region with vertex $\eta$, hence it is strongly restricted. ∎
Proposition 12.7.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $(u_{i})_{i=0}^{q-1}$ be an orthonormal multitype basis at $\eta$ with multitype $(n_{i})_{i=0}^{q-1}.$ Let $f\colon D\to D^{\prime}$ be a holomorphic map. Assume that $\xi\in\partial D$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then
  • (i)
    $\langle f(z)-\eta,u_{i}\rangle=O_{K}(\delta_{D}(z)^{1/n_{i}}),$ for all   $0\leq i\leq q-1;$
  • (ii)
    $\langle f(z)-\eta,u_{i}\rangle=o_{K^{\prime}}(\delta_{D}(z)^{1/n_{i}}),$ for all   $1\leq i\leq q-1.$
Proof.
Let $R>0$, let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$. By Proposition 10.14 there exists $C_{1}\geq 0$ such that for all $z\in A(\varphi,R)$
$\delta_{D^{\prime}}(f(z))\leq C_{1}\delta_{D}(z).$
We now prove (i). We have
$\frac{|\langle f(z)-\eta,u_{i}\rangle|}{\delta_{D}(z)^{1/n_{i}}}\leq C_{1}^{1/n_{i}}\frac{|\langle f(z)-\eta,u_{i}\rangle|}{\delta_{D^{\prime}}(f(z))^{1/n_{i}}}.$
By Corollary 10.12 $f$ sends $K$-regions with vertex $\xi$ into $K$-regions with vertex $\eta$, so the right-hand side is bounded by Remark 7.10, which proves (i).
We prove (ii). Let $\tilde{\rho}\colon D\to\mathbb{D}$ be a left inverse of $\varphi$, and let $1\leq i\leq q-1.$ By Lemma 12.5 there exists $C_{2}\geq 0$ such that $\delta_{D}(\varphi(t))\leq C_{2}|1-\tilde{\rho}(\varphi(t))|$ for all $t\in[0,1)$, so
$\frac{|\langle f(\varphi(t))-\eta,u_{i}\rangle|}{|1-\tilde{\rho}(\varphi(t))|^{1/n_{i}}}\leq C_{2}^{1/n_{i}}\frac{|\langle f(\varphi(t))-\eta,u_{i}\rangle|}{\delta_{D}(\varphi(t))^{1/n_{i}}}\leq(C_{1}C_{2})^{1/n_{i}}\frac{|\langle f(\varphi(t))-\eta,u_{i}\rangle|}{\delta_{D^{\prime}}(f(\varphi(t)))^{1/n_{i}}}\stackrel{{\scriptstyle t\to 1^{-}}}{{\longrightarrow}}0,$
where we used Remark 7.10 and the fact that, thanks to Proposition 12.6, the curve $t\mapsto f(\varphi(t))$ is (strongly) restricted. By point (i) above the holomorphic function $\frac{\langle f(z)-\eta,u_{i}\rangle}{(1-\tilde{\rho}(z))^{1/n_{i}}}$ is $K$-bounded, and thus the Lindelöf principle (Theorem 9.4) yields
$K^{\prime}\textrm{-}\lim_{z\to\xi}\frac{\langle f(z)-\eta,u_{i}\rangle}{(1-\tilde{\rho}(z))^{1/n_{i}}}=0.$
By Lemma 12.5 we have (ii). ∎
Proof of (12.1) in Theorem 12.1.
Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$, and let $\tilde{\rho}\colon D\to\mathbb{D}$ be a left inverse of $\varphi$. Let $0\leq i\leq q-1$ and $v\in\mathbb{C}^{d}\setminus\{0\}.$ Thanks to Lemma 12.5 it is enough to prove that the function
$(1-\tilde{\rho}(z))^{\frac{1}{m_{\xi}(v)}-\frac{1}{n_{i}}}\langle df_{z}(v),u_{i}\rangle$
is $K$-bounded. Let $z\in A(\varphi,R)$. Since $D$ is convex, by [42, Lemma 11.1.2, Proposition 11.1.4, Theorem 11.2.1] there exists [^5] Notice that $\kappa_{\mathbb{D}}(0,1)=2$. a complex geodesic $\psi\colon\mathbb{D}\to D$ such that $\psi(0)=z$ and
$\psi^{\prime}(0)=2v/\penalty 50\kappa_{D}(z,v).$
Let $r\in(0,1)$, then by the Cauchy formula
$\langle df_{z}(v),u_{i}\rangle=\frac{\kappa_{D}(z,v)}{2}\langle df_{z}(\psi^{\prime}(0)),u_{i}\rangle=\frac{\kappa_{D}(z,v)}{4\pi i}\int_{|\zeta|=r}\frac{\langle f(\psi(\zeta))-\eta,u_{i}\rangle}{\zeta^{2}}d\zeta,$
so
$(1-\tilde{\rho}(z))^{1/m_{\xi}(v)-1/n_{i}}\langle df_{z}(v),u_{i}\rangle=$
$\frac{1}{4\pi}\int_{-\pi}^{\pi}\frac{\langle f(\psi(re^{i\theta}))-\eta,u_{i}\rangle}{(1-\tilde{\rho}(\psi(re^{i\theta})))^{1/n_{i}}}\cdot\left(\frac{1-\tilde{\rho}(\psi(re^{i\theta}))}{1-\tilde{\rho}(z)}\right)^{1/n_{i}}\cdot\frac{\kappa_{D}(z,v)(1-\tilde{\rho}(z))^{1/m_{\xi}(v)}}{re^{i\theta}}d\theta.$
Notice that, since $\psi$ is a complex geodesic, $k_{D}(\psi(re^{i\theta}),z)=k_{\mathbb{D}}(0,r)$, so $\psi(re^{i\theta})\in A(\varphi,R_{r})$ where $R_{r}:=R+k_{\mathbb{D}}(0,r)$, which implies by the proof of Proposition 7.4
$\tilde{\rho}(\psi(re^{i\theta}))\in A^{\mathbb{D}}(\mathrm{id}_{\mathbb{D}},R_{r})\subset K^{\mathbb{D}}_{0}(1,e^{R_{r}}).$
It follows that the first and third terms are bounded respectively by Proposition 12.7 and Theorem 8.2 combined with Lemma 12.5. Let us consider the second term. Using (7.8) we have
$\left|\frac{1-\tilde{\rho}(\psi(re^{i\theta}))}{1-\tilde{\rho}(z)}\right|$ $\leq e^{R_{r}}\frac{1-|\tilde{\rho}(\psi(re^{i\theta}))|}{1-|\tilde{\rho}(z)|}$
$\leq 2e^{R_{r}}\exp(k_{\mathbb{D}}(0,\tilde{\rho}(z))-k_{\mathbb{D}}(0,\tilde{\rho}(\psi(re^{i\theta}))))$
$\leq 2e^{R_{r}+k_{\mathbb{D}}(0,r)}.$
hence it is bounded. ∎
12.3. Second step: proof of Theorem 12.3
By the previous argument for all $v\in\mathbb{C}^{d}\setminus\{0\}$ the function $\delta_{D}(z)^{\frac{1}{m_{\xi}(v)}-1}\frac{\partial f}{\partial v}(z)$ is $K$-bounded.
Proof of (i) in Theorem 12.3.
By the Lindelöf Principle (Theorem 9.4) it is sufficient to prove that the limit (i) holds along the normal segment, which by Lemma 6.9 is a strongly restricted curve. First of all we prove that
$\lim_{t\to 0^{+}}\bar{\eta}\frac{\partial f}{\partial n_{\xi}}(\xi-tn_{\xi})$
exists and it is positive. Let $B\subset D$ be an Euclidean ball internally tangent to $\partial D$ at $\xi$. By Proposition 10.15 (noticing that, on the inner normal segment at $\xi$ we have $\delta_{D}=\delta_{B}$ close to $\xi$) it follows that the restriction $f|_{B}\colon B\to\mathbb{D}$ has a regular contact point at $\xi$ with $K$-limit $\eta$. Applying Rudin’s Julia-Wolff-Carathéodory theorem to $f|_{B}$ we obtain the existence and positivity of the limit. Denote by $\bar{\eta}\frac{\partial f}{\partial n_{\xi}}(\xi)$ such limit.
Now let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with $\varphi(0)=p$ and endpoint $\xi$, and define $\nu:=\varphi^{\prime}_{N}(1)n_{\xi}$. Consider the normal segment $\sigma\colon[t_{0},1)\to D$ given by $\sigma(t)=\xi+(t-1)\nu$, then by Proposition 10.10, Lemma 6.9 and the fundamental theorem of calculus we have
$\lambda_{\xi,p}$ $=\lim_{t\to 1^{-}}\exp(k_{\mathbb{D}}(0,t)-k_{\mathbb{D}}(f(\varphi(t)),0))=\lim_{t\to 1^{-}}\exp(k_{\mathbb{D}}(0,t)-k_{\mathbb{D}}(f(\sigma(t)),0))$
$=\lim_{t\to 1^{-}}\frac{1+t}{1-t}\cdot\frac{1-|f(\sigma(t))|}{1+|f(\sigma(t))|}=\lim_{t\to 1^{-}}\frac{1-|f(\sigma(t))|}{1-t}=\lim_{t\to 1^{-}}\frac{1}{1-t}\int_{t}^{1}\frac{{\rm Re}\left(\overline{f(\sigma(s)})\frac{\partial f}{\partial\nu}(\sigma(s))\right)}{|f(\sigma(s))|}ds$
$=\lim_{t\to 1^{-}}\frac{\varphi_{N}^{\prime}(1)}{1-t}\int_{t}^{1}\frac{{\rm Re}\left(\overline{f(\sigma(s)})\frac{\partial f}{\partial n_{\xi}}(\sigma(s))\right)}{|f(\sigma(s))|}ds=\varphi_{N}^{\prime}(1)\bar{\eta}\frac{\partial f}{\partial n_{\xi}}(\xi)=\frac{1}{|\Omega_{\xi}(p)|}\bar{\eta}\frac{\partial f}{\partial n_{\xi}}(\xi).$
Proof of (ii) in Theorem 12.3.
Set $m:=m_{\xi}(v)$ and $s:=1/m$. First of all we prove the result in dimension 2 for $f\colon\mathbb{E}_{m}\to\mathbb{D}$ where $\mathbb{E}_{m}$ is the egg domain, $\xi=(1,0)\in\partial\mathbb{E}_{m}$ is a regular contact point and $v=(0,1)$.
We start writing $f$ as follows:
$f(z_{0},z_{1})=f(z_{0},0)+\frac{\partial f}{\partial z_{1}}(z_{0},0)z_{1}+\sum_{k=2}^{\infty}\frac{1}{k!}\frac{\partial^{k}f}{\partial z_{1}^{k}}(z_{0},0)z_{1}^{k}.$
Define the holomorphic function $g\colon\mathbb{D}\to\mathbb{C}$ as
$g(z_{0})=\frac{1}{2}\frac{\partial f}{\partial z_{1}}(z_{0},0)(1-z_{0})^{s-1},$
where we are using the principal value of the $m$-th root. We need to show $g(t)\to 0$ if $t\to 1^{-}$. Define the holomorphic function $h\colon\mathbb{E}_{m}\to\mathbb{C}$ as
$h(z_{0},z_{1})=f(z_{0},0)+\frac{z_{1}}{2}\frac{\partial f}{\partial z_{1}}(z_{0},0)=f(z_{0},0)+z_{1}(1-z_{0})^{1-s}g(z_{0}).$
We want to prove that $h(\mathbb{E}_{m})\subseteq\mathbb{D}$. Fix $z_{0}\in\mathbb{D}$ and set $r:=\sqrt[m]{1-|z_{0}|^{2}}$. By the Schwarz-Pick lemma applied to $f(z_{0},\cdot)\colon r\mathbb{D}\to\mathbb{D}$, we have
$\frac{\left|\frac{\partial f}{\partial z_{1}}(z_{0},z_{1})\right|}{1-|f(z_{0},z_{1})|^{2}}\leq\frac{r}{r^{2}-|z_{1}|^{2}},$
which implies that at $z_{1}=0$ we have
$\left|\frac{\partial f}{\partial z_{1}}(z_{0},0)\right|\leq\frac{1-|f(z_{0},0)|^{2}}{r}.$
Finally, for all $z_{1}\in r\mathbb{D}$,
$|h(z_{0},z_{1})|$ $=\left|f(z_{0},0)+\frac{z_{1}}{2}\frac{\partial f}{\partial z_{1}}(z_{0},0)\right|\leq|f(z_{0},0)|+\frac{|z_{1}|}{2}\left|\frac{\partial f}{\partial z_{1}}(z_{0},0)\right|$
$\leq|f(z_{0},0)|+\frac{|z_{1}|}{2r}(1-|f(z_{0},0)|^{2})\leq|f(z_{0},0)|+\frac{1}{2}(1-|f(z_{0},0)|^{2})<1,$
so $h(\mathbb{E}_{m})\subseteq\mathbb{D}$.
Fix $a>0$, set $A:=\frac{a^{2}}{a^{2}+1}$ and define the curve
$\zeta_{a}\colon(A,1)\to\mathbb{D},\quad\zeta_{a}(t)=t+ia(1-t).$
Notice that for $t\in(A,1)$ we have $|\zeta_{a}(t)|^{2}<t$, so we can find $\eta_{a}(t)\in\mathbb{C}$ is such that
$1-t<|\eta_{a}(t)|^{m}<1-|\zeta_{a}(t)|^{2}$
and
$\eta_{a}(t)(1-\zeta_{a}(t))^{1-s}g(\zeta_{a}(t))\in\mathbb{R}_{\geq 0}.$
Define the function
$\sigma_{a}\colon(A,1)\to\mathbb{C}^{2},\quad\sigma_{a}(t)=(\zeta_{a}(t),\eta_{a}(t)).$
Notice that $\sigma_{a}((A,1))\subset\mathbb{E}_{m}$. Now we estimate $\limsup_{t\to 1^{-}}|g(\zeta_{a}(t))|.$ Consider the holomorphic function $\ell\colon\mathbb{D}\to\mathbb{D}$ defined as $\ell(\zeta)=f(\zeta,0)$. The point $1\in\partial\mathbb{D}$ is a regular contact point for $\ell$, hence by the one-dimensional Julia–Wolff–Carathéodory theorem (see e.g. [1, Theorem 1.2.7]) there exists $\alpha>0$ such that
$\frac{1-\ell(\zeta_{a}(t))}{1-\zeta_{a}(t)}=\alpha+o(1),$
and thus $\ell(\zeta_{a}(t))=1-(\alpha+o(1))(1-ia)(1-t)$. Since $h(\mathbb{E}_{m})\subseteq\mathbb{D}$ we have
$1$ $\geq{\operatorname{Re}\,}h(\sigma_{a}(t))=1-(\alpha+o(1))(1-t)+|\eta_{a}(t)||1-\zeta_{a}(t)|^{1-s}|g(\zeta_{a}(t))|$
$>1-(\alpha+o(1))(1-t)+|1-ia|^{1-s}|g(\zeta_{a}(t))|(1-t).$
So
$\limsup_{t\to 1^{-}}|g(\zeta_{a}(t))|\leq\frac{\alpha}{|1-ia|^{1-s}}.$
Similarly, we obtain $\limsup_{t\to 1^{-}}|g(\bar{\zeta}_{a}(t))|\leq\frac{\alpha}{|1-ia|^{1-s}},$ and since $g\colon\mathbb{D}\to\mathbb{C}$ is $K$-bounded at $1$ it follows that
$\limsup_{t\to 1^{-}}|g(t)|\leq\frac{\alpha}{|1-ia|^{1-s}}.$
Letting $a\to+\infty$ we have $\lim_{t\to 1^{-}}g(t)=0$.
For the general case, by Proposition 10.15 we can reduce to the case $d=2$ by cutting $D$ with the complex plane $\xi+\text{span}_{\mathbb{C}}\{n_{\xi},v\}$. Notice that the line type remains the same. Now $D$ has the following defining function near the origin
$r(z):={\operatorname{Re}\,}z_{0}+H(z_{1})+R(z)$
where $H\colon\mathbb{C}\to\mathbb{C}$ is a non-negative $m$-homogeneous ($H(tz_{1})=t^{m}H(z_{1})$) convex polynomial that is not zero and
$R(z)=o(|z_{0}|+|z_{1}|^{m}).$
Now we can find $\varepsilon>0$ small enough such that the egg domain
$E=\{z\in\mathbb{C}^{2}:|z_{0}+\varepsilon|^{2}+|z_{1}|^{m}<\varepsilon^{2}\}$
is contained in $D$. We conclude applying the first part of the proof to $f|_{E}$. ∎
This concludes the proof of Theorem 12.3. Combining it with Lemma 11.12, we immediately obtain the following corollary.
Corollary 12.8.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and let $\xi\in\partial D$ be a point of locally finite type. Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with endpoint $\xi$ and let $\tilde{\rho}\colon D\to\mathbb{D}$ be a left inverse of $\varphi$. Then if $v\in\mathbb{C}^{d}$,
$K^{\prime}\textrm{-}\lim_{z\to\xi}d\tilde{\rho}_{z}(v)=\langle v,n_{\xi}\rangle\alpha_{\xi}(\tilde{\rho})=\langle v,n_{\xi}\rangle|\Omega^{D}_{\xi}(\varphi(0))|.$
12.4. Third step: proof of Theorem 12.1
We are left with proving points (i),(ii),(iii) of Theorem 12.1.
Proof of (i) and (ii) in Theorem 12.1.
Let $\varphi\colon\mathbb{D}\to D$ be a complex geodesic with $\varphi(0)=p$ and endpoint $\xi$, and let $\psi\colon\mathbb{D}\to D^{\prime}$ be a complex geodesic with $\psi(0)=p^{\prime}$ and endpoint $\eta$. Let $\tilde{\rho}\colon D^{\prime}\to\mathbb{D}$ be a left inverse of $\psi$. First of all notice that $\xi$ is a regular contact point for the function $\tilde{\rho}\circ f$, with $K$-limit $1\in\partial\mathbb{D}$. Let $0\leq i\leq d-1$. For all $t\in[0,1)$ we have
$\frac{\partial(\tilde{\rho}\circ f)}{\partial v_{j}}(\varphi(t))=d\tilde{\rho}_{f(\varphi(t))}(df_{\varphi(t)}(v_{j}))=\sum_{i=0}^{q-1}\langle df_{\varphi(t)}(v_{j}),u_{i}\rangle d\tilde{\rho}_{f(\varphi(t))}(u_{i}).$
By Theorem 12.3 we have
$\frac{\partial(\tilde{\rho}\circ f)}{\partial v_{0}}(\varphi(t))\stackrel{{\scriptstyle t\to 1^{-}}}{{\to}}\alpha_{\xi}(\tilde{\rho}\circ f)=\alpha_{\eta}(\tilde{\rho})\alpha_{\xi}(f)=\alpha_{\xi}(f)|\Omega_{\eta}^{D^{\prime}}(p^{\prime})|,$
while for all $1\leq j\leq d-1$,
$\frac{\frac{\partial(\tilde{\rho}\circ f)}{\partial v_{j}}(\varphi(t))}{\delta_{D}(\varphi(t))^{1-\frac{1}{m_{j}}}}\stackrel{{\scriptstyle t\to 1^{-}}}{{\to}}0.$
Fix $1\leq i\leq q-1$. Notice that $f(\varphi(t))$ is a restricted curve. We have
$\frac{\langle df_{\varphi(t)}(v_{j}),u_{i}\rangle d\tilde{\rho}_{f(\varphi(t))}(u_{i})}{{\delta_{D}(\varphi(t))^{1-\frac{1}{m_{j}}}}}=\frac{\langle df_{\varphi(t)}(v_{j}),u_{i}\rangle}{\delta_{D}(\varphi(t))^{\frac{1}{n_{i}}-\frac{1}{m_{j}}}}\frac{d\tilde{\rho}_{f(\varphi(t))}(u_{i})}{\delta_{D^{\prime}}(f(\varphi(t)))^{1-\frac{1}{n_{i}}}}\frac{\delta_{D^{\prime}}(f(\varphi(t)))^{1-\frac{1}{n_{i}}}}{\delta_{D}(\varphi(t))^{1-\frac{1}{n_{i}}}}\stackrel{{\scriptstyle t\to 1^{-}}}{{\longrightarrow}}0,$
thanks to (12.1), Theorem 12.3, and Proposition 10.14.
Moreover, $d\tilde{\rho}_{f(\varphi(t))}(u_{0})\stackrel{{\scriptstyle t\to 1^{-}}}{{\to}}\alpha_{\eta}(p^{\prime})=|\Omega_{\eta}^{D^{\prime}}(p^{\prime})|>0.$ It follows that
$\langle df_{\varphi(t)}(v_{0}),u_{0}\rangle\stackrel{{\scriptstyle t\to 1^{-}}}{{\to}}\alpha_{\xi}(f),$
and for all $1\leq j\leq d-1$,
$\frac{\langle df_{\varphi(t)}(v_{j}),u_{0}\rangle}{\delta_{D}(\varphi(t))^{1-\frac{1}{m_{j}}}}\stackrel{{\scriptstyle t\to 1^{-}}}{{\to}}0,$
and the Lindelöf principle (Theorem 9.4) together with Lemma 12.5 yield the result.
Proof of (iii) in Theorem 12.1.
Define
$\varphi_{t}(\zeta)=\xi-\varepsilon(1-\zeta)(1-t)n_{\xi}.$
Notice that there exists $\varepsilon>0$ such that $\varphi_{t}(\mathbb{D})\subset D^{\prime}$ for all $t\in(0,1)$. We want to show that if we set $\sigma(t):=\varphi_{t}(0)=\xi-\varepsilon(1-t)n_{\xi}$, we have
$\lim_{t\to 1^{-}}(1-t)^{1-1/n_{i}}\langle df_{\sigma(t)}(n_{\xi}),u_{i}\rangle=0.$
Fix $r\in(0,1)$. Notice that for all $\theta$ the curve $t\mapsto\varphi_{t}(re^{i\theta})$ is contained in the cone with vertex $\xi$, axis $n_{\xi}$ and amplitude $2\arcsin(r)<\pi$. By [53, Proposition 2.4]
$\left|\log\frac{\delta_{D}(\varphi_{t}(re^{i\theta}))}{\delta_{D}(\varphi_{t}(0))}\right|\leq k_{D}(\varphi_{t}(re^{i\theta}),\varphi_{t}(0))\leq k_{\mathbb{D}}(0,r).$ (12.4)
Using the Cauchy formula we have
$(1-t)^{1-1/n_{i}}\langle df_{\sigma(t)}(n_{\xi}),u_{i}\rangle$ $=\frac{1}{\varepsilon(1-t)^{1/n_{i}}}\langle df_{\varphi_{t}(0)}(\varphi_{t}^{\prime}(0)),u_{i}\rangle$
$=\frac{1}{2\varepsilon\pi}\int_{-\pi}^{\pi}\frac{\langle f(\varphi_{t}(re^{i\theta}))-\eta,u_{i}\rangle}{\delta_{D}(\varphi_{t}(re^{i\theta}))^{1/n_{i}}}\cdot\left(\frac{\delta_{D}(\varphi_{t}(re^{i\theta}))}{(1-t)}\right)^{1/n_{i}}\cdot\frac{1}{re^{i\theta}}d\theta.$
Notice that the first factor in the integral is bounded by (i) in Proposition 12.7. The second factor is bounded by (12.4) since $\delta_{D}(\varphi_{t}(0))=\varepsilon(1-t)$. Finally, by (ii) in Proposition 12.7 the first factor converges pointwise to 0 as $t\to 1^{-}$, so we can conclude by the dominated convergence theorem. ∎
This ends the proof of Theorem 12.1.
12.5. Examples and consequences
Example 12.9.
We now give some examples that show that our results are sharp for holomorphic maps between egg domains. Let $m\geq 2$ be an even integer and let $\mathbb{E}_{m}$ be the egg domain (see (4.1)). Now consider the holomorphic function $\vartheta\colon\mathbb{D}\to\mathbb{D}$ given by
$\vartheta(\zeta)=\exp\left(-\frac{\pi}{2}-i\log(1-\zeta)\right).$
Notice that as $\zeta\to 1$ the function $\vartheta(\zeta)$ spirals around the origin without limit. Moreover
$\vartheta^{\prime}(z)=\frac{i}{1-\zeta}\vartheta(\zeta).$
The function $\varphi\colon\mathbb{D}\to\mathbb{E}_{m}$ given by $\varphi(\zeta)=(\zeta,0)$ is a complex geodesic with endpoint $\xi:=(1,0)$ and the function $\tilde{\rho}\colon\mathbb{E}_{m}\to\mathbb{D}$ given by $\tilde{\rho}(z_{0},z_{1})=z_{0}$ is a left inverse of $\varphi$, so by Lemma 12.5 the function $\delta_{\mathbb{E}_{m}}(z)/(1-z_{0})$ and its reciprocal are $K$-bounded at $\xi$. Notice also the quantity $\frac{z_{1}^{m}}{1-z_{0}}$ is $K$-bounded at $\xi$ by Proposition 12.7.
Let $f\colon\mathbb{E}_{m}\to\mathbb{D}$ be given by
$f(z)=z_{0}+\frac{1}{2}z_{1}^{m}\vartheta(z_{0}).$
Notice that $\xi$ is a regular contact point for $f$. Then
$\frac{\partial f}{\partial z_{0}}(z)=1+\frac{i}{2}\frac{z_{1}^{m}}{1-z_{0}}\vartheta(z_{0})$
and
$(1-z_{0})^{\frac{1}{m}-1}\frac{\partial f}{\partial z_{1}}(z)=\frac{m}{2}\left(\frac{z_{1}^{m}}{1-z_{0}}\right)^{\frac{m-1}{m}}\vartheta(z_{0}).$
The two functions are $K$-bounded at $\xi$ with sharp exponents and have restricted $K$-limit at $\xi$, but they do not have $K$-limit at $\xi$. Indeed they do not have limit along the curve $\gamma_{\lambda}(t)=(t,\lambda\sqrt[m]{1-t^{2}})$, with $\lambda\in\mathbb{D}^{*}$, which $K$-converges to $\xi$ but does not $K^{\prime}$-converge to $\xi$.
Now we discuss two examples of holomorphic mappings $f\colon\mathbb{E}_{m_{1}}\to\mathbb{E}_{m_{2}}$ depending on the order relation between $m_{1}$ and $m_{2}$.
  • Let $2\leq m_{1}\leq m_{2}$. Let $f\colon\mathbb{E}_{m_{1}}\to\mathbb{E}_{m_{2}}$ be given by
    $f(z)=\left(z_{0},\frac{1}{2^{1/m_{1}}}\frac{z_{1}}{(1-z_{0})^{\frac{m_{2}-m_{1}}{m_{1}m_{2}}}}\vartheta(z_{0})\right).$
    Then
    $(1-z_{0})^{1-\frac{1}{m_{2}}}\frac{\partial f_{2}}{\partial z_{0}}(z)=\frac{1}{2^{1/m_{1}}}\left(\frac{m_{2}-m_{1}}{m_{1}m_{2}}+i\right)\left(\frac{z_{1}^{m_{1}}}{1-z_{0}}\right)^{\frac{1}{m_{1}}}\vartheta(z_{0}),$
    $(1-z_{0})^{\frac{1}{m_{1}}-\frac{1}{m_{2}}}\frac{\partial f_{2}}{\partial z_{1}}(z)=\frac{1}{2^{1/m_{1}}}\vartheta(z_{0}).$
  • Let $2\leq m_{2}\leq m_{1}$ and fix $r>0$. Let $f\colon\mathbb{E}_{m_{1}}\to\mathbb{E}_{m_{2}}$ be given by
    $f(z)=\left(\frac{z_{0}+r}{1+r},\frac{1}{2^{1/m_{1}}}\frac{r^{1/m_{2}}}{(1+r)^{2/m_{2}}}z_{1}(1-z_{0})^{\frac{m_{1}-m_{2}}{m_{1}m_{2}}}\vartheta(z_{0})\right).$
    Then
    $(1-z_{0})^{1-\frac{1}{m_{2}}}\frac{\partial f_{2}}{\partial z_{0}}(z)=\frac{1}{2^{1/m_{1}}}\frac{r^{1/m_{2}}}{(1+r)^{2/m_{2}}}\left(\frac{m_{2}-m_{1}}{m_{1}m_{2}}+i\right)\left(\frac{z_{1}^{m_{1}}}{1-z_{0}}\right)^{\frac{1}{m_{1}}}\vartheta(z_{0}),$
    $(1-z_{0})^{\frac{1}{m_{1}}-\frac{1}{m_{2}}}\frac{\partial f_{2}}{\partial z_{1}}(z)=\frac{1}{2^{1/m_{1}}}\frac{r^{1/m_{2}}}{(1+r)^{2/m_{2}}}\vartheta(z_{0}).$
In both cases, $\xi$ is a regular contact point. Notice that for all the functions we have sharp exponents for the $K$-boundedness. Moreover, $(1-z_{0})^{1-\frac{1}{m_{2}}}\frac{\partial f_{2}}{\partial z_{0}}(z)$ has restricted $K$-limit at $\xi$ but does not have $K$-limit at $\xi$. Finally, $(1-z_{0})^{\frac{1}{m_{1}}-\frac{1}{m_{2}}}\frac{\partial f_{2}}{\partial z_{1}}(z)$ does not even have radial limit at $\xi$.
[…]
We conclude with some consequences. The first is an asymptotic estimate, depending only on the multitypes at $\xi$ and $\eta$, of the complex Jacobian of $f$ when the dimension of $D$ is equal to the dimension of $D^{\prime}$. This generalizes the result [59, Corollary, p.178], which shows that if $D$ and $D^{\prime}$ are balls, then the complex Jacobian of $f$ is $K$-bounded.
Corollary 12.10.
Let $D,D^{\prime}\subset\mathbb{C}^{d}$ be $\mathbb{C}$-proper convex domains, and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type with multitype respectively $(m_{j})_{j=0}^{d-1}$ and $(n_{j})_{j=0}^{d-1}$. Let $f\colon D\to D^{\prime}$ be a holomorphic map, and assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$. Then
$\det df_{z}=O_{K}\left(\delta_{D}(z)^{\sum_{j=0}^{d-1}\frac{1}{n_{j}}-\frac{1}{m_{j}}}\right).$
Proof.
Let $(v_{j})_{j=0}^{d-1}$ (resp. $(u_{j})_{j=0}^{d-1}$) be an orthonormal multitype basis at $\xi$ (resp. at $\eta$). The result follows computing the determinant of the Jacobian matrix of $f$ using Theorem 12.1. ∎
Remark 12.11.
If $H$ denotes the harmonic mean, then clearly
$\sum_{j=0}^{d-1}\frac{1}{n_{j}}-\frac{1}{m_{j}}=\frac{d}{H(n_{0},\dots,n_{d-1})}-\frac{d}{H(m_{0},\dots,m_{d-1})}.$
In particular, if the harmonic mean of the multitype at $\eta$ is strictly smaller than the harmonic mean of the multitype at $\xi$, then
$K\textrm{-}\lim_{z\to\xi}\det df_{z}=0$
Theorem 12.1 has an interesting consequence for self-maps.
Corollary 12.12.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and $\xi\in\partial D$ a point of locally finite type and let $f\colon D\to D$ be a holomorphic self-map. If $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\xi$, then for all $p\in D$
$K^{\prime}\textrm{-}\lim_{z\to\xi}\langle df_{z}(n_{\xi}),n_{\xi}\rangle=\lambda_{\xi,p,p}(f).$
We now prove a version of the Julia Lemma for the pluricomplex Poisson kernel.
Proposition 12.13.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains, and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map.
  • (i)
    If the function
    $z\mapsto\frac{\Omega^{D}_{\xi}(z)}{\Omega^{D^{\prime}}_{\eta}(f(z))}$
    is bounded from above, then $\xi$ is a regular contact point for $f$ with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$.
  • (ii)
    If $\xi$ is a regular contact point for $f$ with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$, then
    $\sup_{z\in D}\frac{\Omega^{D}_{\xi}(z)}{\Omega^{D^{\prime}}_{\eta}(f(z))}=\alpha_{\xi}=K^{\prime}\textrm{-}\lim_{z\to\xi}\langle df_{z}(n_{\xi}),n_{\eta}\rangle.$
Proof.
Point (i) follows from the implication $(3)\Rightarrow(4)$ in the Julia Lemma (Theorem 10.6) and from (11.3). We prove (ii). Fix $p\in D,p^{\prime}\in D^{\prime}$. From (10.4) and (11.3) we obtain
$\sup_{z\in D}\left(\log|\Omega^{D^{\prime}}_{\eta}(p^{\prime})|-\log|\Omega^{D^{\prime}}_{\eta}(f(z))|-\log|\Omega^{D}_{\xi}(p)|+\log|\Omega^{D}_{\xi}(z)|\right)=\log\lambda_{\xi,p,p^{\prime}},$
and the result immediately follows. ∎
Remark 12.14.
Point (ii) of the previous proposition yields an interesting estimate from below of $K^{\prime}\textrm{-}\lim_{z\to\xi}\langle df_{z}(n_{\xi}),n_{\eta}\rangle$. As an example, if $f\colon\mathbb{E}_{m}\to\mathbb{E}_{m}$ if a holomorphic self-map such that $\xi=(1,0)$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\xi$, then by Example 11.6 if $f(0)=(z_{0},z_{1})$ is the image of the origin, we have
$K^{\prime}\textrm{-}\lim_{z\to\xi}\langle df_{z}(e_{0}),e_{0}\rangle\geq\frac{|1-z_{0}|^{2}}{1-|z_{0}|^{2}-|z_{1}|^{m}}.$
Corollary 12.15.
Let $D,D^{\prime}\subset\mathbb{C}^{d}$ be $\mathbb{C}$-proper convex domains, and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a biholomorphism and assume that there exists a sequence $(z_{n})$ in $D$ which is $K$-converging to $\xi$ such that $f(z_{n})\to\eta$. Then for all $w\in D$
$\frac{\Omega_{\eta}(w)}{\Omega_{\xi}(f(w))}=\alpha_{\xi}=K^{\prime}\textrm{-}\lim_{z\to\xi}\langle df_{z}(n_{\xi}),n_{\eta}\rangle.$
Proof.
It is enough to notice that, since $f$ is a biholomorphism,
$\log\lambda_{\xi,w,f(w)}=\liminf_{z\to\xi}k_{D}(z,w)-k_{D^{\prime}}(f(z),f(w))=0$
for all $w\in D$. ∎
We conclude with two extrinsic characterizations of regular contact points (notice the analogy with [1, Propositions 1.2.6, 1.2.8]).
Corollary 12.16.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain and $\xi\in\partial D$ be a point of locally finite type. Let $f\colon D\to\mathbb{D}$ be a holomorphic map, then $\xi$ is a regular contact point if and only if
$\lim_{t\to 0^{+}}|f(\xi-tn_{\xi})|=1$ (12.5)
and
$\limsup_{t\to 0^{+}}\left|\frac{\partial f}{\partial n_{\xi}}(\xi-tn_{\xi})\right|<\infty.$ (12.6)
Proof.
Assume that $\xi$ is a regular contact point for $f$, then (12.5) follows from the definition while (12.6) follows from Theorem 12.3.
Conversely, assume (12.5) and (12.6). Let $t_{0}>0$ and $M>0$ such that for all $0\leq t\leq t_{0}$ we have $\left|\frac{\partial f}{\partial n_{\xi}}(\xi-tn_{\xi})\right|\leq M.$ Up to taking a smaller $t_{0}$ we have for all $0\leq t\leq t_{0}$
$\frac{\delta_{\mathbb{D}}(f(\xi-tn_{\xi}))}{t}$ $=\frac{1-|f(\xi-tn_{\xi})|}{t}$
$=\frac{1}{t}\int_{0}^{t}\frac{{\operatorname{Re}\,}(\overline{f(\xi-sn_{\xi})}\frac{\partial f}{\partial n_{\xi}}(\xi-sn_{\xi}))}{|f(\xi-sn_{\xi})|}ds$
$\leq\frac{1}{t}\int_{0}^{t}\left|\frac{\partial f}{\partial n_{\xi}}(\xi-sn_{\xi})\right|ds\leq M,$
which implies by Proposition 10.15 that $\xi$ is a regular contact point.
Corollary 12.17.
Let $D\subset\mathbb{C}^{d}$ and $D^{\prime}\subset\mathbb{C}^{q}$ be $\mathbb{C}$-proper convex domains and let $\xi\in\partial D,\eta\in\partial D^{\prime}$ be points of locally finite type. Let $f\colon D\to D^{\prime}$ be a holomorphic map, then $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$ if and only if
$\lim_{t\to 0^{+}}f(\xi-tn_{\xi})=\eta$ (12.7)
and
$\limsup_{t\to 0^{+}}|\langle df_{\xi-tn_{\xi}}(n_{\xi}),n_{\eta}\rangle|<\infty.$ (12.8)
Proof.
Assume that $\xi$ is a regular contact point with $K\textrm{-}\lim_{z\to\xi}f(z)=\eta$ . Then (12.7) is obvious and (12.8) follows from (i) in Theorem 12.1.
Conversely, assume (12.7) and (12.8). Assume $\eta=0$, $n_{\eta}=e_{0}$. Let $\pi\colon\mathbb{C}^{q}\to\mathbb{C}$ be the projection on the first component. By the convexity of $D^{\prime}$ we have $\pi(D^{\prime})\subseteq\mathbb{H}$. Notice that
$\frac{\partial(\pi\circ f)}{\partial n_{\xi}}(z)=\langle df_{z}(n_{\xi}),n_{\eta}\rangle$
so by Corollary 12.16 $\xi$ is a regular contact point for the function $\mathscr{C}^{-1}\circ\pi\circ f\colon D\to\mathbb{D}$ (and thus also for $\pi\circ f\colon D\to\mathbb{H}$). Now fix $p\in D$ and $p^{\prime}\in D^{\prime}$, then
$\liminf_{t\to 0^{+}}k_{D}(p,\xi-tn_{\xi})-k_{D^{\prime}}(p^{\prime},f(\xi-tn_{\xi}))\leq\liminf_{t\to 0^{+}}k_{D}(p,\xi-tn_{\xi})-k_{\mathbb{H}}(\pi(p^{\prime}),\pi(f(\xi-tn_{\xi})))<+\infty,$
hence $\lambda_{\xi,p,p^{\prime}}(f)<+\infty$. Finally, the Julia Lemma (Proposition 10.6) yields the result. ∎
Our results also give us information at points of infinite type.
Corollary 12.18.
Let $D\subset\mathbb{C}^{d}$ be a $\mathbb{C}$-proper convex domain, and let $\xi\in\partial D$ be a smooth point of infinite line type. Let $f\colon D\to\mathbb{D}$ be a holomorphic function such that
$\lim_{t\to 0^{+}}|f(\xi-tn_{\xi})|=1$
and
$\limsup_{t\to 0^{+}}\left|\frac{\partial f}{\partial n_{\xi}}(\xi-tn_{\xi})\right|<\infty.$
Let $v\in\mathbb{C}^{d}$ with $m_{\xi}(v)=+\infty$, then for all $s<1$ and for all sequences $(z_{n})$ converging non-tangentially to $\xi$ we have
$\frac{\frac{\partial f}{\partial v}(z_{n})}{\delta_{D}(z_{n})^{s}}\stackrel{{\scriptstyle n\to+\infty}}{{\longrightarrow}}0.$
Proof.
First of all assume that $\xi$ is the origin and $n_{\xi}=e_{0}$. Let $r\colon\mathbb{C}^{d}\to\mathbb{R}$ be a smooth convex defining function of $D$ (for instance, the signed distance function), and define
$r_{L}(z):=r(z)+\sum_{j=1}^{d-1}|z_{j}|^{L}$
where $L$ is an even integer. Notice that the origin is a point of locally finite type for $D_{L}:=\{r_{L}<0\}$ with line type $L$. Moreover, since $r\leq r_{L}$ we have $D_{L}\subset D$. Now for all $v\in\mathbb{C}^{d}$, by definition we have
$m^{D_{L}}_{\xi}(v)=\min\{m^{D}_{\xi}(v),L\}.$
By Corollary 12.16 the origin is a regular contact point for the map $f|_{D_{L}}:D_{L}\to\mathbb{D}$. Then by Theorem 12.3 and recalling that a sequences which converges non-tangentially to $\xi$ is contained in a $K$-region of $D_{L}$ (Corollary 7.15), we have that for all sequence $(z_{n})$ converging non-tangentially to $\xi$
$\frac{\partial f}{\partial v}(z_{n})\delta_{D}(z_{n})^{\frac{1}{\min\{m_{\xi}(v),L\}}-1}$
is bounded. We conclude letting $L\to+\infty$. ∎
Acknowledgement. We want to thank the anonymous referee for the careful reading of the paper and for many useful comments.